### v1

# Quantum optical coherence theory based on Feynman’s path integral

Abstract
The first, second and higher-order interference of classical and nonclassical light is interpreted in quantum optical coherence theory based on Feynman’s path integral. Compared to the existed classical optical coherence theory based on Maxwell’s electromagnetic theory and Glauber’s quantum optical coherence theory based on wave mechanics formulation of quantum mechanics, quantum optical coherence theory based on Feynman’s path integral provides a new tool to study optical coherence and has the advantage of understanding the connection between the physics and mathematical calculations. For instance, based on the results of transient first-order interference of two independent light beams, it is predicted that the well-accepted electric field model for thermal light introduced by many optical textbooks may not be accurate. The physics of two-photon bunching of thermal light and Hong-Ou-Mandel dip of entangled photon pairs is the same, which can be interpreted by constructive and destructive two-photon interference, respectively. The second-order interference of two independent light beams are introduced in detail. Interpreting the interference of light in quantum optical coherence theory based on Feynman’s path integral is helpful to understand the physics of light, and may eventually lead to the answer of the question puzzling us for a long time: what is photon?
I Introduction
Ever since the interference of light introduced by Young in 1801 Young (1802), optical coherence has been one of the most important topics in the development of physics. With the contributions of generations of scientists, there are two main theories of optical coherence Born and Wolf (2013); Mandel and Wolf (1995). One is classical optical coherence theory based on Maxwell’s electromagnetic theory Born and Wolf (2013). The other one is Glauber’s quantum optical coherence theory based on wave mechanics formulation of quantum mechanics Glauber (1963a, c). It was proved by Glauber and Sudarshan independently that classical and quantum optical coherence theories are mathematically equivalent for classical light Sudarshan (1963); Glauber (1963b). However, classical optical coherence theory can not interpret non-classical phenomena such as photoelectric effect Hertz (1887); Einstein (1905), photon antibuching Miller and Mishkin (1967); Stoler (1974); Kimble et al. (1977), Bell inequality violation Bell (1964, 1966) and so on. On the other hand, quantum optical coherence theory can interpret both classical and non-classical phenomena.
Since Glauber’s quantum optical coherence theory can explain all the optical coherence phenomena and be a great success Glauber (2006), what is the reason to introduce another quantum optical coherence theory based on Feynman’s path integral? This is the question we asked ourselves since the first day we employed Feynman’s path integral to explain the subwavelength interference of light Liu and Zhang (2010). Similar question bothered Young and Feynman Young (1802); Feynman (1948), too. In the beginning of Young’s paper on the theory of light and colours Young (1802), he explained why he employed wave theory to interpret light and colours: Although the invention of plausible hypotheses, independent of any connection with experimental observations, can be of very little use in the promotion of natural knowledge; yet the discovery of simple and uniform principles, by which a great number of apparently heterogeneous phenomena are reduced to coherent and universal laws, must ever be allowed to be of considerable importance towards the improvement of the human intellect.
When Feynman introduced path integral formulation of quantum mechanics, he also faced the same question. There had been wave mechanics and matrix mechanics formulations of quantum mechanics. What is the reason to introduce a third and equivalent formulation of quantum mechanics. Feynman’s answer to this question is Feynman (1948): The formulation is mathematically equivalent to the more usual formulations. There are, therefore, no fundamentally new results. However, there is a pleasure in recognizing old things from a new point of view. Also, there are problems for which the new point of view offers a distinct advantage.
The answers offered by Young and Feynman are also the reasons why we introduced quantum optical coherence theory based on Feynman’s path integral besides Glauber’s quantum optical coherence theory. As introduced in Sect. III, quantum optical coherence theory based on Feynman’s path integral can indeed offer deeper insight about the first- and second-order interference of two independent light beams, which is helpful to understand the physics of coherence and the concept of photon.
The remaining parts of the paper are organized as follows. Basic concepts in quantum optical coherence theory based on Feynman’s path integral are introduced in Sect. II. The first- and second-order interference of light are introduced in Sects. III and IV, respectively. Generalizing the method developed in Sects. III and IV to the third- and higher-order interference of light can be found in Sect. V. Section VI introduces the quantum atomic coherence theory based on Feynman’s path integral by analogy of the optical one. Section VII summarizes the conclusions.
II Basic concepts in quantum optical coherence theory based on Feynman’s path integral
II.1 History of quantum optical coherence theory based on Feynman’s path integral
The history of optical coherence began in the 17th century when Grimaldi discovered diffraction of light Hecht (2012). With the contributions from Young Young (1802, 1804), Fresnel Fresnel (1819), and many other scientists Hecht (2012); Born and Wolf (2013), classical optical coherence theory has been established based on Maxwell’s electromagnetic theory Maxwell (1865). We will not introduce the history of optical coherence since there are many good reviews about this topic Born and Wolf (2013); Hecht (2012); Mandel and Wolf (1965).
The history of quantum optical coherence theory dates back to Dirac and Feynman. In his famous book on quantum mechanics, Dirac discussed the first-order interference of light with the concept of photon and concluded “Each photon then interferes only with itself. Interferences between two different photons never occurs Dirac (1930).” Feynman employed path integral to discuss Young’s double-slit interference of light in his lectures Feynman et al. (2011). All the optical coherence discussions were limited to the first-order before 1956. Both quantum and classical optical coherence theories can explain the first-order interference of light. The need for quantum optical coherence theory was not so urgent. In 1956, Hanbury Brown and Twiss found that photons emitted by thermal light source are not independent Brown and Twiss (1956a, b), which is known as two-photon bunching of thermal light. People began to realize that there are second- and higher-order interference of light besides the first-order. In 1963, Glauber published a series of papers explaining optical coherence in quantum theory Glauber (1963a, c, b), which laid the foundation of quantum optical coherence theory. Glauber was awarded 2005 Nobel physics prize for his contribution to quantum theory of optical coherence Glauber (2006). Glauber’s quantum optical coherence theory was employed by most scientists and introduced in many quantum optics books Mandel and Wolf (1995); Scully and Zubairy (1997); Loudon (2000); Shih (2020).
Glauber’s quantum optical coherence theory is based on wave mechanics formulation of quantum mechanics Schrödinger (1926). There are two more equivalent formulations of quantum mechanics called matrix mechanics Heisenberg (1925, 1949) and Feynman’s path integral Feynman (1948); Feynman et al. (2010). In principle, one can establish quantum optical coherence theory based on matrix mechanics and Feynman’s path integral formulations, too. In the following parts of this paper, we will introduce quantum optical coherence theory based on Feynman’s path integral.
Quantum optical coherence theory based on Feynman’s path integral was first introduced by Feynman himself. In his lectures on physics, Feynman examined Young’s double-slit interference of electrons, which is similar as the Young’s double-slit interference of light Feynman et al. (2011). Shortly after Hanbury Brown and Twiss’s experiments Brown and Twiss (1956a, b), Fano employed Feynman’s path integral to interpret two-photon bunching of thermal light Fano (1961). As shown in Fig. 1, there are two different ways for two photons emitted by atoms a and b and absorbed by atoms c and d. One is photon emitted by atom a is absorbed by atom c and photon emitted by atom b is absorbed by atom d. The second ways is emitted by atom a is absorbed by atom d and photon emitted by atom b is absorbed by atom c. Two-photon bunching of thermal light is interpreted as the superposition of these two different two-photon probability amplitudes corresponding these two different ways, which is later named as two-photon interference Shih (2020). In his lecture on quantum electrodynamics, Feynman also employed similar idea to interpret two-photon bunching of thermal light Feynman (2006).
Figure 1: Interpretation of two-photon bunching of thermal light in Feynman’s path integral. From Fano (1961)
As mentioned before, Glauber’s quantum optical coherence theory was employed to study coherence in most researches in quantum optics. There are limited number of studies employing Feynman’s path integral to study optical coherence. Hillery and Zubairy employed Feynman’s path integral to discuss some problems in nonlinear optics and calculated the path-integral representation for the coherent-state propagator Hillery and Zubairy (1982). Yabuki employed Feynman’s path integral to discuss the nonlinear contribution in Young’s double-slit interferometer of light Yabuki (1986). Barut and Basri discussed single-photon interference in single- and double-slit interferometers with Feynman’s path integral Barut and Basri (1992). Braidotti et. al. employed Feynman’s path integral to discuss non-paraxial optics problem Braidotti et al. (2018). Wallner employed Feynman’s path integral to discuss the interference of entangled photon pairs Wallner (2021).
We have employed employed Feynman’s path integral to discuss the first- and second-order interference of light since 2010 Liu and Zhang (2010, 2011); Liu et al. (2013, 2014b, 2014a, 2015b, 2015a); Bai et al. (2016); Zhou et al. (2017, 2019a); Luo et al. (2021a, b). Some interesting conclusions have been drawn. For instance, it is predicted that there is no transient first-order interference pattern by superposing two independent thermal light beams Bai et al. (2016), which is different from the observed transient first-order interference pattern by superposing two independent laser light beams Javan et al. (1962); Magyar and Mandel (1963); Lipsett and Mandel (1963). This distinction leads to an interesting conclusion that electric field models for laser and classical light within the coherence time should be different, which is contrary to the usual electric model introduced in classical optics textbooks Born and Wolf (2013). Detail discussions about this problem can be found in Sect. III.
The researches on optical coherence with Feynman’s path integral did by other groups Fano (1961); Feynman et al. (2011); Feynman (2006); Hillery and Zubairy (1982); Yabuki (1986); Barut and Basri (1992); Braidotti et al. (2018); Wallner (2021) and our group Liu and Zhang (2010, 2011); Liu et al. (2013, 2014b, 2014a, 2015b, 2015a); Bai et al. (2016); Zhou et al. (2017, 2019a); Luo et al. (2021a, b) are all about special interference phenomena of light. No systematic study about quantum optical coherence theory based on Feynman’s path integral is reported. This paper will present a comprehensive introduction about quantum optical coherence theory based on Feynman’s path integral by re-visiting the first-, second-, third- and higher-order interference of light.
II.2 Principles of quantum optical coherence theory based on Feynman’s path integral
The key to discuss optical coherence in Feynman’s path integral is superposition principle. As Feynman pointed out in his lectures, it contains the only mystery (of quantum mechanics) Feynman et al. (2011). Feynman summarized how to calculate the probability of an event with superposition principle Feynman et al. (2011):
(I) The probability of an event in an ideal experiment is given by the square of the absolute value of a complex number $\phi$ which is called the probability amplitude: $P$=probability,
$\phi$=probability amplitude,
$P=|\phi|^{2}$
(II) When an event can occur in several alternative ways, the probability amplitudes for the event is the sum of the probability amplitudes for each way considered separately. There is interference: $\phi=\phi_{1}+\phi_{2}$,
$P=|\phi_{1}+\phi_{2}|^{2}$
(III) If an experiment is performed which is capable of determining whether one or another alternative is actually taken, the probability of the event is the sum of the probabilities for each alternative. The interference is lost: $P=P_{1}+P_{2}$
In order to calculate the probability of an event with the above assumptions, there are two premises need to be figured out. The first one is how to calculate the probability amplitude $\phi$ for each alternative way. The second one is how to decide whether two different alternative ways for an event are distinguishable or not.
II.2.1 Probability amplitude for a path
The probability amplitude for a path is first given by Dirac by analogy of classical theory of dynamics Dirac (1930). Dirac assumed that the solution of schrödinger wave equation is as follows,
$\langle q_{t_{b}}|q_{t_{a}}\rangle=e^{iS/\hbar},$ (1)
where $q_{t_{a}}$ and $q_{t_{b}}$ are canonical coordinates of particle at time $t_{a}$ and $t_{b}$, respectively, $\hbar=h/(2\pi)$ and $h$ is Planck constant, $S$ is classical action given by Taylor and Taylor (2005)
[…]
$x(t)$ represents the path, $L(x,\dot{x},t)$ is the Lagrangian of the particle taking the path $x(t)$, $\dot{x}$ is short for $\partial x/\partial t$. Equation (1) means that the probability amplitude for a particle starting from $q_{t_{a}}$ at time $t_{a}$ and ending at $q_{t_{b}}$ at time $t_{b}$ equals $e^{iS/\hbar}$.
Based on the above assumption, Dirac concluded that the probability amplitude can be expressed as a result of integral over all the possible paths
[…]
×$\times\langle q_{t_{2}}|q_{t_{1}}\rangle dq_{t_{1}}\langle q_{t_{1}}|q_{t_{a}}\rangle,$
where $t_{j}=t_{a}+(j-1)\times(t_{b}-t_{a})/m$ ($j=1,2,...,m$) and $q_{t_{j}}$ is the corresponding canonical coordinate at time $t_{j}$. Equation (2) is in fact path integral. The same equation can be found in Feynman’s book Feynman et al. (2010). The difference between Dirac and Feynma’s work on path integral is that Dirac gave the probability amplitude by analogy of classical dynamics and he used the word, analogous to, while Feynman showed that the probability amplitude in quantum mechanics is equal to Eq. (1) Feynman et al. (2010).
The probability amplitude for photon, usually called photon’s Feynman propagator, has simple form in Feynman’s Gauge in momentum space Greiner and Reinhardt (2013)
$D_{F}^{\mu\nu}(k)=\frac{-g_{\mu\nu}}{k^{2}+i\epsilon},$ (3)
where $g^{\mu\nu}$ is metric tensor in four-dimensional Minkowski spacetime, $k$ is four-dimension wave vector, $(k_{0},k_{1},k_{2},k_{3})$, $k_{0}=\omega$, $(k_{1},k_{2},k_{3})$ are three components of usual wave vector, $k^{2}=k^{\mu}k_{\mu}=k_{0}^{2}-k_{1}^{2}-k_{2}^{2}-k_{3}^{2}$, $\epsilon$ is an infinitesimal quantity introduced to ensure convergence of the integral.
Photon’s Feynman propagator in coordinate space can be obtained via Fourier transform of Eq. (3)
$D_{F}^{\mu\nu}(x;y)=\int\frac{d^{4}k}{(2\pi)^{4}}e^{-ik\cdot(y-x)}D_{F}^{\mu\nu}(k),$ (4)
where $x$ and $y$ are four-dimensional Minkowski coordinate vector, $x=(x_{0},x_{1},x_{2},x_{3})$, $x_{0}$ is time $t$, $(x_{1},x_{2},x_{3})$ is the three components of usual coordinate vector $\vec{x}$. Equation (4) can be further simplified as Greiner and Reinhardt (2013)
$D_{F}(x;y)=\int\frac{d^{3}k}{(2\pi)^{3}}\frac{i}{(2\omega_{k})^{2}}e^{-i[\omega(t_{y}-t_{x})-\vec{k}\cdot(\vec{y}-\vec{x})]},$ (5)
in which $t_{y}>t_{x}$ is assumed.
Comparing photon’s Feynman propagator Eq. (5) with Green function in classical optics Born and Wolf (2013)
[…]
these two functions have similar expression with different contribution for each momentum component. The reason is that all the paths are taken into account in Eq. (5) with equal amplitude, while all the paths are taken into account in Eq. (6) with unequal amplitudes.
It has been proved by Feynman that if the Lagrangian of a particle has the form of Feynman et al. (2010)
$L(\dot{x},x,t)$ (7)
$=$ $a(t)\dot{x}^{2}+b(t)\dot{x}x+c(t)x^{2}+d(t)\dot{x}+e(t)x+f(t),$
where one dimension $x$ is assumed for simplicity. The particle’s Feynman propagator can be expressed as
$K(x_{a},x_{b})=e^{iS_{cl}(x_{a},x_{b})/\hbar}F(t_{a},t_{b}),$ (8)
where $x_{a}$ and $x_{b}$ are the starting and ending positions of the particle, respectively, $S_{cl}(a,b)$ is the classical action of the particle between $x_{a}$ and $x_{b}$, $F(t_{a},t_{b})$ is a function of the starting and ending time. Equation (8) means that the probability amplitude of particle going from $x_{a}$ to $x_{b}$ can be related to the classical propagation function with an extra function only dependent on the starting and ending time. The Lagrangian of free photon has the form of Eq. (7) Srednicki (2007). Equation (8) should be valid for photons. Classical Green function can be used as photon’s Feynman propagator in certain cases.
II.2.2 Distinguishable and indistinguishable paths
Indistinguishable paths is a unique concept in quantum theory. All the paths in classical theory are distinguishable. One reason is that measurement are different in quantum and classical theories Wheeler and Zurek (2014). In classical theory, measurement precision can be arbitrarily high (at least in principle) without introducing unpredictable influence to the measured system. In quantum theory, measurement precision can not be arbitrarily high due to the measurement will introduce unpredictable influence on the measured system. Another reason is that there is no identical particles in classical theory. As Griffiths and Schroeter mentioned in their book that all electrons are utterly identical, in a way that no two classical objects can never be, identical particles is a unique concept in quantum theory Griffiths and Schroeter (2018).
Figure 2: Distinguishable and indistinguishable paths. S1 and S2 are two sources can emit quantum or classical particles. D is a single-particle detector.
Taking Fig. 2 as an example to illustrate distinguishable and indistinguishable paths. S1 and S2 are two sources that can emit either quantum or classical particles. D is a single-particle detector. If the detector D detects a particle. The particle comes from either S1 or S2. In classical theory, the particles emitted by these two sources are always distinguishable. Even if all the properties such as mass, volume, color and so on, are identical [^1] This assumption is impossible in classical theory. For instance, two particles having the same weight in classical theory is only valid in some precision. Since classical particles are infinitely divisible, one can always find higher precision in which these two particles having different weights., the momentum difference between these two cases can be employed to distinguish these two paths. Different paths are always distinguishable in classical theory. Things become different in quantum theory. Assuming the particles emitted by S1 and S2 are identical. The only difference between these two paths is the momentum difference. The measurement precision of momentum in quantum theory is limited by Heisenberg uncertainty principle Heisenberg (1949)
$\Delta p_{x}\Delta x\geq h,$ (9)
where $\Delta p_{x}$ and $\Delta x$ are the uncertainties of momentum and position of the measured particle, respectively. The uncertainty of position is usually determined by the detection mechanism of the employed detector. For instance, the position uncertainty of photon detection is usually at the size of atom absorbing the photons via photoelectric effect, which is usually at $10^{-10}$ m range for metal atoms Shackelford (2016). The paths with momentum difference less than $h/10^{-10}$ are indistinguishable. Different paths can be indistinguishable in quantum theory. Heisenberg uncertainty principle should be employed to analyze the path difference when detail measuring apparatus is taken into account.
II.3 What is photon?
The concept of photon can trace back to Planck Plank (1901) and Einstein Einstein (1905). The answer to the question of what is photon is still not satisfactory. However, it does not stop photons appear everywhere even when there is no need for them Schleich (2003). As stated by Einstein in 1951: All these fifty years of conscious brooding have brought me no nearer to the answer to the question, -What are light quanta? Nowadays every Tom, Dick and Harry thinks he knows it, but he is mistaken Afshar et al. (2007). Despite significant progress in our understanding, it remains an open question Roychoudhuri and Roy (2003). In quantum optical coherence theory based on Feynman’s path integral, the concept of photon is unavoidable. It is impossible for us to answer the question of what is photon right now. However, it is helpful to clarify what we are talking about when the concept photon is employed.
We think photon is a particle, which is described by photon wave function $\psi(\vec{r},t)$. The probability to find a photon at $(\vec{r},t)$ in a volume $dV$ is given by $|\psi(\vec{r},t)|^{2}dV$. The initial phase of a photon is determined by the emission time.
II.3.1 Photon is a particle
Wave-particle duality is well-accepted for photon. However, when we talk about photon in this paper, we try to picture it as a particle like electron, proton and so on. The particle is described by a photon wave function $\psi(\vec{r},t)$, which is different from the concept of particle in classical theory.
The evidences for that photon is a particle are photoelectric effect Hertz (1887) and photon antibunching effect Paul (1982). Hertz observed the following phenomena when shining light on the surface of a metal:
(I) The kinetic energy $E$ of the ejected electron is equal to $\hbar\nu-\Phi$, where $\nu$ is the frequency of incidental light, and $\Phi$ is the work function of the metal.
(II) The rate of electron ejection is proportional to the square of the incident electric filed.
(III) The time delay $T$ between the light shining on the metal and the ejected photoelectron is very short.
It is experimentally proved by Lawrence and Beams that $T$ is shorter than 3 ns Lawrence and Beams (1928). Forrester et al. further proved that $T$ is much shorter than 0.1 ns Forrester et al. (1955). The first two phenomena can be interpreted by wave theory. However, the last phenomenon can only be interpreted by treating photon is a particle Muthukrishnan et al. (2017).
The second evidence for that photon is a particle is photon antibuching effect Paul (1982). The physics of photon antibunching is that a photon can never trigger two separated single-photon detectors simultaneously. The key to observe photon antibunching is to generate single-photon state. Kimble et al. Kimble et al. (1977) and Grangier et al. Grangier et al. (1986) observed photon antibunching effect with different ways to generate single-photon state. Classical electromagnetic wave theory can not interpret photon antibunching effect Paul (1982); Klyshko (1996).
There are also other evidence for photon is a particle. For instance, when a beam of light with low intensity is incident to a interferometer, the first-order interference pattern is accumulated point by point on the observation plane Dimitrova and Weis (2008).
II.3.2 Photon’s wave function
Strictly speaking, there is no wave function since for photon there is no position operator for photon Bohm (2012). However, one can define photon’s position operator from the point view of detection. Photon is usually detected by photoelectric effect. The position resolution is limited by the size of atom. As long as the position accuracy of photon’s wave function is not arbitrary high, one can define and use photon’s wave function by analogy of the wave function of massive particle.
The wave function of a single photon emitted by a two-level atom is Scully and Zubairy (1997)
$\psi(\vec{r},t)=K\frac{\sin\eta}{r}\theta(t-\frac{r}{c})\exp{[-i(\omega+i\frac{\Gamma}{2})(t-\frac{r}{c})]},$ (10)
where $K$ is a normalization constant, $r=|\vec{r}-\vec{r}_{0}|$ is the distance between the position of the atom $\vec{r}_{0}$ and the position of photon $\vec{r}$, $\eta$ is the angle of $\vec{r}$ comparing to the electric dipole of atom, $\theta(t-\frac{r}{c})$ is a step function, which means that photon can only be detected when photon propagates at position $\vec{r}$, $\omega$ is the frequency of photon, $\Gamma$ is the width of up-level energy of two-level system, $\exp{[-i(\omega+i\frac{\Gamma}{2})(t-\frac{r}{c})]}$ means that the probability to detect the photon is decayed by $e^{-\Gamma t}$.
Equation (10) only gives the temporal part of photon’s wave function. The spatial part of photon’s wave function can be treated the same as the spatial structure of classical electric field Scully and Zubairy (1997). For instance, the electric field emitted by a point source is spherical wave in classical theory. The spatial wave function of single photon emitted by a point source is also spherical.
II.3.3 Photon’s phase
Photon’s phase is defined by analogy of the phase of electric field in classical electromagnetic theory Dirac (1930). Classical electromagnetic field can be expressed as Born and Wolf (2013)
$\vec{E}(\vec{r},t)=\vec{E}_{0}\exp{[-i(\omega t-\vec{k}\cdot\vec{r})+i\varphi_{0}]},$ (11)
where $E_{0}$ is the amplitude of the field, $\omega t$ is the phase related to field propagating in time, $\vec{k}\cdot\vec{r}$ is the phase related to the field propagating in space, $\varphi_{0}$ is the initial phase of the field, $\varphi=-\omega t+\vec{k}\cdot\vec{r}+\varphi_{0}$ is the phase of electric field in classical theory.
Dirac gave a definition of quantum phase Dirac (1927). However, the phase operator defined by Dirac encounters difficulties when it is expanded in terms of number states. Several other definitions of quantum phase were also given Susskind and Glogower (1964); Barnett and Pegg (1986); Pegg and Barnett (1997). The usually employed phase operator is defined by phase shifts Pegg and Barnett (1997); Loudon (2000)
$|\varphi_{m}\rangle=\frac{1}{\sqrt{1+r}}\sum_{n=0}^{r}e^{in\varphi_{m}}|n\rangle,$ (12)
where $\varphi_{m}=2\pi m/(r+1)$ is a discrete phase to simplify the calculation, $m=0,1,2...r$, and $r$ is a very large positive integer. When $r$ goes to infinity, the discrete phase becomes continuous.
The uncertainties of phase ($\Delta\varphi$) and number of photons ($\Delta N$) satisfies the following uncertainty principle Pegg and Barnett (1997)
$\Delta N\Delta\varphi\geq\frac{1}{2}.$ (13)
For number state $|N\rangle$, the uncertainty of photon number is zero, which means that the phase of photon in number state is complete uncertain. It can be proved that the phase distribution function,
$P(\varphi)=\frac{1}{2\pi}|\sum_{0}^{\infty}c_{n}\exp{(-in\varphi)}|^{2},$ (14)
of single-photon state equals $1/2\pi$, which means the phase of photon in single-photon state is equally distributed between 0 and $2\pi$ Pegg and Barnett (1997).
Feynman assumed that the initial phase of photon is determined by the emission time Feynman (2006). Based on the phase relationship between different photons, light can be categorized into two groups. The first group is that the initial phases of photons in a light beam are random. The photons in this type of light are usually emitted by spontaneous emission. Chaotic light, thermal light, number state are examples of this type of light. The second group is that the initial phases of photons in a light beam are identical. The photons in this type of light are usually emitted by stimulated emission. Single-mode laser, amplified spontaneous emission belongs to the second type.
II.4 Coherence volume and degeneracy factor of light
The key to apply superposition principle in Feynman’s path integral to calculate optical coherence is distinguishing different paths for an event are distinguishable or not. Heisenberg’s uncertainty principle should be employed to judge whether different paths are distinguishable or not. Indistinguishable path is close related to indistinguishable photons. Photons within the same coherence volume are indistinguishable Martienssen and Spiller (1964). The uncertainties of coordinates ($\Delta x$, $\Delta y$, and $\Delta z$) and momentum ($\Delta p_{x}$, $\Delta p_{y}$, and $\Delta p_{z}$) of photons within the same coherence volume satisfies the following equation,
$\Delta x\Delta p_{x}=h$
$\Delta y\Delta p_{y}=h$ (15)
$\Delta z\Delta p_{z}=h.$
Assuming the light beam propagating in $z$ direction and the frequency bandwidth of light is $\Delta v$, it can be proved that the coherence volume equals the transverse coherence area times the longitudinal coherence length, $c/\Delta v$ Martienssen and Spiller (1964).
The average number of photons within one coherence volume is called degeneracy factor of the light. For blackbody radiation, the degeneracy factor is given by
$\delta=\frac{1}{e^{h\nu/k_{B}T}-1},$ (16)
where $h$ is Planck constant, $v$ is the frequency of radiation, $k_{B}$ is Boltzmann constant, $T$ is temperature of blackbody. The degeneracy factor of usual thermal light is much less than 1. For instance, the degeneracy factor of sun light at 532 nm equals 0.0073. On the other hand, the degeneracy factor of laser light can be much larger than 1. For instance, the degeneracy factor a 1-mW single-mode continuous-wave He-Ne laser light equals $3\times 10^{9}$ Mandel and Wolf (1995).
III First-order interference of light
The first-order interference of light can be well interpreted in classical optical coherence theory based on Maxwell’s electromagnetic theory Born and Wolf (2013). Dirac interpreted the first-order interference of light in the language of photon and gave the well-known statement about the first-order interference of light: “Each photon then interferes only with itself. Interferences between two different photons never occurs Dirac (1930).” The contents in this section are mainly based on Dirac’s statement. For a comprehensive review about the first-order interference of light, please refer to Refs. Born and Wolf (2013); Hecht (2012); Hariharan (2003).
III.1 First-order interference of one light beam
The first-order interference of one light beam means that light emitted by one source is split into two or more components and then combined together to observe interference. It corresponds to that each photon interferes only with itself in Dirac’s statement. Taking Mach-Zehnder interferometer shown in Fig. 3 as an example to show how to employ Feynman’s path integral to discuss the first-order interference of one light beam.
Figure 3: Mach-Zehnder interferometer. S is light source. BS1 and BS2 are two non-polarizing 1:1 beam splitter. M1 and M2 are two mirrors. D is a single-photon detector.
The event for the single-photon detector D detecting a photon in Fig. 3 has two different paths. The first one is the detected photon taking the upper path (transmitting through beam splitter BS1, reflected by mirror M1, and reflected by beam splitter BS2) and recorded as path 1 for simplicity. The probability amplitude for photon taking path 1 is written as $A_{1}(S,D)$. The second path is the photon taking the lower path (reflected by BS1 and M2, transmitting through BS2) and recorded as path 2. The probability amplitude for photon taking path 2 is written as $A_{2}(S,D)$. If these two different paths to trigger a photon detection event are indistinguishable, the probability function for the $j$th detected photon is
$P_{j}(t)$ $=$ $|A_{1}(S,D)+A_{2}(S,D)|^{2}$ (17)
$=$ $|e^{i\varphi_{j}}K_{1}(S,D)+e^{i\varphi_{j}}K_{2}(S,D)|^{2},$
where $j$ is a positive integer, $\varphi_{j}$ is the initial phase of the $j$th detected photon, $K_{1}(S,D)$ is photon’s Feynman propagator from the source to detector via path 1, $A_{1}(S,D)=e^{i\varphi_{j}}K_{1}(S,D)$, $\varphi_{j}$ is the initial phase of the $j$th detected photon. Similar definition holds for $K_{2}(S,D)$. The first-order interference of one light beam is due to photon interferes with itself. The initial phases are the same for two different probability amplitudes.
The first-order interference pattern is proportional to the sum of the probability functions of all the detected photon,
$I(t)\propto\sum_{j=1}^{N}P_{j}(t),$ (18)
where $N$ is the number of detected photon. For a large enough number of photons,
$I(t)\propto\langle P_{j}(t)\rangle,$ (19)
where delimited-⟨$\langle$..$\rangle$ … ⟩ is ensemble average by taking all possible realizations Mandel and Wolf (1995). Substituting Eq. (17) into Eq. (19), the first-order interference pattern in Fig. 3 is proportional to
$I(t)\propto\langle|K_{1}(S,D)+K_{2}(S,D)|^{2}\rangle.$ (20)
The initial phase of photon, $\varphi_{j}$, does not contribute to the first-order interference pattern in Fig. 3. No matter whether thermal or laser light is employed, as long as the frequency bandwidth and spatial distribution are the same, the observed first-order interference pattern will be identical. The reason is that the first-order interference is a result of photon interferes with itself. The initial phases of different paths for the $j$th detected photon are identical.
For a point light source, photon’s Feynman propagator can be approximated as
$K(\vec{r}_{0},t_{0};\vec{r},t)\propto\frac{1}{|\vec{r}-\vec{r}_{0}|}e^{-i[\omega_{0}(t-t_{0})-\vec{k}\cdot(\vec{r}-\vec{r}_{0})]},$ (21)
where $(\vec{r}_{0},t_{0})$ is the space-time coordinate of the photon being emitted, $(\vec{r},t)$ is the space-time coordinate of the photon being detected, $\omega_{0}$ is the angular frequency, $\vec{k}$ is wave vector. For simplicity, single-frequency is assumed in the following calculations. Substituting Eq. (21) into Eq. (20) and considering the temporal part, the first-order interference pattern in Fig. 3 is
$I(t_{1}-t_{2})$ $\propto$ $P_{j}(t_{1}-t_{2})$ (22)
$\propto$ $1+\cos[\omega_{0}(t_{1}-t_{2})],$
where $t_{1}$ and $t_{2}$ are the travailing time of photon from S to D via path 1 and 2, respectively. The first-order interference pattern is proportional to the probability function of the $j$th detected photon in Fig. 3, which is the same as the first-order interference pattern obtained in classical optical coherence theory Born and Wolf (2013).
The first-order interference pattern of light with finite frequency bandwidth can be calculated in the same way. Assuming S is a point single-photon source, the photon wave function can be expressed as Bialynicki-Birula (1996); Davis et al. (2018)
[…]
where $f(\omega)$ is the frequency distribution function, $\hat{a}^{\dagger}(\omega)$ is the creation operator, $|0\rangle$ is vacuum state.
When the time difference between the photon taking two paths, $|t_{1}-t_{2}|$, is larger than the coherence time of the employed single-photon state, the two different paths to trigger a photon detection event at D are distinguishable. The probability function for the $j$th detected photon equals
$P_{j}(t)$
$=$ $|e^{i\varphi_{j}}K_{1}(S,D)|^{2}+|e^{i\varphi_{j}}K_{2}(S,D)|^{2}$
[…]
where $t$ is the time of photon being detected by detector, $t_{1}=t-t_{01}$, $t_{01}$ is the emitting time of photon detected by D via path 1, $t_{2}=t-t_{02}$ and $t_{02}$ is the emitting time of photon detected by D via path 2. The probability of the $j$th detected photon is the sum of the probabilities of the photon taking two different paths. There is no first-order interference.
When the time difference, $|t_{1}-t_{2}|$, is less than the coherence time of the employed single-photon state, these two different paths to trigger a photon detection event at D are indistinguishable. The probability function of the $j$th detected photon is
$P_{j}(t)$ $=$ $|e^{i\varphi_{j}}K_{1}(S,D)+e^{i\varphi_{j}}K_{2}(S,D)|^{2}$
[…]
Assume the frequency distribution function is Gaussian,
$f(\omega)=A\exp{[-\frac{(\omega-\omega_{0})^{2}}{2\sigma^{2}}]},$ (26)
where $A$ is a normalization constant, $\sigma^{2}$ is variance of Gaussian function, $\omega_{0}$ is the central frequency of photon. Equation (III.1) can be simplified as Shih (2020)
$P_{j}(t_{1}-t_{2})=1+\frac{e^{-\sigma^{2}(t_{1}^{2}+t_{2}^{2})/2}}{e^{-\sigma^{2}t_{1}^{2}}+e^{-\sigma^{2}t_{2}^{2}}}\cos{\omega_{0}(t_{1}-t_{2})}.$ (27)
The first-order interference pattern can not be observed with only one photon. The first-order interference pattern of single-photon state can be observed by preparing many identical single-photon states and put these single photon states into the interferometer one by one. By collecting large enough number of photons, the first-order interference pattern with single-photon source in Fig. 3 is
$I(t_{1}-t_{2})$ $\propto$ $\langle P_{j}(t_{1}-t_{2})\rangle$ (28)
$=$ $1+\frac{e^{-\sigma^{2}(t_{1}^{2}+t_{2}^{2})/2}}{e^{-\sigma^{2}t_{1}^{2}}+e^{-\sigma^{2}t_{2}^{2}}}\cos{\omega_{0}(t_{1}-t_{2})},$
which is the same the probability of the $j$th detected photon.
Figure 4: First-order interference pattern of single-photon states in Mach-Zehnder interferometer
Figure 4 shows the first-order interference pattern of single-photon state with Gaussian spectrum in a Mach-Zehnder interferometer. $\tau=t_{1}-t_{2}$ is the time difference between path 1 and 2, I is the intensity of light. When the time difference equals 0, the visibility of the first-order interference pattern can reach 100%. As the time difference becomes larger, the visibility of first-order interference pattern decreases. When the time difference is larger than the coherence length of light, the visibility goes to zero, which is consistent with Eq. (III.1). The simulated result in Fig. 4 is consistent with experimental results with single photons Braig et al. (2003)
Similar method can be employed to calculate the spatial first-order interference pattern of one light beam in Feynman’s path integral.
III.2 First-order interference of two independent light beams
The first-order interference of one light beam confirms Dirac’s statement that photon only interferes with itself. However, based on Popper’s philosophy that a scientific conclusion can not be proved to be true Popper (2005). The first-order interference of light does not prove Dirac’s statement is correct. On the contrary, any scientific conclusion should be possible to be disproved. Dirac’s statement can be disproved if one can observe the first-order interference pattern with two independent light beams.
When Young first discussed interference of light, he pointed out that there is first-order interference pattern of two light beams from different origins by analogy of sound and water waves Young (1802): When two Undulations, from different Origins, coincide either perfectly or very nearly in Direction, their joint effect is a Combination of the Motions belonging to each. Later, Young realized that it is impossible to observe first-order interference pattern of light emitted by two different origins. He observed the first-order interference pattern of sunlight by splitting one beam into two and then combining these two beams together Young (1804). It is now well-accepted that in order to observe the first-order interference pattern with two light beams, these two light beams must have the polarization and frequency, and fixed relative phase Hecht (2012). The easiest way to meet these condition is these two light beams are splitting from the same one. Most of the first-order interferometers are designed in this way Born and Wolf (2013).
Thing becomes different when laser was invented Maiman et al. (1960); Hecht (2010). Several groups superposed two light beams emitted by two independent lasers and observed the first-order interference pattern within the coherence time Javan et al. (1962); Magyar and Mandel (1963); Lipsett and Mandel (1963), which is usually called transient first-order interference pattern. Does the observed transient first-order interference pattern with two independent laser light beams mean that the second part of Dirac’s statement is disproved? Further experiments shows that the first-order interference pattern exists when the intensities of these two laser light beams are so low that there is only one photon in the interferometer in most time Pfleegor and Mandel (1968); Radloff (1968); Hariharan et al. (1993). It is thought by Pfleegor and Mandel that the observed transient first-order interference pattern does not conflict with Dirac’s statement Pfleegor and Mandel (1967). It was the detection of the photon makes the photon into a superposition state and give interference Pfleegor and Mandel (1967). What Glauber said about these experiments is easier to understand Glauber (1995): The things that interfere in quantum mechanics are not particles. They are probability amplitudes for certain events. It is the fact that probability amplitudes add up like a complex numbers that is responsible for all quantum mechanics interferences.
III.2.1 First-order interference of two independent laser light beams
In the first-order interference of two independent laser light beams shown in Fig. 5, there are two different paths for a single-photon detector D detecting a photon. The first one is that the detected photon comes from laser S1 and the corresponding probability amplitude is $A_{1}(x)$. The second one is that the detected photon comes from laser S2 and the corresponding probability amplitude is $A_{2}(x)$. If these two laser light beams have the same polarization, frequency bandwidth, intensity, these two different ways can only be distinguished by momentum difference. As introduced in Sect. II.2.2, when the distance between two lasers satisfies
$d\leq\frac{\lambda L}{\Delta x},$ (29)
these two different paths are indistinguishable, where $L$ is the distance between source and observation planes, $\lambda$ is the wavelength of photon, and $\Delta x$ is the position uncertainty of photon detection. The probability function for the $j$th detected photon is
$P_{j}(x)$ $=$ $|A_{j1}(x)+A_{j2}(x)|^{2}$ (30)
$=$ $|e^{i\varphi_{j1}}K_{j1}(x)+e^{i\varphi_{j2}}K_{j2}(x)|^{2},$
where $\varphi_{j1}$ and $\varphi_{j2}$ are the initial phases of the $j$th detected photon emitted by S1 and S2, respectively, $K_{j1}(x)$ and $K_{j2}(x)$ are the corresponding photon’s Feynman propagators.
Figure 5: The first-order interference of two independent light beams. S1 and S2 are two independent light sources. The distance between two sources is $d$. The distance between the source plane and the observation plane, O, equals $L$. D is a single-photon detector.
For a point light source, Eq. (30) can be simplified as
$P_{j}(x)\propto 1+\cos[\frac{2\pi d}{\lambda L}x+(\varphi_{j1}-\varphi_{j2})],$ (31)
where paraxial approximation and one-dimension are employed to simplify the calculation. The initial phases of photons in single-mode continuous-wave laser light are identical within the coherence volume. The initial phases of photons in different coherence volumes are random. The initial phases of two independent lasers, $\varphi_{j1}$ and $\varphi_{j2}$, are independent. However, the relative phase, $\varphi_{j1}-\varphi_{j2}$, is fixed within one coherence volume for two independent lasers. The observed intensity is proportional to the sum of the probabilities of all the detected photons
$I(x)$ $\propto$ $\sum_{j}^{N}P_{j}(x)$ (32)
$\propto$ $\langle 1+\cos[\frac{2\pi d}{\lambda L}x+(\varphi_{j1}-\varphi_{j2})]\rangle.$
When the observation time is shorter than the coherence time, $\varphi_{j1}-\varphi_{j2}$ is a random but fixed number. The observed interference pattern of two independent lasers is
$I(x)$ $\propto$ $1+\cos[\frac{2\pi d}{\lambda L}x+(\varphi_{j1}-\varphi_{j2})],$ (33)
where the first-order interference pattern can be observed.
When the observation time is much longer than the coherence time, the sum of random phases averages to a constant. The observed intensity
$I(x)$ $\propto$ $1,$ (34)
where the first-order interference pattern can not be observed for a long observation time.
The results above are consistent with Magyar and Mandel’s experiments Magyar and Mandel (1963). Even through the physics of the interference of two independent lasers is in question for a long time Mandel (1964); Louradour et al. (1993); Wallace (1994); Davis and Parigger (1994); Paul (1986), the experimental results are certain and the predictions in quantum optical coherence theory is consistent with the one obtained in classical theory.
III.2.2 First-order interference of two independent thermal light beams
Things become interesting for the first-order interference of two independent thermal light beams. As mentioned in Sect. II.3, there are two major differences between photons in laser and thermal light. The first one is that the initial phases of photons in thermal light are random within the coherence volume and the initial phases of photons in laser light are identical within one coherence volume. The second difference is that the degeneracy factor of thermal light is usually much less than 1 and the degeneracy factor of laser light can be much larger than one. The latter difference indicates that it may be impossible to detect large enough number of photons to see whether there is transient first-order interference pattern of two independent thermal light beams within the coherence time. No transient first-order interference pattern of two independent thermal light beams has been reported as the one for two independent lasers Magyar and Mandel (1963).
Theoretically, we can assume the degeneracy factor of thermal light is much larger than one and discuss whether there exists transient first-order interference pattern or not. The scheme for the first-order interference of two independent thermal light beams is also represented by Fig. 5, where S1 and S2 are two identical and independent thermal light sources. There are two different paths for the single-photon detector D detecting a photon. With the same process as the one in Sect. III.2.1, the probability function for the $j$th detected photon is
$P_{j}(x)\propto 1+\cos[\frac{2\pi d}{\lambda L}x+(\varphi_{j1}-\varphi_{j2})],$ (35)
where paraxial approximation and one-dimension are assumed. The probability function for $N$ detected photon is
$P_{N}(x)$ $=$ $\sum_{j}^{N}1+\cos[\frac{2\pi d}{\lambda L}x+(\varphi_{j1}-\varphi_{j2})].$ (36)
Since the initial phases of photons in thermal light are random, $\varphi_{j1}-\varphi_{j2}$ is random for each detected photon. For a large enough number of photons, Eq. (36) can be simplified as Bai et al. (2016)
$P_{N}(x)\propto 1+\frac{1}{\sqrt{N}}\cos(\frac{2\pi d}{\lambda L}x+\varphi),$ (37)
where $\varphi$ is the result of the sum of large random phases. The visibility of the first-order interference pattern after detecting $N$ photon is
$V_{N}=\frac{P_{Nmax}-P_{Nmin}}{P_{Nmax}+P_{Nmin}}=\frac{1}{\sqrt{N}},$ (38)
which decreases as the number of detected photons increases. Figure 6 shows the relationship between the simulated visibility of first-order interference pattern of two independent thermal light beams and the number of detected photons. $V$ is the visibility and $N$ is the number of detected photons. The red line is theoretical value given by Eq. (38) and square dots are numerical results. Theoretical and numerical results are consistent Bai et al. (2016).
Figure 6: The visibility of first-order interference pattern of two independent thermal light beams vs. the number of detected photons Bai et al. (2016).
The first-order interference pattern can not be observed with only one or several photons due to one photon only triggers a photo-electron Dimitrova and Weis (2008). There must be large enough number of photons to retrieve the first-order interference pattern. However, the visibility decreases as the number of detected photon increases via Eq. (38). The first-order interference pattern of two independent thermal light beams may not be observable, which is similar as the first-order interference pattern of two independent laser light beams can not be observed for a long observation time.
III.2.3 First-order interference of two independent non-classical light beams
Feynman’s path integral can also be employed to calculate the first-order interference of two independent non-classical light beams. Assume S1 and S2 in Fig. 5 are two identical and independent single-photon sources. There are two different paths for single-photon detector D detecting a photon, which are emitted by S1 and S2. The probability amplitudes for these two paths are $A_{1}(x)$ and $A_{2}(x)$, respectively. When the distance between these two sources is less than $\lambda L/\Delta x$, these two difference paths can not be distinguished by momentum difference. However, this condition alone does not guarantee that these two paths are indistinguishable as the one in laser and thermal light cases.
When there is only one photon in the interferometer, these two difference paths are always distinguishable. One can measure the status of the single-photon sources to ensure the detected photon comes from S1 or S2 without changing the detection results. The probability function for the $j$th detected photon is
$P_{j}(x)$ $=$ $|A_{j1}(x)|^{2}+|A_{j2}(x)|^{2}$ (39)
$=$ $|e^{i\varphi_{j1}}K_{j1}(x)|^{2}+|e^{i\varphi_{j2}}K_{j2}(x)|^{2}.$
The probability equals the sum of the probabilities for different paths. There is no interference.
When these two single-photon sources emit photons simultaneously, there are two photons in the interferometer and these two different paths are distinguishable when all the other parameters are indistinguishable. The probability function for the $j$th detected photon is
$P_{j}^{\prime}{}(x)=|e^{i\varphi_{j1}}K_{j1}(x)+e^{i\varphi_{j2}}K_{j2}(x)|^{2}.$ (40)
The probability function after detecting $N$ photons is the sum of the probabilities correspond to these two cases,
$P_{N}(x)=\sum_{j=1}^{N_{1}}P_{j}(x)+\sum_{j=1}^{N_{2}}P_{j}^{\prime}{}(x),$ (41)
where $N_{1}$ is the number of photons when only one source emitted photons, $N_{2}$ is the number of photons when two sources emitted photons simultaneously, $N_{1}+N_{2}=N$. With the same method as the one for the first-order interference of two independent thermal light beams, Eq. (41) can be simplified as
$P_{N}(x)\propto 1+\frac{N_{2}}{N}\frac{1}{\sqrt{N_{2}}}\cos(\frac{2\pi d}{\lambda L}x+\varphi).$ (42)
$N_{2}$ is much less than $N$ for two independent single-photon sources. The visibility of the first-order interference pattern of two independent single-photon states is much less than the one of two-independent thermal light beams. It is impossible to observe the first-order interference pattern of two independent single-photon states for either short or long time comparing to the coherence time.
With the same method above, one can calculate the first-order interference of two independent number states with different numbers, one thermal and one laser light beams, and other conditions. No transient first-order interference pattern can be observed except the one of two independent laser light beams. No first-order interference pattern can be observed for two independent light beams with long observing time.
III.2.4 Remakes about the classical model of laser and thermal light field
With the results above, it is interesting to re-visit the electric model for laser and thermal light in classical electromagnetic theory. For simplicity, quasi-monochromatic light and plane wave are assumed. The electric field for single-mode continuous-wave laser light at space-time coordinate $(\vec{r},t)$ is Loudon (2000); Born and Wolf (2013)
$\vec{E}(\vec{r},t)=\vec{E}_{j}e^{-i(\omega_{0}t-\vec{k}\cdot\vec{r}-\varphi_{j})},$ (43)
where $\vec{E}_{j}$ is the amplitude, $\omega_{0}$ is the central angular frequency, $\vec{k}$ is the wave vector, $\varphi_{j}$ is the initial phase of the field. $\vec{E}_{j}$ and $\varphi_{j}$ are constant within one coherence time and change randomly between different coherence times.
The same model as Eq. (43) is employed to describe thermal light field within one coherence time. For instance, in Born and Wolf’s book, they wrote: In a monochromatic wave field the amplitude of the vibrations at any point P is constant, while the phase varies linearly with time. This is no longer the case in a wave field produced by a real source: the amplitude and phase undergo irregular fluctuation, the rapidity of which depends essentially on the effective width $\Delta\nu$ of the spectrum. The complex amplitude remains substantially constant only during a time interval $\Delta t$ which is small compared to the reciprocal of the effective spectral width $\Delta\nu$; in such a time interval the change of the relative phase of any two Fourier components is much less than $2\pi$ and the addition of such components represents a disturbance which in this time interval behaves like a monochromatic wave with the mean frequency. However, this is not true for a longer time interval. The characteristic time $\Delta t=1/\Delta v$ is of the order of the coherence time introduced in §7.5.8. They thought the amplitude and phase of electric field produced by a real source undergoes irregular fluctuations for a long time interval. The amplitude and phase of electric field can be seen as constant within the coherence time. In other words, the electric field produced by real source, i.e., laser or thermal light sources, can be described by Eq. (43). The same electric field model is employed to describe laser and thermal light field within the coherence time. This model is employed by Forrester to calculate the first-order interference of two independent thermal light beams, too Forrester (1956).
In classical theory, it is straightforward to obtain the conclusion that there is transient first-order interference pattern by superposing two independent thermal light beams as the one of two independent laser light beams. This conclusion is in conflict with the conclusion obtained in Feynman’s path integral in Sect. III.2.2. Experiments should be employed to verify the different conclusions. However, due to the very low degeneracy factor of thermal light, it is very difficult to observe transient first-order interference pattern of two independent thermal light beams directly.
Forrester et al. claimed they observed interference between independently generated light waves Forrester et al. (1955). Detail analysis of their experiments indicates that their experiments is a type of one beam interference including nonlinear effect as the slit in a Young’s double-slit interferometer, which is similar as the ones with single atom Grangier et al. (1985); Eichmann et al. (1993) and atom ensembles Afzelius et al. (2007) as the slit in a Young’s double-slit interferometer. Forrester et al.’s experiments can not be interpreted as the first-order interference of two independent thermal light beams.
If there is no transient first-order interference pattern with two independent thermal light beams, the classical model for thermal light within the coherence time may not be the same as the one of laser light. This result is helpful for us to understand the physics of thermal light and worthy of further experimental studies.
III.3 First-order interference of multiple independent light beams
The same method can be employed to calculate the first-order interference of multiple independent light beams. Figure 7 shows the first-order interference of three light beams emitted by three independent light sources, S1, S2, and S3.
Figure 7: First-order interference of three independent light beams. S1, S2, and S3 are three independent light sources. D is a single-photon detector in the observation plane. $L$ is the distance between the source and observation planes.
There are three different paths for the single-photon detector D detecting a photon. The first one is that the detected photon is emitted by S1 and the corresponding probability amplitude is $A_{1}(S,D))$. The second and third paths can be defined similarly and the corresponding probability amplitudes are $A_{2}(S,D)$ and $A_{3}(S,D)$, respectively. For simplicity, these three light sources are assumed to be identical single-mode continuous-wave lasers with equal intensities. The polarizations and frequency distributions of the photons emitted by these three lasers are assumed to be the same. When the distance between $S_{1}$ and S3 is less than $\lambda L/\Delta x$, these three different paths to trigger a photon detection event are indistinguishable. The probability function for the $j$th detected photon is
$P_{j}(x)$
$=$ $|A_{1}(S,D)+A_{2}(S,D)+A_{3}(S,D)|^{2}$
$=$ $|e^{i\varphi_{j1}}K_{1}(S,D)+e^{i\varphi_{j2}}K_{2}(S,D)+e^{i\varphi_{j3}}K_{3}(S,D)|^{2},$
where $\varphi_{j1}$, $\varphi_{j2}$ and $\varphi_{j3}$ are the initial phases of the $j$th detected photon emitted by S1, S2 and S3, respectively. $K_{1}(S,D)$, $K_{2}(S,D)$ and $K_{3}(S,D)$ are the corresponding photon’s Feyman propagators for these three paths.
Equation (7) can be simplified as
$P_{j}(x)$ $=$ $\frac{3}{2}+\cos[\frac{2\pi d_{12}}{\lambda L}(x-\frac{d_{23}}{2})+(\varphi_{1j}-\varphi_{2j})]$ (45)
$+\cos[\frac{2\pi d_{23}}{\lambda L}(x+\frac{d_{12}}{2})+(\varphi_{2j}-\varphi_{3j})]$
$+\cos[\frac{2\pi d_{13}}{\lambda L}x+(\varphi_{1j}-\varphi_{3j})],$
where paraxial approximation and one-dimension are assumed to simplify the calculation, $d_{12}$ is the distance between S1 and S2, similar definitions hold for $d_{13}$ and $d_{23}$. The relative position for these three sources is shown in Fig. 8.
Figure 8: Relative positions of three sources.
Since these three light sources are independent lasers, $\varphi_{j1}$, $\varphi_{j2}$ and $\varphi_{j3}$ are random. The relative phases are fixed if all the detected photons are within one coherence volume. The probability for $N$ detected photon is
$P_{N}(x)\propto P_{j}(x),$ (46)
where transient first-order interference pattern between three independent laser light beams can be observed. Figure 9 shows transient first-order interference pattern of three laser light beams with equal initial phases. When $d_{12}\neq d_{23}$, the visibility of the transient first-order interference pattern can not reach 100%. The visibility can reach 100% when $d_{12}=d_{23}$ as shown in Fig. 9(b).
Figure 9: Transient first-order interference pattern for (a) $d_{12}\neq d_{23}$ and (b) $d_{12}=d_{23}$.
When the observation time is much longer than the coherence time, the probability for $N$ detected photons can be approximated as a constant due to the random initial phases.
The same method can be employed to calculate the first-order interference of three and more different types of independent light beams. It can be concluded that only when more than two laser light beams are employed, transient first-order interference pattern can be observed.
IV Second-order interference of light
As mentioned above, nearly all the studies of optical coherence are about the first-order interference of light before 1956 Born and Wolf (2013). In 1956, Hanbury Brow and Twiss discovered two-photon bunching of thermal light Brown and Twiss (1956b, a), which is one of the most famous second-order interference phenomena in quantum optics. The history of second-order coherence of light can trace back to 1940s when several groups measured the correlation of photon pair generated in the annihilation of a positron-electron pair Hanna (1948); Bleuler and Bradt (1948); Wu and Shaknov (1950). The second-order interference of light is measured with coincidence count detection system, which is proposed by Bothe in 1925 Bothe and Geiger (1925).
The second-order interference of light has an unique role in the development of optical coherence theory. The first-order interference of light exists in nature. For instance, soap bubbles, transparent wings of insects, some butterfly wings, etc. appear colored streaks when exposed to sunlight is due to the first-order interference of light. The second-order interference of light does not exist in nature. It is a human invented phenomenon. Special detection scheme is required to observe the second-order interference of light. The second-order coherence is qualitatively different from the first-order interference. It may be reason why people took more than 200 years since the first observed phenomenon of first-order interference of light to the first observed phenomenon of second-order interference of light. Extending the second-oder coherence to the third- and higher-order coherence is natural. The third- and higher-order coherence is quantitative different from the second-order coherence.
Figure 10: (a) First-, (b) second- and (c) third-order interference. S is light source. IT1, IT2, IT3 are first-, second- and third-order interferometers, respectively. D is photon detector. CC is coincidence count detection scheme.
Figure 10 shows the schemes for the first-, second- and third-order interference of light. For comparison, the first-order interference of light shown in Fig. 10(a) is illustrated first. In a typical experiment of first-order interference of light, light generated by the light source S is split into more than one components in the first-order interferometer IT1. Then the split light beams are recombined to observe the intensity distribution over space or time by a single photo-detector. Young’s double-slit interferometer, Mach-Zehnder interferometer, Michelson interferometer are the typical first-order interferometers Born and Wolf (2013). The first-order interference of two and multiple independent light beams introduced in Sects. III.2 and III.3 belong to the first-order interference of light.
A typical scheme for the second-order interference of light is shown in Fig. 10(b). Light emitted by source S is split into two or more different components in the second-order interferometer (IT2). Then the output light beams of IT2 are detected by two detectors (D1 and D2). Hanbury Brown-Twiss (HBT) interferometerBrown and Twiss (1956b, a), Hong-Ou-Mandel (HOM) interferometer Hong et al. (1987); Shih and Alley (1988); Alley and Shih (1986), Franson interferometer Franson (1989) are some typical second-order interferometers. The output signals of these two detectors are sent into coincidence count detection system (CC) to measure the correlation of these two signals. Based on what types of photo-detector is employed, there are two work principles for CC. The first one is for normal photo-electric detector. The output of the detector is electric current or voltage proportional to the intensity of the incidental light. The measured correlation is
$\Gamma^{(2)}(\vec{r}_{1},t_{1};\vec{r}_{2},t_{2})=\langle I_{1}(\vec{r}_{1},t_{1})I_{2}(\vec{r}_{2},t_{2})\rangle,$ (47)
where $I_{1}(\vec{r}_{1},t_{1})$ is the intensity of incidental light at space-time coordinate $(\vec{r}_{1},t_{1})$ detected by D1, $I_{2}(\vec{r}_{2},t_{2})$ is the intensity of incidental light at space-time coordinate $(\vec{r}_{2},t_{2})$ detected by D2, delimited-⟨$\langle$..$\rangle$ … ⟩ is ensemble average. For ergodic and stationary system, ensemble average can be approximated by time average Mandel and Wolf (1995). Equation (47) measures the correlation of intensity fluctuations detected by these two detectors. The CC system based on normal photo-electric detector was employed by Hanbury Brown and Twiss in their correlation experiments Brown and Twiss (1956b, a).
The second one is for single-photon detector. The output of single-photon detector is a voltage pulse. Equation (47) can not be employed to calculate the correlation of the outputs from single-photon detectors. A different coincidence count principle should be employed. When the time difference between these two photon detection events is less than a time window $T$, these two photon detection events is a two-photon coincidence count. Otherwise, these two photon detection is not a two-photon coincidence count. The time window $T$ should be less than the coherence time due to these two photons are indistinguishable within the same coherence volume. There are also other coincidence count detection system, such as two-photon absorption Boitier et al. (2009), sum frequency generation Pe’er et al. (2007), which are also based on the same principle as the one for single-photon detector.
The second-order interference shown in Fig. 10(b) can be extended to the third- and higher-order interference. Figure 10(c) shows the scheme for the third-order interference of light. The light beam emitted by source S is split into multiple components and detected by three detectors, D1, D2 and D3. The output of these three detectors are sent into a three-photon coincidence count system (CC) to measure the third-order correlation. Based on different types of detectors are employed, there are also two different types of principle for three-photon coincidence count system. When normal photo-detectors are employed, the third-order correlation is given by
$\Gamma^{(3)}(\vec{r}_{1},t_{1};\vec{r}_{2},t_{2};\vec{r}_{3},t_{3})=\langle I_{1}(\vec{r}_{1},t_{1})I_{2}(\vec{r}_{2},t_{2})I_{3}(\vec{r}_{3},t_{3})\rangle,$ (48)
where the meanings of the symbols in Eq. (48) are similar as the ones in Eq. (47). When single-photon detectors are employed, the time differences between these three photon detection events are less than a time window $T$, the three photon detection events is a three-photon coincidence count. Otherwise, the three photon detection events is not a three-photon coincidence count. Similar method can be employed to define the fourth- and higher-order interference of light.
IV.1 Second-order interference of one light beam
Figure 11 shows the second-order interference one light beam in a HBT interferometer. The light beam emitted by source S is split into two by a 1:1 non-polarizing beam splitter BS. The output light beams of BS are detected by two single-photon detectors, D1 and D2. The outputs of these two detectors are sent into two-photon coincidence count system CC to measure the second-order correlation of light.
Figure 11: One light beam in a HBT interferometer. S is light source. BS is 1:1 nonpolarizing beam splitter. D1 and D2 are two single-photon detectors. CC is two-photon coincidence count system.
IV.1.1 Second-order interference of one thermal light beam
Assuming S is a thermal light source in Fig. 11, the scheme shown in Fig. 11 is equivalent to the scheme employed by Hanbury Brown and Twiss discovering two-photon bunching of thermal light in 1956 Brown and Twiss (1956b, a). There are two different paths for D1 and D2 detecting a two-photon coincidence count in Fig. 11. The first path is photon a is detected by D1 and photon b is detected by D2. The probability for the path is $1/2$ and the corresponding two-photon probability amplitude is $A(a,D_{1};b,D_{2})$. The second path is photon a is detected by D2 and photon b is detected by D1. The probability for the path is $1/2$ and the corresponding two-photon probability amplitude is $A(a,D_{2};b,D_{1})$. When photons a and b are in the same coherence volume of thermal light, these two different paths to trigger a two-photon coincidence count are indistinguishable. The two-photon probability function for the $j$th detected photon pair is Feynman et al. (2010, 2011)
$P^{(2)}_{j}(\vec{r}_{1},t_{1};\vec{r}_{2},t_{2})$ (49)
$=$ $|A_{j}(a,D_{1};b,D_{2})+A_{j}(a,D_{2};b,D_{1})|^{2}$
$=$ $|e^{i(\varphi_{aj}+\varphi_{bj})}K_{j}(a,D_{1};b,D_{2})$
$+e^{i(\varphi_{aj}+\varphi_{bj})}K_{j}(a,D_{2};b,D_{1})|^{2}$
$=$ $|K_{j}(a,D_{1};b,D_{2})+K_{j}(a,D_{2};b,D_{1})|^{2},$
where $\varphi_{aj}$ and $\varphi_{bj}$ are the initial phases of photons a and b in the $j$th detected photon pair, respectively, $K_{j}(a,D_{1};b,D_{2})$ is photon’s Feynman propagator for photon a emitted by S going to D1 and photon b emitted by S going to D2. The meaning of $K_{j}(a,D_{2};b,D_{1})$ is defined similarly. Based on the principle for the probability amplitude of independent events introduced by Feynman, $K_{j}(a,D_{1};b,D_{2})=K_{j}(a,D_{1})K_{j}(b,D_{2})$. Equation (49) can be simplified as
$P^{(2)}_{j}(\vec{r}_{1},t_{1};\vec{r}_{2},t_{2})$ (50)
$=$ $|K_{j}(a,D_{1})K_{j}(b,D_{2})+K_{j}(a,D_{2})K_{j}(b,D_{1})|^{2}.$
The two-photon probability function of $N$ detected photon pairs is the sum of the probabilities of $N$ detected photon pairs,
$P^{(2)}(\vec{r}_{1},t_{1};\vec{r}_{2},t_{2})$
$=$ $\sum_{j=1}^{N}P^{(2)}_{j}(\vec{r}_{1},t_{1};\vec{r}_{2},t_{2})$
$\propto$ $\langle P^{(2)}_{j}(\vec{r}_{1},t_{1};\vec{r}_{2},t_{2})\rangle$
$=$ $\langle|K_{j}(a,D_{1})K_{j}(b,D_{2})+K_{j}(a,D_{2})K_{j}(b,D_{1})|^{2}\rangle,$
where delimited-⟨$\langle$..$\rangle$ … ⟩ is ensemble average. The normalized two-photon probability function,
$p^{(2)}(\vec{r}_{1},t_{1};\vec{r}_{2},t_{2})=\frac{P^{(2)}(\vec{r}_{1},t_{1};\vec{r}_{2},t_{2})}{P^{(1)}(\vec{r}_{1},t_{1})P^{(1)}(\vec{r}_{2},t_{2})},$ (52)
is employed to eliminate the influence of the first-order interference on the second-order interference, where $P^{(1)}(\vec{r}_{1},t_{1})$ and $P^{(1)}(\vec{r}_{2},t_{2})$ are single-photon probability function at space-time coordinates $(\vec{r}_{1},t_{1})$ and $(\vec{r}_{2},t_{2})$, respectively. $P^{(1)}(\vec{r}_{1},t_{1})$ is written as $P(\vec{r}_{1},t_{1})$ if no misunderstanding exists. Two-photon probability function $P^{(2)}(\vec{r}_{1},t_{1};\vec{r}_{2},t_{2})$ and normalized two-photon probability function $p^{(2)}(\vec{r}_{1},t_{1};\vec{r}_{2},t_{2})$ correspond to the second-order coherence function $G^{(2)}(\vec{r}_{1},t_{1};\vec{r}_{2},t_{2})$ and normalized second-order coherence function $g^{(2)}(\vec{r}_{1},t_{1};\vec{r}_{2},t_{2})$ in Glauber’s quantum optical coherence theory, respectively Glauber (1963a, c).
Let us first calculate temporal two-photon bunching of thermal light. Assuming S is a point light source, D1 and D2 are at symmetrical positions of BS to simplify the calculation. In this case, photon’s Feynman propagator is
[…]
where the spatial part is ignored, $\omega_{0}$ is the central angular frequency, $\Delta\omega$ is angular frequency distribution function of thermal light, $t_{1}$ is the time of photon detection event at D1, $t_{a}$ is the emitting time of photon a. Substituting Eq. (53) into Eq. (IV.1.1) and considering $f(\omega)$ equals 1 within $[\omega_{0}-\Delta\omega/2,\omega_{0}+\Delta\omega/2]$ and equals 0 in other domains, the temporal two-photon probability function is Shih (2020)
$P^{(2)}(t_{1}-t_{2})\propto 1+\text{sinc}^{2}\frac{\Delta\omega(t_{1}-t_{2})}{2}.$ (54)
It is easy to prove that $P^{(1)}(t_{1})$ and $P^{(2)}(t_{1})$ are constant for thermal light. The normalized temporal two-photon probability function, $p^{(2)}(t_{1}-t_{2})$ is proportional to $P^{(2)}(t_{1}-t_{2})$. In the following, if the normalized probability function is proportional to the corresponding probability function, only probability function is calculated.
Figure 12: Temporal two-photon bunching of thermal light. $t_{1}-t_{2}$ is the relative position difference between D1 and D2. Lines 1, 2, 3 correspond to the frequency bandwidths equal $\Delta\omega$, $2\Delta\omega$, and $4\Delta\omega$, respectively.
Figure 12 shows the temporal two-photon probability functions of thermal light with different frequency bandwidths. When two detectors are at symmetrical positions, the probability of detecting two photons equal 2. As the time difference, $|t_{1}-t_{2}|$, increases, the probability of detecting two photons decreases. When the time difference between two photon detection events is much longer than the coherence time, the probability of detecting two photons equal 1, which means that these two photon detection events are independent.
The spatial two-photon bunching of thermal light can be calculated in the similar way. Assuming one dimension light source is employed, the light emitted by S is single frequency light, and the distance between the source plane and D1 is equal the one of source plane and D2. photon’s Feynman propagator in this case can be approximated by Green function in classical optics Goodman (2005)
[…]
where $d$ is the length of thermal light source, $s(x_{s})$ is the distribution function of S, $\vec{r}_{1}$ is position vector of D1, $\vec{r}_{s}$ is the position vector of S emitting photon a, $\vec{k}_{s1}$ is wave vector of light. Substituting Eq. (55) into Eq. (IV.1.1) and considering $s(x_{s})$ equals 1 within $[-d/2,d/2]$ and equals 0 in other domains, spatial two-photon probability function of thermal light in a HBT interferometer is Shih (2020)
$P^{(2)}(x_{1}-x_{2})\propto 1+\text{sinc}^{2}\frac{d}{L\lambda}(x_{1}-x_{2}).$ (56)
Based on the results above, two-photon bunching of thermal light can be interpreted as the result of two-photon interference, which is the superposition of two-photon probability amplitudes for two different and indistinguishable paths to trigger a two-photon coincidence count Fano (1961); Feynman (2006); Glauber (1995). Two-photon interference in the second-order interference of light can not be used to disprove Dirac’s statement that different photons never interferes with each other in the first-order interference of light Paul (1986). By analogy, Dirac’s statement about the first-order interference of light can be extended to the second-order interference of light: Each photon pair only interferes with itself and interferes between two different photon pairs never occurs Shih (2020).
IV.1.2 Second-order interference of one superbunching light beam
$g^{(2)}(0)$ equals 2 for thermal light. $g^{(2)}(0)$ measured in the HBT interferometer is the degree of second-order coherence of light, which is usually employed to categorize different types of light Loudon (2000). $g^{(2)}(0)$ equals 1 for single-mode continuous-wave laser light. When $g^{(2)}(0)$ is less than 1, the light belongs to non-classical light. When $g^{(2)}(0)$ is larger than 2, it is superbunching light Ficek and Swain (2005). Superbunching light can be either classical or nonclassical. For instance, entangled photon pairs is non-classical superbunching light Klyshko (1996) and superbunching pseudothermal light is classical light Zhou et al. (2017). This section will concentrate on the second-order interference of a classical superbunching light beam in HBT interferometer.
There are two different methods to increase the degree of second-order coherence of classical light. The first one is adding more two-photon paths. The second one is the two-photon probability amplitudes changes with time.
Figure 13: Increase the degree of second-order coherence by adding more two-photon paths.
Figure 13 shows how to increase the degree of second-order coherence by adding more two-photon paths. Figure 13(a) shows there are two different ways for photon a and b to trigger a two-photon coincidence count in D1 and D2. The first one is photon a is detected by D1 and photon b is detected by D2. The corresponding two-photon probability amplitude is $A(a,D_{1};b,D_{2})$. The second one is photon a is detected by D2 and photon b is detected by D1. The corresponding two-photon probability amplitude is $A(a,D_{2};b,D_{1})$. If these two different paths to trigger a two-photon coincidence count are indistinguishable, the two-photon probability function is
$P^{(2)}(\vec{r}_{1},t_{1};\vec{r}_{2},t_{2})$ (57)
$=$ $\langle|A(a,D_{1};b,D_{2})+A(a,D_{2};b,D_{1})|^{2}\rangle.$
When these two different paths are distinguishable, the two-photon probability function is
$P^{(2)}(\vec{r}_{1},t_{1};\vec{r}_{2},t_{2})$ (58)
$=$ $\langle|A(a,D_{1};b,D_{2})|^{2}\rangle+\langle|A(a,D_{2};b,D_{1})|^{2}\rangle,$
which will contribute to the background of two-photon probability function. Figure 13(a) and Eqs. (57) are valid for two-photon bunching of thermal light. $g^{(2)}(0)$ equals 2 for thermal light. There is another way to estimate the value of $g^{(2)}(0)$. Assuming all the photons have equal probability reaching D1 and D2, $|A(a,D_{1};b,D_{2})|$ equals $|A(a,D_{2};b,D_{1})|$. The are four terms in Eq. (57) and $P^{(2)}(0)$ equals $4|A|^{2}$, where $A$ is short for $|A(a,D_{1};b,D_{2})|$. There are 2 terms in Eq. (58) and $P^{(2)}(0)$ equals $2|A|^{2}$. The degree of second-order coherence equals the normalized two-photon probability function when these two detectors are at the same space-time coordinate,
$g^{(2)}(0)=p^{(2)}(0)=\frac{4|A|^{2}}{2|A|^{2}}=2.$ (59)
The reason why Eq. (58) can be treated as the background is that these two paths are distinguishable. The two photon detection events are independent and $P^{(2)}(\vec{r}_{1},t_{1};\vec{r}_{2},t_{2})$ equals $P^{(1)}(\vec{r}_{1},t_{1})P^{(1)}(\vec{r}_{2},t_{2})$.
Similar method can be employed to calculate the degree of second-order coherence of light in Fig. 13(b). There are four different paths for photons a and b triggering a two-photon coincidence count. The first one is photon a transmitting through $\alpha$ is detected by D1 and photon b transmitting through $\beta$ is detected by D2. The corresponding two-photon probability amplitude is $A(a,\alpha,D_{1};b,\beta,D_{2})$. The other three paths can be defined similarly and the corresponding two-photon probability amplitudes are $A(a,\alpha,D_{2};b,\beta,D_{1})$, $A(a,\beta,D_{1};b,\alpha,D_{2})$ and $A(a,\beta,D_{2};b,\alpha,D_{1})$. When these four different paths to trigger a two-photon coincidence count are indistinguishable, the two-photon probability function
$P^{(2)}(\vec{r}_{1},t_{1};\vec{r}_{2},t_{2})$
$=$ $\langle|A(a,\alpha,D_{1};b,\beta,D_{2})+A(a,\alpha,D_{2};b,\beta,D_{1})$
$+A(a,\beta,D_{1};b,\alpha,D_{2})+A(a,\beta,D_{2};b,\alpha,D_{1})|^{2}\rangle.$
When these four different paths are distinguishable, the two-photon probability function
$P^{(2)}(\vec{r}_{1},t_{1};\vec{r}_{2},t_{2})$
$=$ $\langle|A(a,\alpha,D_{1};b,\beta,D_{2})|^{2}\rangle+\langle|A(a,\alpha,D_{2};b,\beta,D_{1})|^{2}\rangle$
$+\langle|A(a,\beta,D_{1};b,\alpha,D_{2})|^{2}\rangle+\langle|A(a,\beta,D_{2};b,\alpha,D_{1})|^{2}\rangle$
Again, assuming photons have equal probability being detected by D1 and D2. The ratio between Eqs. (IV.1.2) and (IV.1.2) at the same space-time coordinate equals the degree of second-order coherence,
$g^{(2)}(0)=p^{(2)}(0)=\frac{16|A^{\prime}{}|^{2}}{4|A^{\prime}{}|^{2}}=4,$ (62)
where $|A^{\prime}{}|$ is short for $|A(a,\alpha,D_{1};b,\beta,D_{2})|$.
Similar method can be employed to calculate the degree of second-order coherence when there are $2N$ different paths to trigger a two-photon coincidence count. The corresponding degree of second-order coherence equals $2^{N}$, where $N$ is a positive integer Zhou et al. (2017).
Figure 14: Experimental scheme to observe two-photon superbunching with $2N$ different paths Zhou et al. (2017). P1, P2, …, PN are $N$ pinholes. RG1, RG2, …, RGN are $N$ rotating groundglasses.
Figure 14 shows the experimental scheme to observe two-photon superbunching with $2N$ different paths. Single-mode continuous-wave laser is employed as the light source. After passing through pinhole P1, the light is scattered by a rotating groundglass, RG1. Then the scattered light is filtered with another pinhole, P2. The size of P2 should be less than the coherence area of scattered light at P2 plane. Then the filtered light is scattered by RG2. Repeating the process $N$ times and superbunching pseudothermal light with $g^{(2)}(0)$ equaling $2^{N}$ can be generated.
Figure 15: Normalized second-order coherence functions with one (a), two (b), and three (c) rotating groundglasses Bai et al. (2017). $t_{1}-t_{2}$ is the time difference between two photon detection events at D1 and D2. $g^{(2)}(t_{1}-t_{2})$ is the normalized second-order coherence function in Glauber’s quantum optical coherence theory, which is equivalent to the normalized two-photon probability function.
Figure 15(a)-(c) shows the measured temporal second-order coherence function of light with one, two, and three rotating groundglasses, respectively Bai et al. (2017). $g^{(2)}(0)$ equals $1.96\pm 0.01$ in Fig. 15(a), in which there is one rotating groundglass and two-photon bunching is observed. The reason why it is less than 2 is due to part of the un-scattered laser light contribute to the two-photon coincidence count. The source with one rotating groundglass is pseudothermal light source invented by Martienssen and Spiller Martienssen and Spiller (1964). $g^{(2)}(0)$ equals $3.80\pm 0.04$ in Fig. 15(b) when there are two rotating groundglass. When there are three rotating groundglass, $g^{(2)}(0)$ equals $7.10\pm 0.07$ in Fig. 15(c). Two-photon superbunching is observed in Figs. 15(b) and 15(c).
Another method to increase the degree of second-order coherence of light is the two-photon probability amplitudes changes with time. Figure 16 shows the scheme for observing two-photon superbunching by changing the two-photon probability amplitudes. The light source shown in Fig. 16 as the light source in a HBT interferometer shown in Fig. 11. There are two different paths for photons a and b to trigger a two-photon coincidence count in a HBT interferometer. The two-photon probability amplitudes are $A(a,D_{1};b,D_{2})$ and $A(a,D_{2};b,D_{1})$.
Figure 16: Simplified superbunching pseudothermal light sourc Zhou et al. (2019b). IM is intensity modulator and RG is rotating groudglass.
The difference between usual pseudothermal light source and the one shown in Fig. 16in the HBT interferometer is $A(a,D_{1};b,D_{2})$ and $A(a,D_{2};b,D_{1})$ will change with time in the latter due to that intensity modulator (IM) is employed to modulate the intensity of incidental laser light. When these two different paths to trigger a two-photon coincidence count are indistinguishable, the temporal two-photon probability function
$P^{(2)}(t_{1};t_{2})$ $=$ $\langle|\sqrt{P_{a1}(t_{a1})P_{b2}(t_{b2})}K_{a1}K_{b2}$ (63)
$+\sqrt{P_{a2}(t_{a2})P_{b1}(t_{b1})}K_{a2}K_{b1}|^{2}\rangle$
where $P_{a1}(t_{a1})$ is the probability that photon a is the detected by D1 at time $t_{a1}$ and the meanings of other symbols are defined similarly.
Assuming point light source is employed and with the same method above, Eq. (63) can be simplified as Zhou et al. (2019b)
$P^{(2)}(t_{1}-t_{2})$ $\propto$ $p^{(2)}(t_{1}-t_{2})$
$\propto$ $\gamma(t_{1}-t_{2})[1+\text{sinc}^{2}\frac{\Delta\omega(t_{2}-t_{1})}{2}],$
where $\gamma(t_{1}-t_{2})=\langle I(t_{1})I(t_{2})\rangle/(\langle I(t_{1})\rangle\langle I(t_{2})\rangle)$ is the correlation function of laser light introduced by IM, $I(t_{1})$ and $I(t_{2})$ are the intensities of light before RG, respectively. When $\gamma(t_{1}-t_{2})$ equals 1, Eq. (IV.1.2) describes normal two-photon bunching of thermal light. When $\gamma(t_{1}-t_{2})$ is larger than 1, two-photon suerpbunching can be observed.
Figure 17: Observed two-photon superbunching with changing probability amplitudes Zhou et al. (2019b).
Figure 17 shows the experimental measured temporal second-order coherence function of superbunching pseudotheral light with the scheme shown in Fig. 16. Two-photon superbuching with $g^{(2)}(0)$ equaling 20.45 is observed Zhou et al. (2019b).
There are also other different schemes to increase the degree of second-order coherence of light Hong et al. (2012); Straka et al. (2018); Zhang et al. (2019, 2020), which can also be categorized into the above two methods.
IV.1.3 Second-order interference of one laser light beam
Assuming S is a single-mode continuous-wave laser in Fig. 11. Same method can be employed to calculate two-photon probability function of laser light in a HBT interferometer. There are two different paths for photons a and b to trigger a two-photon coincidence count. The probability amplitudes for these two paths are $A(a,D_{1};b,D_{2})$ and $A(a,D_{2};b,D_{1})$. If these two different ways are indistinguishable, the two-photon probability is
$P^{(2)}(\vec{r}_{1},t_{1};\vec{r}_{2},t_{2})$ (65)
$=$ $\langle|A(a,D_{1};b,D_{2})+A(a,D_{2};b,D_{1})|^{2}\rangle.$
It seems that two-photon bunching should also be observed for single-mode continuous-wave laser light in a HBT interferometer. However, it was proved that photon detection events are independent when a single-mode continuous-wave laser light beam in a HBT interferometer Arecchi (1965); Freed and Haus (1965). It is predicted in Glauber’s quantum optical coherence theory that the degree of second-order coherence of coherent state equals 1 Glauber (1963a), which indicates that photon detection events are independent in single-mode continuous-wave laser light.
The reason why there is no two-photon bunching of single-mode continuous-wave laser light lies at the difference between the superposition principles in classical and quantum theories. In classical theory, assuming there are two fields at space-time coordinate $(\vec{r},t)$, the final field is the superposition of these two fields,
$\vec{E}(\vec{r},t)=\vec{E}_{1}(\vec{r},t)+\vec{E}_{2}(\vec{r},t).$ (66)
If $\vec{E}_{1}(\vec{r},t)$ is identical to $\vec{E}_{2}(\vec{r},t)$, the superposition will give $2\vec{E}_{1}(\vec{r},t)$. The intensity at that point will be four times of the one of $\vec{E}_{1}(\vec{r},t)$.
This is not true for the superposition in quantum theory. Assuming there are two wave functions at space-time coordinate $(\vec{r},t)$, the final wave function is the superposition of these two,
$\Psi(\vec{r},t)=\Psi_{1}(\vec{r},t)+\Psi_{2}(\vec{r},t).$ (67)
If $\Psi_{1}(\vec{r},t)$ is identical to $\Psi_{2}(\vec{r},t)$, the superposition will not give $2\Psi_{1}(\vec{r},t)$. The results of superposition of two identical states equals one of the state, not two times of the individual state Bohm (2012). The wave function can be related to the probability amplitude Feynman et al. (2010),
[…]
The superposition of two probability amplitudes for two identical paths will not give two times of each probability amplitude, but only one probability amplitude. The two different paths in Eq. (65) are identical for single-mode continuous-wave laser light. The two-photon probability function is
$P^{(2)}(\vec{r}_{1},t_{1};\vec{r}_{2},t_{2})=\langle|A(S,D_{1};S,D_{2})|^{2}\rangle,$ (69)
no second-order interference exists and $A(S,D_{1};S,D_{2})$ is the probability amplitude for two photons emitted by S going to D1 and D2, respectively.
Comparing the two-photon probability functions of thermal and laser light in a HBT interferometer, the reason why these two cases are different is due to the initial phases of photons in thermal light are random and the initial phases of photons in laser light are identical within one coherence volume. This difference makes two different paths, $A(a,D_{1};b,D_{2})$ and $A(a,D_{2};b,D_{1})$ different for thermal light and identical for laser light. Hence two-photon bunching can be observed with thermal light and can not be observed for single-mode continuous-wave laser light.
IV.2 Second-order interference of two light beams
The first second-order interference of two light beams experiment are correlation measurement of photon pairs generated in the annihilation of positron-electron pairs Hanna (1948); Bleuler and Bradt (1948); Wu and Shaknov (1950). Pfleegor and Mandel employed the second-order correlation of light to prove there exist transient first-order interference pattern for two independent lasers at low light level Pfleegor and Mandel (1967). The most famous second-order interference of two light beams is the second-order interference of entangled photon pairs in a HOM interferometer Hong et al. (1987); Shih and Alley (1988); Alley and Shih (1986). Many other second-order interference of light experiments were performed, please see Refs. Belinski and Klyshko (1993); Paul (1986); Ou (2007); Shih (2020) for details. This section will concentrate on applying Feynman’s path integral to discuss the second-order interference of two light beams in a HOM interferometer.
Figure 18: Second-order interference of two light beams in a HOM interferometer. S1 and S2 are two light sources. BS is a 1:1 nonpolarizing beam splitter. D1 and D2 are two photo detectors. CC is coincidence count detection system.
The second-order interference of two light beams in a HOM interferometer is shown in Fig. 18. Two light beams emitted by source S1 and S2 are incident to two input ports of a 1:1 beam splitter (BS). The output light beams from BS are detected by two single-photon detectors, D1 and D2. The output of these two detectors are sent into a two-photon coincidence count detection system to measure the second-order correlation.
IV.2.1 Second-order interference of entangled photon pairs
Assume two photons in the same entangled photon pair are in the scheme shown in Fig. 18, which is similar as the original second-order interference with entangled photon pairs Hong et al. (1987); Shih and Alley (1988); Alley and Shih (1986). There are two different paths for photons a and b to trigger a two-photon coincidence count in D1 and D2. For simplicity, assume photon a is emitted by S1 and photon b is emitted by S2. The first path is photon a is detected by D1 and photon b is detected by D2. The probability for the path is 1/2 and the corresponding two-photon probability amplitude is $A_{a,D_{1};b,D_{2}}$. The second path is photon a is detected by D2 and photon b is detected by D1. The probability for the path is 1/2 and the corresponding two-photon probability amplitude is $A_{a,D_{2};b,D_{1}}$. If these two different paths are indistinguishable, the two-photon probability function is
$P^{(2)}(\vec{r}_{1},t_{1};\vec{r}_{2},t_{2})$ $=$ $\langle|A_{a,D_{1};b,D_{2}}+A_{a,D_{2};b,D_{1}}|^{2}\rangle.$ (70)
There is an extra $\pi/2$ phase of the reflected photon compared to the transmitted photon introduced by BS Loudon (2000), which can not be ignored in HOM interferometer as the one in HBT interferometer. Equation (70) can be simplified as
$P^{(2)}(\vec{r}_{1},t_{1};\vec{r}_{2},t_{2})$ (71)
$=$ $\langle|e^{i\varphi_{a}}K_{aD_{1}}e^{i\varphi_{b}}K_{bD_{2}}+e^{i\varphi_{a}}e^{i\pi/2}K_{aD_{2}}e^{i\varphi_{b}}e^{i\pi/2}K_{bD_{1}}|^{2}\rangle$
$=$ $\langle|K_{aD_{1}}K_{bD_{2}}-K_{aD_{2}}K_{bD_{1}}|^{2}\rangle,$
where $\varphi_{a}$ and $\varphi_{b}$ are the initial phases of photon a and b, respectively, $K_{aD_{1}}$ is the photon’s Feynman propagator for photon a emitted by S1 goes to D1 and the meanings of other symbols are defined similarly.
Since $\varphi_{a}$ and $\varphi_{b}$ are random for entangled photon pairs, the probability to detect one photon at D1 or D2 is a constant. The normalized two-photon probability function is proportional to the two-photon probability function, $P^{(2)}(\vec{r}_{1},t_{1};\vec{r}_{2},t_{2})$.
Both temporal and spatial second-order interference can be calculated. First calculate the temporal case. Assume D1 and D2 are in the symmetrical positions about BS, all the properties such as polarization, frequency are the same for photons a and b. With the same method as the one in Sect. IV, Eq. (71) can be simplified as
$P^{(2)}(t_{1}-t_{2})\propto 1-\text{sinc}^{2}\frac{\Delta\omega(t_{2}-t_{1})}{2},$ (72)
where $t_{1}$ and $t_{2}$ are the photon detection time at D1 and D2, respectively, $\Delta\omega$ is frequency bandwidth of entangle photon pairs. Point light sources are assumed to simplify the calculation.
Figure 19: HOM dip for entangled photon pairs with different frequency bandwidths. 1, 2, and 3 correspond to $\Delta\omega$, $2\Delta\omega$, and $4\Delta\omega$, respectively.
When the time difference between these two photon detection events is zero, $P^{(2)}(t_{1}-t_{2})$ equals zero, which means it is impossible to detect two photons at D1 and D2 simultaneously. These two photons are in the same output port of BS. When $|t_{1}-t_{2}|$ increases, $P^{(2)}(t_{1}-t_{2})$ increases from zero to one. $P^{(2)}(t_{1}-t_{2})$ equals 1 when $|t_{1}-t_{2}|$ is much longer than the coherence time, which indicates the detections of these two photons are independent.
Similar method can be employed to calculate spatial two-photon probability function
$P^{(2)}(x_{1}-x_{2})\propto 1-\text{sinc}^{2}\frac{d}{\lambda L}(x_{1}-x_{2}),$ (73)
where $x_{1}$ and $x_{2}$ are the transverse coordinates of D1 and D2, respectively, $d$ is the size of one dimension source. Paraxial approximation and one-dimension are assumed to simplify the results. Experiments with entangled photon pairs confirmed the result in Eq. (73) Kim et al. (2006); Lee and van Exter (2006); Devaux et al. (2019).
It is interesting to comparing two-photon bunching of thermal light in a HBT interferometer and HOM dip of entangled photon pair in a HOM interferometer. Two-photon bunching of thermal light shown in Fig. 12 looks like the contrary effect of HOM dip shown in Fig. 19. Photons have the tendency to bunch together when these two detectors are at symmetrical positions in Fig. 12 while photons have the tendency to avoid each other when two detectors are at symmetrical positions in Fig. 19. Even through different phenomena are observed with types of light in different interferometers, the physics of these two phenomena are the same. As shown in Eqs. (IV.1.1) and (71), they are results of two-photon interference. The only difference is that two-photon bunching of thermal light is caused by constructive two-photon interference and HOM dip with entangled photon pairs is caused by destructive two-photon interference.
Entangled photon pairs are not necessary to observe HOM dip shown in Fig. 19. The same HOM dip can be observed with two independent and identical single-photon states as the input of HOM interferometer. The theoretical calculations are exactly the same as the one for entangled photon pairs and the results are experimentally verified by several groups Legero et al. (2004); Beugnon et al. (2006); Flagg et al. (2010); Qian et al. (2016).
IV.2.2 Second-order interference of two classical light beams
The same method can be employed to calculate the second-order interference of two classical light beams in a HOM interferometer. Assume S1 and S2 are two independent and identical single-mode continuous-wave lasers. The polarizations and intensities of these two laser light beams are the same. There are three different ways to trigger a two-photon coincidence count in this case. The first one is these two photons are both emitted by S1. The probability is 1/4. There is one path and the corresponding two-photon probability amplitude is $A_{11,12}$. The second way is these two photons are both emitted by S2. The probability is 1/4. There is one path and the corresponding two-photon probability amplitude is $A_{21,22}$. The third way is the detected two photons are emitted by S1 and S2 one each. The probability is 1/2 and there are two different paths in this way. The corresponding two-photon probability amplitudes are $A_{11,22}$ and $A_{12,21}$. If these four different paths are indistinguishable, the two-photon probability function is
$P^{(2)}(\vec{r}_{1},t_{1};\vec{r}_{2},t_{2})$ (74)
$=$ $\langle|\frac{1}{\sqrt{4}}A_{11,12}+\frac{1}{\sqrt{4}}A_{21,22}+\frac{1}{\sqrt{4}}(A_{11,22}+A_{12,21})|^{2}\rangle$
$=$ $\langle|\frac{1}{\sqrt{4}}e^{i\varphi_{11,12}}K_{11,12}+\frac{1}{\sqrt{4}}e^{i\varphi_{21,22}}K_{21,22}$
$+\frac{1}{\sqrt{4}}(e^{i\varphi_{11,22}}K_{11,22}+e^{i\varphi_{12,21}}K_{12,21})|^{2}\rangle,$
where $\varphi_{11,12}$ is the phase related to the initial phase of these two photons and extra $\pi/2$ phase due to the reflection on BS. The last $1/\sqrt{4}$ in Eq. (74) is the product of $1/\sqrt{2}$ and $1/\sqrt{2}$, where the first $1/\sqrt{2}$ is due to the probability of these two paths equals 1/2 and the second $1/\sqrt{2}$ is the normalization parameter. Table 1 summarizes the phases for these four different paths, where $\varphi_{1}$ and $\varphi_{2}$ are the initial phases of photons emitted by S1 and S2, respectively.
Table 1: Phases for different two-photon paths.
$\varphi_{11,12}$ $\varphi_{1}+\varphi_{1}+\pi/2$
$\varphi_{21,22}$ $\varphi_{2}+\varphi_{2}+\pi/2$
$\varphi_{11,22}$ $\varphi_{1}+\varphi_{2}$
$\varphi_{12,21}$ $\varphi_{1}+\varphi_{2}+\pi$
With the phases summarized in Table 1, Eq. (74) can be simplified as
$P^{(2)}(\vec{r}_{1},t_{1};\vec{r}_{2},t_{2})$ (75)
$\propto$ $\langle|e^{i(2\varphi_{1}+\pi/2)}K_{11,12}+e^{i(2\varphi_{2}+\pi/2)}K_{21,22}$
$+e^{i(\varphi_{1}+\varphi_{2})}(K_{11,22}-K_{12,21}|^{2}\rangle.$
Since these two lasers are independent, Eq. (75) can be simplified as
$P^{(2)}(\vec{r}_{1},t_{1};\vec{r}_{2},t_{2})$
$=$ $\langle|K_{11,12}|^{2}\rangle+\langle|K_{21,22}|^{2}\rangle+\langle|K_{11,22}-K_{12,21}|^{2}\rangle.$
Assuming S1 and S2 are point light sources, temporal two-photon probability function can be obtained via simplifying Eq. (IV.2.2)
$P^{(2)}(t_{1}-t_{2})$ $\propto$ $1+1+2[1-\cos\Delta\omega(t_{1}-t_{2})]$ (77)
$\propto$ $1-\frac{1}{2}\cos\Delta\omega(t_{1}-t_{2}),$
where $\Delta\omega=\omega_{1}-\omega_{2}$ is frequency difference between these two laser light beams. There is second-order interference pattern by superposing two independent laser light beams. Unlike the transient first-order interference pattern of two independent laser light beams, the second-order interference pattern in Eq. (77) is observed over long time average.
Figure 20: Temporal second-order interference pattern of two independent laser light beams in a HOM interferometer Liu et al. (2016)
Figure 20 shows the temporal second-order interference pattern with two tunable single-mode continuous-wave lasers. The frequency bandwidth of laser light is 100 kHz. The frequency difference between these two lasers are 189 MHz, 461 MHz, 1.11 GHz, and 3.0 GHz in Figs. 20(a)-(d), respectively. Periodic interference pattern is observed in Figs. 20(a)-(b). There are also periodic structure in Figs. 20(c)-(d). The reason why it can not be observed is the period of the interference pattern is too small to observe in the present time scale Liu et al. (2016).
The same method can be employed to calculate the second-order interference of two independent thermal and laser light beams. Assume S1 is single-mode continuous-wave laser and S2 is thermal light source in Fig. 18. The intensities, polarizations, frequencies of the light beams emitted by these two sources are the same. There are three different ways for D1 and D2 detecting a two-photon coincidence count. The first way is these two photons are emitted by laser S1. The probability is 1/4 and there is one path, $A_{11,12}$. The second way is these two photons are emitted by thermal light source S2. The probability is 1/4 and there are two paths, $A_{2a1,2b2}$ and $A_{2a2,2b1}$. The third way is these two photons are emitted by S1 and S2 each. The probability is 1/2 and there are two different paths, $A_{11,22}$ and $A_{12,21}$. If the five different paths are indistinguishable, the two-photon probability function equals
$P^{(2)}(\vec{r}_{1},t_{1};\vec{r}_{2},t_{2})$ (78)
$=$ $\langle|\frac{1}{\sqrt{4}}A_{11,12}+\frac{1}{\sqrt{8}}(A_{2a1,2b2}+A_{2a2,2b1})$
$+\frac{1}{\sqrt{4}}(A_{11,22}+A_{12,21})|^{2}\rangle,$
With the same method as the one for Eq. (74), Eq. (78) can be simplified as Liu et al. (2014b)
$P^{(2)}(\vec{r}_{1},t_{1};\vec{r}_{2},t_{2})$ (79)
$=$ $2\langle|K_{11,12}|^{2}\rangle+\langle|K_{2a1,2b2}+K_{2a2,2b1}|^{2}\rangle$
$+2\langle|K_{11,22}-K_{12,21}|^{2}\rangle.$
With paraxial approximation, one dimension spatial two-photon probability function can be obtained via Eq. (79),
$p^{(2)}(x_{1}-x_{2})$
$\propto$ $1+\frac{1}{4}\text{sinc}^{2}\frac{\pi l}{L\lambda}(x_{1}-x_{2})[1-2\cos\frac{2\pi d}{L\lambda}(x_{1}-x_{2})],$
where $l$ is the length of thermal light source, $d$ is the distance between the central points of thermal light source and the virtual source of laser relative to BS Liu et al. (2014b).
Figure 21: Second-order interference pattern of thermal and laser light (b) and the second-order coherence function of laser and thermal light beams in the interferometer alone (a) Liu et al. (2014b).
Figure 21(b) shows the second-order interference pattern of thermal and laser light beams in a HOM interferometer, which is consistent with Eq. (IV.2.2). For comparison, Fig. 21(a) shows the measured second-order coherence functions when one of the incidental light beams is blocked, which is equivalent to laser and thermal light in a HBT interferometer Liu et al. (2014b).
The second-order interference between two independent thermal light beams Liu et al. (2013), classical and nonclassical light beams Hong et al. (1988) and other two independent light beams can also be calculated in the same method above. The main difference between the second-order of two independent classical and nonclassical light beams is that the visibility of the second-order interference pattern of two independent nonclassicla light beams can reach 100%, while the maximal visibility of the second-order interference pattern between two independent classical light beam is 50% Mandel (1983). The reason is clearly shown in Eqs. (70), (74), and (78). In the second-order interference of two independent classical light beams, it is impossible to distinguish one source emitting two photons from two sources emitting a photon each, which will contribute to the background coincidence counts. If special two-photon coincidence count system can distinguish these two types of coincidence count, the visibility of second-order interference pattern of two independent classical light beams can exceed 50%. For instance, Kaltenbaek et al. observed the visibility of second-order interference pattern with two independent laser light beams exceeding 85% by employing sum frequency generation to distinguish these two different types of coincidence count. With two non-independent classical light beams and normal two-photon coincidence count system, Sadana et al. observed near 100% visibility second-order interference pattern Sadana et al. (2019). The results do not conflict with Mandel’s conclusion due to Mandel’s conclusion is valid for two independent light beams and normal two-photon coincidence count detection system Mandel (1983).
IV.3 Second-order interference of multiple light beams
Feyman’s path integral can be employed to calculate the second-order interference of more than two independent light beams. Figure 22 shows the second-order interference of three independent light beams. S1, S2 and S3 are three independent light sources. D1 and D2 are two single-photon detectors. The output of two detectors are sent into a two-photon coincidence count system to measure the second-order correlation, which is not shown in the figure.
Figure 22: Second-order interference of three independent light beams
Assume S1, S2 and S3 are three independent and identical single-photon sources. There are three different ways to trigger a two-photon coincidence count. The first one is the two detected photons are emitted by S1 and S2. The probability is 1/3. There are two different paths and the corresponding two-photon probability amplitudes are $A_{11,22}$ and $A_{12,21}$. The second way is these two photons are emitted by S1 and S3. The probability is 1/3. There are two paths and the corresponding two-photon probability amplitudes are $A_{11,32}$ and $A_{12,31}$. The third way is these two photons are emitted by S2 and S3. The probability is 1/3. There are two paths and the corresponding two-photon probability amplitudes are $A_{21,32}$ and $A_{22,31}$. If only one single-photon source emit a photon, it is impossible to trigger a two-photon coincidence count. If two single-photon sources emit photons simultaneously, these three different ways are distinguishable. One can measure the status of the single-photon source to ensure which two sources emitted photons. The probability for three independent single-photon sources simultaneously emitting photons is much smaller than the one of two single-photon sources simultaneously emitting photons, which will be ignored in the calculation.
If every two different paths to trigger a two-photon coincidence count are indistinguishable, the two-photon probability function is
$P^{(2)}(\vec{r}_{1},t_{1};\vec{r}_{2},t_{2})$ (81)
$=$ $\langle|\frac{1}{\sqrt{6}}(A_{11,22}+A_{12,21})|^{2}+|\frac{1}{\sqrt{6}}(A_{11,32}+A_{12,31})|^{2}$
$+|\frac{1}{\sqrt{6}}(A_{21,32}+A_{22,31})|^{2}\rangle.$
With the same method above, Eq. (81) can be simplified by considering the three single-photon sources are independent,
$P^{(2)}(\vec{r}_{1},t_{1};\vec{r}_{2},t_{2})$ (82)
$=$ $\langle|K_{11}K_{22}+K_{12}K_{21}|^{2}\rangle+\langle|K_{11}K_{32}+K_{12}K_{31}|^{2}\rangle$
$+\langle|K_{21}K_{32}+K_{22}K_{31}|^{2}\rangle,$
where $K_{11}$ is the photon’s Feynman propagator from S1 to D1 and the meanings of other symbols are defined similarly. Assuming S1, S2 and S3 are point source and in the same source plane, the distance between the source and the observation planes is much larger than one between sources, the spatial two-photon probability function can be obtained via Eq. (82)
$P^{(2)}(x_{1}-x_{2})$
$=$ $3+\cos\frac{2\pi d_{12}}{L\lambda}(x_{1}-x_{2}-\frac{d_{23}}{2})$
$+\cos\frac{2\pi d_{13}}{L\lambda}(x_{1}-x_{2})+\cos\frac{2\pi d_{23}}{L\lambda}(x_{1}-x_{2}+\frac{d_{12}}{2}),$
where paraxial approximation and one-dimension have been assumed, $\lambda$ is the wavelength of light and single-frequency is employed to simplify the calculation, $d_{12}$ is the distance between S1 and S2 and relative positions of these three sources are shown in Fig. 8. The meanings of other symbols are defined similarly.
Figure 23: Second-order interference pattern of three independent light beams. $x_{1}-x_{2}$ is the relative position of D1 and D2.
Figure 23 shows the second-order interference pattern of three single-photon sources shown in Fig. 22. The patterns will change as the distances between these three sources change. Compared to the first-order interference pattern of three light beams shown in Fig. 9, the visibility of the second-order interference pattern with three independent light beams can not reach 100% even when single-photon states are employed.
The second-order interference of three independent laser light beams can also be calculated in the same method. Assume S1, S2 and S3 are three independent and single-mode continuous-wave lasers and the polarizations of these three laser light beams are identical. There are six different ways to trigger a two-photon coincidence count in Fig. 22. The first way is these two photons are both emitted by S1. The probability is $I_{1}^{2}/(I_{1}+I_{2}+I_{3})^{2}$, where $I_{1}$, $I_{2}$, and $I_{3}$ are the intensities of laser light beams emitted by S1, S2 and S3, respectively. There is one path and the corresponding two-photon probability amplitude is $A_{11,12}$. The second ways is these two photons are both emitted by S2. The probability is $I_{2}^{2}/(I_{1}+I_{2}+I_{3})^{2}$. There is one path and the two-photon probability amplitude is $A_{21,22}$. The third way is these two photons are both emitted by S3. The probability is $I_{3}^{2}/(I_{1}+I_{2}+I_{3})^{2}$. There is one path and the probability amplitude is $A_{31,32}$. The fourth way is these two photons are emitted by S1 and S2 one each. The probability is $2I_{1}I_{2}/(I_{1}+I_{2}+I_{3})^{2}$. There are two paths and the probability amplitudes are $A_{11,22}$ and $A_{12,21}$. The fifth way is these two photons are emitted by S1 and S3 one each. The probability is $2I_{1}I_{3}/(I_{1}+I_{2}+I_{3})^{2}$ and the two paths are $A_{11,32}$ and $A_{12,31}$. The sixth way is these two photons are emitted by S2 and S3 one each. The probability is $2I_{2}I_{3}/(I_{1}+I_{2}+I_{3})^{2}$ and the two paths are $A_{21,32}$ and $A_{22,31}$. If all the nine different paths to trigger a two-photon coincidence count are indistinguishable, the two-photon probability function is
$P^{(2)}(\vec{r}_{1},t_{1};\vec{r}_{2},t_{2})$
$\propto$ $\langle|I_{1}A_{11,12}+I_{2}A_{21,22}+I_{3}A_{31,32}$
$+\sqrt{I_{1}I_{2}}(A_{11,22}+A_{12,21})$
$+\sqrt{I_{1}I_{3}}(A_{11,32}+A_{12,31})+\sqrt{I_{2}I_{3}}(A_{21,32}+A_{22,31})|^{2}\rangle,$
where the common factor $1/(I_{1}+I_{2}+I_{3})$ has been ignored, $\sqrt{I_{1}I_{2}}=\sqrt{2I_{1}I_{2}}\times 1/\sqrt{2}$.
Considering three lasers are independent, Eq. (IV.3) can be simplified as
$P^{(2)}(\vec{r}_{1},t_{1};\vec{r}_{2},t_{2})$
$=$ $I_{1}^{2}\langle|K_{11}K_{12}|^{2}\rangle+I_{2}^{2}\langle|K_{21}K_{22}|^{2}\rangle+I_{3}^{2}\langle|K_{31}K_{32}|^{2}\rangle$
$+I_{1}I_{2}\langle|K_{11}K_{22}+K_{12}K_{21}|^{2}\rangle$
$+I_{1}I_{3}\langle|K_{11}K_{32}+K_{12}K_{31}|^{2}\rangle+I_{2}I_{3}\langle|K_{21}K_{32}+K_{22}K_{31}|^{2}\rangle.$
Assuming the intensities of these three light beams are equal and single frequency light is emitted, Eq. (IV.3) can be simplified as
$P^{(2)}(x_{1}-x_{2})$
$\propto$ $4.5+\cos\frac{2\pi d_{12}}{L\lambda}(x_{1}-x_{2}-\frac{d_{23}}{2})$
$+\cos\frac{2\pi d_{13}}{L\lambda}(x_{1}-x_{2})+\cos\frac{2\pi d_{23}}{L\lambda}(x_{1}-x_{2}+\frac{d_{12}}{2}),$
where paraxial approximation and one-dimension are assumed to simplify the calculations. Comparing the second-order interference pattern of three independent single-photon states, Eq. (IV.3) and the second-order interference pattern of three independent single-mode continuous-wave lasers, the interference patterns are the same except the constant changes from 3 to 4.5. The reason for this difference is that laser light can not avoid two photons emitted from the same source with normal two-photon coincidence count detection system, which is exactly the same as the one for the difference between the visibility of the second-order interference pattern of two independent classical and non-classical light beams Mandel (1983).
IV.4 Subwavelength interference
In the above parts of Sect. IV, we have introduced how to employ Feynman’s path integral to calculate the second-order interference of one, two, and multiple light beams. As an example of how quantum optical coherence theory based on Feynman’s path integral will help us to understand the physics of optical coherence, subwavelength interference is analyzed in Feynman’s path integral.
Taking the first- and second-order interference of light emitted by two independent sources shown in Fig. 5 as an example to illustrate subwavelength interference. The transient first-order interference pattern of two independent single-mode continuous-wave lasers is
$P^{(1)}(x)\propto 1+\cos(\frac{2\pi d}{\lambda L}x+\varphi).$ (87)
Putting two detectors in the detection plane in Fig. 5, the second-order interference pattern of two independent single-mode continuous-wave lasers is proportional to
$p^{(2)}(x_{1}-x_{2})\propto 1+\frac{1}{2}\cos\frac{2\pi d}{\lambda L}(x_{1}-x_{2}).$ (88)
Comparing Eq. (88) with Eq. (87) , if these two detectors are moving in the opposite direction with the same speed, the period of the second-order interference pattern will be half of the one of the first-order interference pattern, which is called subwavelength interference.
Subwavelength interference can be understood better with the concept of de Broglie wavelength. The de Broglie wavelength of a photon equals $h/p$, where $h$ is Planck constant and $p$ is the momentum of the photon. If two photons can be treated as a whole, the corresponding de Broglie wavelength equals $h/(2p)$, which is only half of the wavelength of one photon. The same goes for the de Broglie wavelength of $N$ photon equals $h/(Np)$, where $N$ is a positive integer larger than 1 Jacobson et al. (1995). Two-photon subwavelength interference was first observed with entangled photon pairs Fonseca et al. (1999); Boto et al. (2000); D’Angelo et al. (2001). Then it was found that subwavelength interference can be observed with thermal light Scarcelli et al. (2004); Xiong et al. (2005); Zhai et al. (2005) and laser light Liu and Zhang (2010).
Figure 24: Scheme for subwavelength interference with two laser light beams Liu and Zhang (2010)
Figure 24 shows the scheme for observing two-photon subwavelength with two independent laser light beams Liu and Zhang (2010), which is equivalent to a two point source interferometer. Two independent laser light beams was obtained by splitting a single-mode continuous-wave laser light beam by a beam splitter (BS1) and randomize the phase of one beam by a mirror mounted on a piezoelectric ceramic (PM). One beam splitter (BS2), two single-photon detectors (D1 and D2), and a two-photon coincidence count detection system (CC) are employed to measure the second-order interference pattern.
Figure 25: The observed normal second-order interference pattern Liu and Zhang (2010).
Figure 25(b) shows the normal second-order interference pattern by fixing the position of D2 and scanning the position of D1. The period of the observed second-order interference pattern is 800 $\mu$m. The visibility is $(46\pm 2)\%$, which does not exceed 50%. Figure 25(a) shows the single-photon counting rates observed by two single-photon detectors, respectively. There is no first-order interference pattern observed.
Figure 26: Two-photon subwavelength interference pattern Liu and Zhang (2010).
Figure 26(b) shows the observed two-photon subwavelength interference pattern by by scanning the positions of D1 and D2 in the opposite directions. The period of the second-order interference pattern is 400 $\mu$m, which is exactly half of the one shown in Fig. 25(b) and indicates subwavelength interference is observed. The visibility of the second-order interference pattern equals $(48\pm 4)\%$. One may argue that it is meaningless to observe subwavelength by scanning two detectors in the opposite directions Liu et al. (2014c). However, one can obtain subwavelength interference with two detectors scanning in the same direction for classical light with suitable changes on the experimental scheme Cao et al. (2010), which is similar as the one with entangled photon pairs Fonseca et al. (1999); D’Angelo et al. (2001). As shown below, there is no such big difference between the observed subwavelength interference patterns for detectors scanning in the same direction and opposite directions.
In order to see the physics of subwavelength interference, take the second-order interference of two single-mode continuous-wave laser with the same phase as an example. Assume the intensities and polarizations of these two laser light beams are the same. There are four different paths to trigger a two-photon coincidence count in Fig. 24. If these four difference paths are indistinguishable , the two-photon probability function is
$P^{(2)}(x_{1};x_{2})$ (89)
$=$ $\langle|A_{11,12}+A_{21,22}+A_{11,22}+A_{12,21}|^{2}\rangle$
$=$ $|e^{i2\varphi}(K_{11}K_{12}+K_{21}K_{22}+K_{11}K_{22}+K_{12}K_{21}|^{2}\rangle$
$=$ $|K_{11}K_{12}+K_{21}K_{22}+K_{11}K_{22}+K_{12}K_{21}|^{2},$
where $\varphi$ is the initial phases of photons in these two beams. Equation (89) can be re-arranged as
$P^{(2)}(x_{1};x_{2})$ (90)
$\propto$ $4I_{1}I_{2}$
$+I_{1}(K_{12}K_{22}^{*}+K_{22}K_{12}^{*}+K_{22}K_{12}^{*}+K_{12}K_{22}^{*})$
$+I_{2}(K_{11}K_{21}^{*}+K_{11}K_{21}^{*}+K_{11}K_{21}^{*}+K_{21}K_{11}^{*})$
$+K_{11}K_{12}K_{21}^{*}K_{22}^{*}+K_{21}K_{22}K_{11}^{*}K_{12}^{*}$
$+K_{11}K_{22}K_{12}^{*}K_{21}^{*}+K_{12}K_{21}K_{11}^{*}K_{22}^{*},$
where $I_{1}=|K_{11}|^{2}=|K_{21}|^{2}$ and $I_{2}=|K_{12}|^{2}=|K_{22}|^{2}$, which equals half of the intensities detected by D1 and D2, respectively. Assume point light sources and single-frequency light beams are employed, Eq. (90) can be simplified as Liu and Zhang (2010)
$P^{(2)}(x_{1};x_{2})$ (91)
$\propto$ $4I_{1}I_{2}$
$+4I_{1}I_{2}\cos\frac{2\pi d}{\lambda L}x_{2}$
$+4I_{1}I_{2}\cos\frac{2\pi d}{\lambda L}x_{1}$
$+2I_{1}I_{2}\cos\frac{2\pi d}{\lambda L}(x_{1}+x_{2})$
$+2I_{1}I_{2}\cos\frac{2\pi d}{\lambda L}(x_{1}-x_{2}).$
The last two lines on the right hand side of Eq. (91) corresponds to two-photon subwavelength when two detectors scanning in the same direction and opposition directions, respectively. $2I_{1}I_{2}\cos\frac{2\pi d}{\lambda L}(x_{1}+x_{2})$ corresponds to $K_{11}K_{12}K_{21}^{*}K_{22}^{*}+K_{21}K_{22}K_{11}^{*}K_{12}^{*}$ in Eq. (90), which is a result of two-photon interference for these two photons emitted by the same source. It is consistent with the condition for observing subwavelength interference with entangled photon pairs Fonseca et al. (1999); D’Angelo et al. (2001). $2I_{1}I_{2}\cos\frac{2\pi d}{\lambda L}(x_{1}-x_{2})$ corresponds to $K_{11}K_{22}K_{12}^{*}K_{21}^{*}+K_{12}K_{21}K_{11}^{*}K_{22}^{*}$ in Eq. (90), which is a result of two-photon interference for these two photons emitted by one source each. If one can eliminate other terms and only keep one of the last two terms in Eq. (91), one can observe subwavelength interference. The experiment reported in Ref. Liu and Zhang (2010) employed random phases between these two lasers to eliminate all the other terms except the constant and the last term in Eq. (91). Subwavelength interference is observed when two detectors are scanning in the opposite directions. Equations (90) and (91) shows that there is no intrinsic difference between the subwavelength interference when these two detectors scanning in the same direction and opposite directions. They are results of two-photon interference with different two-photon paths.
V Third- and higher-order interference of light
As mentioned above, the third- and higher-order interference of light is an natural extension of the second-order interference of light. The method introduced above can also be employed to calculate the third- and higher-order interference of light.
The studies of third- and higher-order interference dates back to 1963 when Glauber gave a unified definition of $N$th-order coherence function of light Glauber (1963a, c). Greenberg et al. generalized entangled photon pair to three-photon entanglement Greenberger et al. (1989, 1990). It is difficult to generate entangled three-photon directly. Bouwmeester et al. employed two entangled photon pairs to generate entangled three-photon by post-selection Bouwmeester et al. (1999). Mitchell et al. employed a laser photon and an entangled photon pair to generate entangled three-photon and observed three-photon subwavelength interference Mitchell et al. (2004). In 2010, Hübel et al. employed cascade periodically poled nonlinear crystals to generate entangled three-photon directly Hübel et al. (2010). Based on the work of Bouwmeester et al. Bouwmeester et al. (1999), Walther et al. observed four-photon subwavelength interference with two entangled photon pairs Walther et al. (2004) and Pan’s group observed as high as 12-photon subwavelength interference Zhao et al. (2004); Pan et al. (2012); Zhong et al. (2018).
Ghost imaging was originally realized with entangled photon pairs Pittman et al. (1995) and later with thermal light Gatti et al. (2004); Cheng and Han (2004); Cai and Zhu (2005); Valencia et al. (2005) via the second-order interference of light. Ghost imaging via the third- and higher-order interference of light was discussed for entangled three-photon Wen et al. (2007) and thermal light Ou and Kuang (2007); Bai and Han (2007); Agafonov et al. (2008); Zhou et al. (2010); Cao et al. (2008); Chan et al. (2009); Liu et al. (2009).
In the following, third-order interference of light is taken as an example to show how to employ Feyman’s path integral theory to calculate the third- and high-order interference of light.
V.1 Third-order interference of one light beam
Figure 27 shows the third-order HBT interferometer by analogy of HBT interferometer shown in Fig. 11. The light emitted by light source (S) is split into two via a 1:2 non-polarizing beam splitter BS1. The transmitted light beam is split into two via a 1:1 non-polarizing beam splitter. The intensities detected by D1, D2, and D3 are equal. TCC is short for three-photon coincidence count and is usually written as CC for convenience. The distance between the source and detector planes are equal.
Figure 27: Third-order HBT interferometer. TCC is short for three-photon coincidence count and is usually written as CC for convenience. The meanings of other symbols are the same as the one in HBT interferometer.
Assume S is thermal light source in Fig. 27. There are six different paths to trigger a three-photon coincidence count. Figure 28 shows the six different ways. For instance, Fig. 28(a) show the path for photon a is detected by D1, photon b is detected by D2 and photon c is detected by D3. If all the six different ways are indistinguishable, the three-photon probability function is
$P^{(3)}(\vec{r}_{1},t_{1};\vec{r}_{2},t_{2};\vec{r}_{3},t_{3})$ (92)
$=$ $\langle|A_{a1,b2,c3}+A_{a1,b3,c2}+A_{a2,b1,c3}$
$+A_{a2,b3,c1}+A_{a3,b1,c2}+A_{a3,b2,c1}|^{2}\rangle.$
The initial phases of photons in thermal light are random, Eq. (92) can be simplified as
$P^{(3)}(\vec{r}_{1},t_{1};\vec{r}_{2},t_{2};\vec{r}_{3},t_{3})$ (93)
$=$ $\langle|K_{a1}K_{b2}K_{c3}+K_{a1}K_{b3}K_{c2}+K_{a2}K_{b1}K_{c3}$
$+K_{a2}K_{b3}K_{c1}+K_{a3}K_{b1}K_{c2}+K_{a3}K_{b2}K_{c1}|^{2}\rangle,$
where $K_{a1}$ is the photon’s Feynman propagator from S to D1.
Figure 28: Third-order HBT interferometer. TCC is short for three-photon coincidence count and is usually written as CC for convenience. The meanings of other symbols are the same as the one in HBT interferometer.
With the same method as the one for thermal light in a HBT interferometer introduced in Sect. IV.1.1, the spatial three-photon probability function is Liu and Shih (2009)
$P^{(3)}(x_{1};x_{2};x_{3})$
$\propto$ $1+\text{sinc}^{2}\frac{\pi d}{L\lambda}(x_{1}-x_{2})$
$+\text{sinc}^{2}\frac{\pi d}{L\lambda}(x_{2}-x_{3})+\text{sinc}^{2}\frac{\pi d}{L\lambda}(x_{3}-x_{1})$
$+2\text{sinc}\frac{\pi d}{L\lambda}(x_{1}-x_{2})\text{sinc}\frac{\pi d}{L\lambda}(x_{2}-x_{3})\text{sinc}\frac{\pi d}{L\lambda}(x_{3}-x_{1}),$
where paraxial approximation and one-dimension have been employed, $\lambda$ is the wavelength of thermal light and single frequency light is assumed, $d$ is the length of one-dimension thermal light source, $L$ is the distance between the source and detector planes. Figure 29(a) shows the simulated spatial third-order coherence function of thermal light in a third-order HBT interferometer. Figure 29(b) shows the measured three-photon coincidence counts vs. the transverse distances between these three detectors. The theoretical and experimental results are consistent with each other.
Figure 29: Simulated Liu and Shih (2009) and measured Zhou et al. (2010) spatial third-order coherence functions of thermal light in the third-order interferometer.
Temporal three-photon probability function can be obtained in the same way Zhou et al. (2010)
$P^{(3)}(t_{1};t_{2};t_{3})$
$\propto$ $1+\text{sinc}^{2}\frac{\Delta\omega}{2}(t_{1}-t_{2})$
$+\text{sinc}^{2}\frac{\Delta\omega}{2}(t_{2}-t_{3})+\text{sinc}^{2}\frac{\Delta\omega}{2}(t_{3}-t_{1})$
$+2\text{sinc}\frac{\Delta\omega}{2}(t_{1}-t_{2})\text{sinc}\frac{\Delta\omega}{2}(t_{2}-t_{3})\text{sinc}\frac{\Delta\omega}{2}(t_{3}-t_{1}),$
where $\Delta\omega$ is the frequency bandwidth of thermal light and point thermal light source is assumed.
The degree of third order coherence, $g^{(3)}(0)$, of thermal light equals 6, which means that the probability to detect three photons when these three detectors are in the same space-time coordinate is 6 times of the one when these three detectors are in different coherence volumes of thermal light. It is three-photon bunching of thermal light. Three-photon superbunching can be observed with superbunching pseudothermal light Zhou et al. (2019a) and the degree of $N$th-oder coherence of thermal light equals $N!$ Liu and Shih (2009).
The third-order interference of single-mode continuous-wave laser light in the third-order HBT interferometer is simple. All the six different paths shown in Fig. 28 to trigger a three-photon coincidence count are identical. The superposition of these six probability amplitudes will not give three-photon interference. The reason is the same as the one for there is no two-photon interference for single-mode continuous-wave laser light in a HBT interferometer. No three-photon bunching can be observed when single-mode continuous-wave laser light in the third-order interferometer.
V.2 Third-order interference of two independent light beams
Figure 30 shows the scheme for the third-order interference of two independent light beams. Assume S1 is a single-photon source and S2 is a single-mode continuous-wave laser. There are two different ways to trigger a three-photon coincidence count in Fig. 30. The first way is one photon is emitted by S1 and two photons are emitted by S2. The probability is $3I_{1}I_{2}^{2}/(3I_{1}I_{2}^{2}+I_{2}^{3})$, where $I_{1}$ and $I_{2}$ are the intensities emitted by S1 and S2, respectively. There are three different paths and the corresponding three-photon probability amplitudes are $A_{11,22,23}$, $A_{12,21,23}$, and $A_{13,21,22}$, respectively. The second way is these three photons are all emitted by S2. The probability is $I_{2}^{3}/(I_{1}I_{2}^{2}+I_{2}^{3})$. There is one path and the corresponding three-photon probability is $A_{21,22,23}$.
Figure 30: Scheme for the third-order interference of two independent light beams.
If these four different paths are indistinguishable, the three-photon probability function is
$P^{(3)}(\vec{r}_{1},t_{1};\vec{r}_{2},t_{2};\vec{r}_{3},t_{3})$ (96)
$=$ $\langle|\sqrt{\frac{3I_{1}I_{2}^{2}}{3I_{1}I_{2}^{2}+I_{2}^{3}}}\frac{1}{\sqrt{3}}(A_{11,22,23}+A_{12,21,23}+A_{13,21,22})$
$+\sqrt{\frac{I_{2}^{3}}{3I_{1}I_{2}^{2}+I_{2}^{3}}}A_{21,22,23}|^{2}\rangle.$
The initial phases of photons emitted by S1 are random and emitted by S2 are identical within one coherence volume, Eq. (96) can be simplified as
$P^{(3)}(\vec{r}_{1},t_{1};\vec{r}_{2},t_{2};\vec{r}_{3},t_{3})$ (97)
$=$ $\frac{3I_{1}}{3I_{1}+I_{2}}\langle|K_{11}K_{22}K_{23}+K_{12}K_{21}K_{23}+K_{13}K_{21}K_{22}|^{2}\rangle$
$+\frac{I_{2}}{3I_{1}+I_{2}}\langle|K_{21}K_{22}K_{23}|^{2}\rangle.$
Assume S1 and S2 are point light sources. With the same method and approximations above, one-dimension spatial three-photon probability function is
$P^{(3)}(x_{1};x_{2};x_{3})$ (98)
$\propto$ $3I_{1}^{2}[3+2\cos\frac{2\pi d}{L\lambda}(x_{1}-x_{2})+2\cos\frac{2\pi d}{L\lambda}(x_{2}-x_{3})$
$+2\cos\frac{2\pi d}{L\lambda}(x_{1}-x_{3})]+I_{2}^{2}.$
There is third-order interference pattern. The visibility of the third-order interference pattern is dependent on the ratio between the intensities of light emitted by single-photon source and single-mode continuous-wave laser.
Similar method can be employed to calculate the third-order interference of two other types of light beams.
V.3 Third-order interference of multiple independent light beams
Figure 31 shows the scheme for the third-order interference of three independent light beams. Assume S1, S2, and S3 are three identical and independent single-photon sources. The intensities, polarizations, and the frequencies are the same for the photons emitted by these three sources. There are six different paths to trigger a three-photon coincidence count. The corresponding three probability amplitudes are $A_{11,22,33}$, $A_{11,23,31}$, $A_{12,21,33}$, $A_{12,23,31}$, $A_{13,22,31}$, and $A_{13,21,32}$, where $A_{11,22,33}$ is the probability amplitude for photon emitted by S1 is detected by D1, photon emitted by S2 is detected by D2, and photon emitted by S3 is detected by D3. The meanings of other symbols are defined similarly.
Figure 31: Scheme for the third-order interference of three independent light beams
If six different paths are indistinguishable, three-photon probability function is
$P^{(3)}(\vec{r}_{1},t_{1};\vec{r}_{2},t_{2};\vec{r}_{3},t_{3})$ (99)
$=$ $\langle|A_{11,22,33}+A_{11,23,31}+A_{12,21,33}$
$+A_{12,23,31}+A_{13,22,31}+A_{13,21,32}|^{2}\rangle$
$=$ $\langle|e^{i(\varphi_{1}+\varphi_{2}+\varphi_{3})}(K_{11}K_{22}K_{33}+K_{11}K_{23}K_{31}+K_{12}K_{21}K_{33}$
$+K_{12}K_{23}K_{31}+K_{13}K_{22}K_{31}+K_{13}K_{21}K_{32})|^{2}\rangle,$
where $\varphi_{1}$, $\varphi_{2}$, and $\varphi_{3}$ are the initial phases of photons emitted by S1, S2, and S3, respectively.
With the same method and approximations above, one-dimension spatial three-photon probability function is
$P^{(3)}(x_{1};x_{2};x_{3})$
$=$ $6+6\cos\frac{2\pi d_{12}}{L\lambda}(x_{1}-x_{2})$
$+$ $6\cos\frac{2\pi d_{13}}{L\lambda}(x_{1}-x_{3})+6\cos\frac{2\pi d_{23}}{L\lambda}(x_{2}-x_{3})$
$+$ $2\cos\frac{2\pi d_{13}}{L\lambda}x_{1}\cos\frac{2\pi d_{12}}{L\lambda}(x_{2}-\frac{d_{23}}{2})\cos\frac{2\pi d_{23}}{L\lambda}(x_{3}+\frac{d_{12}}{2})$
$+$ $2\cos\frac{2\pi d_{13}}{L\lambda}x_{1}\cos\frac{2\pi d_{12}}{L\lambda}(x_{3}-\frac{d_{23}}{2})\cos\frac{2\pi d_{23}}{L\lambda}(x_{2}+\frac{d_{12}}{2})$
$+$ $2\cos\frac{2\pi d_{13}}{L\lambda}x_{2}\cos\frac{2\pi d_{12}}{L\lambda}(x_{1}-\frac{d_{23}}{2})\cos\frac{2\pi d_{23}}{L\lambda}(x_{3}+\frac{d_{12}}{2})$
$+$ $2\cos\frac{2\pi d_{13}}{L\lambda}x_{2}\cos\frac{2\pi d_{12}}{L\lambda}(x_{3}-\frac{d_{23}}{2})\cos\frac{2\pi d_{23}}{L\lambda}(x_{1}+\frac{d_{12}}{2})$
$+$ $2\cos\frac{2\pi d_{13}}{L\lambda}x_{3}\cos\frac{2\pi d_{12}}{L\lambda}(x_{1}-\frac{d_{23}}{2})\cos\frac{2\pi d_{23}}{L\lambda}(x_{2}+\frac{d_{12}}{2})$
$+$ $2\cos\frac{2\pi d_{13}}{L\lambda}x_{3}\cos\frac{2\pi d_{12}}{L\lambda}(x_{2}-\frac{d_{23}}{2})\cos\frac{2\pi d_{23}}{L\lambda}(x_{1}+\frac{d_{12}}{2}).$
In order to see the third-order interference pattern more clearly, assuming $d_{12}=d_{23}=d$ and $x_{2}=x_{3}$, Eq. (V.3) can be simplified as
$P^{(3)}(x_{1};x_{2};x_{3}=x_{2})$
$=$ $1+\cos\frac{2\pi d}{L\lambda}(x_{1}-x_{2})$
$+\frac{1}{3}\cos\frac{4\pi d}{L\lambda}x_{1}\cos\frac{2\pi d}{L\lambda}(x_{2}-\frac{d}{2})\cos\frac{2\pi d}{L\lambda}(x_{2}+\frac{d}{2})$
$+\frac{1}{3}\cos\frac{4\pi d}{L\lambda}x_{2}\cos\frac{2\pi d}{L\lambda}(x_{1}-\frac{d}{2})\cos\frac{2\pi d}{L\lambda}(x_{2}+\frac{d}{2})$
$+\frac{1}{3}\cos\frac{4\pi d}{L\lambda}x_{2}\cos\frac{2\pi d}{L\lambda}(x_{2}-\frac{d}{2})\cos\frac{2\pi d}{L\lambda}(x_{1}+\frac{d}{2}).$
Figure 32 shows the third-order interference pattern given by Eq. (V.3).
Figure 32: Third-order interference pattern of three independent single-photon states.
Similar method can be employed to calculate the third-order interference other three types of light beams and higher-order interference of light.
VI Quantum atomic coherence theory based on Feynman’s path integral
The history of atomic coherence traces back to de Broglie in 1924 when he proposed matter wave assumption De Broglie (1924) by analogy of Planck’s Planck (1901) and Einstein’s Einstein (1905) quanta assumption of light. The de Broglie wavelength of matter wave equals
$\lambda_{D}=h/p.$ (102)
De Broglie’s assumption was soon verified by Davisson et al. and Thomson et al. with diffraction pattern of electrons Davisson and Germer (1927); Thomson and Reid (1927). Due to the very short de Broglie wavelength of matter wave, it is difficult to observe the first-order interference pattern like the one of light. In 1953, Marton et al. observed first-order interference pattern of electrons in a Mach-Zehnder interferometer Marton et al. (1953). Then the first-order interference pattern of neutron Maier-Leibnitz and Springer (1962), atom Keith et al. (1991), molecule Arndt et al. (1999) were also reported. The first-order interference pattern of two independent Bose Einstein Condensations (BEC) was also reported Andrews et al. (1997).
The first second-order interference experiments with matter wave was done by Yasuda and Shimizu, in which they observed two-atom bunching with cold atoms Yasuda and Shimizu (1996). Burt et al. observed the degree of third-order coherence of cold atoms and BEC, finding that the ratio between these two equals $7.4\pm 2.6$ Burt et al. (1997). The second-order interference of fermions was also reported by several groups Henny et al. (1999); Oliver et al. (1999); Iannuzzi et al. (2006); Rom et al. (2006); Bocquillon et al. (2013).
Quantum atomic coherence theory is mainly based on Glauber’s quantum optical coherence theory Adams et al. (1994); Meystre (2001); Cronin et al. (2009). Feynman himself introduced how to employ Fenman’s path integral to calculate the first-order interference of electrons in a Young’s double-slit interferometer Feynman et al. (2010, 2011). Storey and Cohen-Tannoudji provided a good introduction about how to employ Feynman’s path integral to calculate the first-order interference of atoms Storey and Cohen-Tannoudji (1994). Employing Feynman’s path integral to calculate the first-order interference of matter wave can be found in Refs. Deng (2006, 2008); Gondran and Gondran (2005), too.
This section will summarize how to employ Feynman’s path integral to calculate the first-, second, and higher-order interference matter waves, with special attentions on the analogy between cold atoms (BEC) and thermal light (laser light).
VI.1 First-order interference of matter wave
VI.1.1 First-order interference of one matter wave beam
Figure 33 shows the scheme for the first-order interference of one matter wave beam. S is a particle source. P1 and P2 are two pinholes with equal size. O is the observation plane. For simplicity, free atoms are assumed.
Figure 33: Scheme for the first-order interference of one matter wave beam.
There are two different paths to trigger a particle detection event in the observation plane. The first one is the particle going through P1. The probability is 1/2 and the probability amplitude is $A_{S1x}$. The second path is the particle going through P2. The probability is 1/2 and the probability amplitude is $A_{S2x}$. If these two different paths are indistinguishable, the probability function in the observation plane is
$P^{(1)}(x,t)$ $=$ $\langle|A_{S1x}+A_{S2x}|^{2}\rangle$ (103)
$=$ $\langle|e^{i\varphi_{j}}K_{S1x}+e^{i\varphi_{j}}K_{S2x}|^{2}\rangle$
$=$ $\langle|K_{S1}K_{1x}+K_{S2}K_{2x}|^{2}\rangle,$
where $\varphi_{j}$ is the initial phase of the $j$th detected particle, $K_{S1x}$ is the particle’s Feynman propagator from S to $x$ in the observation plane via P1, the meanings of other symbols are defined similarly.
Particle’s Feynman propagator can be calculated in the same method as the one for photon and has simpler form for free particles Feynman et al. (2010)
$K(x_{a},t_{a};x_{b},t_{b})=\sqrt{\frac{m}{2\pi i\hbar(t_{b}-t_{a})}}\exp\frac{im(x_{b}-x_{a})^{2}}{2\hbar(t_{b}-t_{a})},$ (104)
where $m$ is the mass of particle, $\hbar=h/(2\pi)$, $(x_{a},t_{a})$ and $(x_{b},t_{b})$ are the starting and ending space-time coordinates, respectively. For simplicity, one-dimension is considered.
Assuming S is in the symmetrical position of P1 and P2, there is no path difference from S to two pinholes. The path difference is introduced by the pinholes to the observation plane. Substituting Eq. (104) into Eq. (103), it is straightforward to have one-dimension spatial first-order interference pattern in the Fraunhofer range,
$P^{(1)}(x)\propto 1+\cos\frac{2\pi d}{L\lambda_{D}}x,$ (105)
in which paraxial approximation has been employed, $\lambda_{D}=h/(mv)$ is the de Broglie wavelength of particle, $v$ is the speed of atom and the speed of all atoms are assumed to be the same, $d$ is the distance between two pinholes, and $L$ is the distance between the pinhole and detection planes.
There is first-order interference pattern when these two different paths are indistinguishable for particles, no matter these particles are fermions or bosons, in cold atom state or BEC state. Just like there is no difference for the first-order interference pattern of photons in a Young’s double-slit interferometer for thermal and laser light. The initial phases of particles does not contribute to the first-order interference pattern of one beam.
VI.1.2 First-order interference of two independent matter wave beams
Figure 34 shows the scheme for the first-order interference of two independent matter wave beams. S1 and S2 are two independent particle sources. O is the observation plane.
Figure 34: Scheme for the first-order interference of two independent matter wave beams.
There are two different paths to trigger a particle detection event in the observation plane. The first one is the detected particle is emitted S1. The probability is 1/2 and the probability amplitude is $A_{1x}$. The second one is the detected particle is emitted S2. The probability is 1/2 and the probability amplitude is $A_{2x}$. If these two different paths are indistinguishable, the probability function in the observation plane is
$P^{(1)}(x,t)$ $=$ $\langle|A_{1x}+A_{2x}|^{2}\rangle$ (106)
$=$ $\langle|e^{i\varphi_{1j}}K_{1x}+e^{i\varphi_{2j}}K_{2x}|^{2}\rangle,$
where $\varphi_{1j}$ and $\varphi_{2j}$ are the initial phases of particles emitted by S1 and S2, respectively. With the same approximations above, spatial probability function can be obtained via Eq. (106),
$P^{(1)}(x)$ $=$ $\langle 1+\cos[\frac{2\pi d}{L\lambda_{D}}x+(\varphi_{1j}-\varphi_{2j})]\rangle.$ (107)
By analogy of the experimental results of the degree of second- and third-order coherence of cold atoms and BEC to thermal and laser light Yasuda and Shimizu (1996); Burt et al. (1997); Lopes et al. (2015), it is reasonable to assume the coherence properties of cold atoms are similar as the ones of thermal light and the coherence properties of BEC are similar as the one of laser light. The initial phases of particles in cold atoms are random. The initial phases of particles in one BEC are identical.
For the superposition of two independent cold atom sources, the initial phases of particles are random. Equation (107) is simplified as
$P^{(1)}(x)$ $\propto$ $1+\frac{1}{\sqrt{N}}\cos[\frac{2\pi d}{L\lambda_{D}}x+\varphi],$ (108)
where $N$ is the number of detected particles and $\varphi$ is a random phase due to the sum of $\varphi_{1j}-\varphi_{2j}$ for $N$ detected particles. The first-order interference pattern can not observed.
For the superposition of of two independent BECs, the initial phases of particles in one BEC are identical. Equation (107) is simplified as
$P^{(1)}(x)$ $=$ $1+\cos[\frac{2\pi d}{L\lambda_{D}}x+(\varphi_{1j}-\varphi_{2j})].$ (109)
$\varphi_{1j}-\varphi_{2j}$ is a random but fixed number for two independent BECs. The first-order interference pattern can be observed Andrews et al. (1997), just like the transient first-order interference pattern by superposing two independent laser light beams Magyar and Mandel (1963).
Similar method can be employed to calculate the first-order interference of multiple independent matter wave beams.
VI.2 Second-order interference of matter wave
VI.2.1 Second-order interference of one matter wave beam
The scheme for the second-order interference of one matter wave beam is similar as the one shown in Fig. 11 except the source S is a particle source, BS is a beam splitter for particles and these two detectors are particle detectors.
First discuss the second-order interference of one boson particle beam. Assuming S is a cold atom source, there are two different paths to trigger a two-particle coincidence count. The first one is particle a is detected by D1 and particle b is detected by D2. The probability is 1/2 and the corresponding two-particle amplitude is $A_{a1,b2}$. The second path is particle a is detected by D2 and particle b is detected by D1. The probability is 1/2 and the corresponding two-particle amplitude is $A_{a2,b1}$. If these two different paths are indistinguishable, the two-particle probability function is
$P^{(2)}(\vec{r}_{1},t_{1};\vec{r}_{2},t_{2})$ (110)
$=$ $\langle|A_{a1,b2}+A_{a2,b1}|^{2}\rangle$
$=$ $\langle|e^{i\varphi_{a}}K_{a1}e^{i\varphi_{b}}K_{b2}+e^{i\varphi_{a}}K_{a2}e^{i\varphi_{b}}K_{b1}|^{2}\rangle$
$=$ $\langle|K_{a1}K_{b2}+K_{a2}K_{b1}|^{2}\rangle.$
Substituting particle’s Feynman propagator into Eq. (110), the spatial two-particle probability function is
$P^{(2)}(x_{1}-x_{2})\propto 1+\text{sinc}^{2}\frac{\pi D}{L\lambda_{D}}(x_{1}-x_{2}),$ (111)
where $D$ is the length of S and single speed particles are assumed. The temporal two-particle probability function is
$P^{(2)}(t_{1}-t_{2})\propto 1+\text{sinc}^{2}\frac{\Delta\omega}{2}(t_{1}-t_{2}),$ (112)
where $\Delta\omega=2mv_{0}\Delta v/\hbar$ is the angular frequency bandwidth, $v_{0}$ is the central speed, $\Delta v$ is the speed bandwidth. Point source is assumed to simplify the calculation.
When these two detectors are at the same space-time coordinates, the probability to detect two particles is two times of the one when these two detectors are in different coherence volumes. There is two-particle bunching for cold boson particles in a HBT interferometer Yasuda and Shimizu (1996), just like there is two-photon bunching for thermal light in a HBT interferometer.
Assume S is a BEC in Fig. 11, there is only one path to trigger a two-particle coincidence count. Since the initial phases of particles in a BEC are identical, just like the one for single-mode continuous-wave laser light. No two-particle bunching can be observed, which is consistent with Burt et al.’s experiments Burt et al. (1997).
Then discuss the second-order interference of one fermion particle beam. There are also two different ways to trigger a two-particle coincidence count and the corresponding two-particle probability amplitude are $A_{a1,b2}$ and $A_{a2,b1}$. If these two different ways are indistinguishable, the two-particle probability function is
$P^{(2)}(\vec{r}_{1},t_{1};\vec{r}_{2},t_{2})$ (113)
$=$ $\langle|A_{a1,b2}-A_{a2,b1}|^{2}\rangle$
$=$ $\langle|e^{i\varphi_{a}}K_{a1}e^{i\varphi_{b}}K_{b2}-e^{i\varphi_{a}}K_{a2}e^{i\varphi_{b}}K_{b1}|^{2}\rangle$
$=$ $\langle|K_{a1}K_{b2}-K_{a2}K_{b1}|^{2}\rangle,$
where the minus sign is due to fermions having exchange antisymmetry Griffiths and Schroeter (2018). By analogy of cold boson particles in a HBT interferometer, one-dimension spatial two-particle probability function is
$P^{(2)}(x_{1}-x_{2})\propto 1-\text{sinc}^{2}\frac{\pi D}{L\lambda_{D}}(x_{1}-x_{2}),$ (114)
where $D$ is the length of S and single speed is assumed. The temporal two-particle probability function of fermions in a HBT interferometer is
$P^{(2)}(t_{1}-t_{2})\propto 1-\text{sinc}^{2}\frac{\Delta\omega}{2}(t_{1}-t_{2}),$ (115)
where the meanings of symbols are similar as the ones in Eq. (112).
When two detectors are at the same space-time coordinates, the probability to detect two fermions equals 0. As the distance between these two detectors increases, the probability to detect two fermions increases to 1. There is two-particle antibunching effect for fermion particles in a HBT interferometer Henny et al. (1999); Jeltes et al. (2007), in contrast to the two-particle bunching effect for bosons in a HBT interferometer Jeltes et al. (2007); Brown and Twiss (1956b, a).
VI.2.2 Second-order interference of two independent matter wave beams
Figure 18 shows the scheme for the second-order interference of two independent matter wave beams. S1 and S2 are two independent particle sources. BS is a particle beam splitter. D1 and D2 are two single-particle detectors. CC is two-particle coincidence count detection system.
First discuss the second-order interference of two single-particle sources. Assuming S1 and S2 are two identical single-particle sources, there are two different paths to trigger a two-particle coincidence count. The first one is the particle emitted by S1 is detected by D1 and the particle emitted by S2 is detected by D2. The probability is 1/2 and the two-particle probability amplitude is $A_{11,22}$. The second path is the particle emitted by S1 is detected by D2 and the particle emitted by S2 is detected by D1. The probability is 1/2 and the two-particle probability amplitude is $A_{12,21}$. If these different paths are indistinguishable, the two-particle probability function for bosons is
$P^{(2)}(\vec{r}_{1},t_{1};\vec{r}_{2},t_{2})$ (116)
$=$ $\langle|A_{11,22}+A_{12,21}|^{2}\rangle$
$=$ $\langle|e^{i\varphi_{1}}K_{11}e^{i\varphi_{2}}K_{22}+e^{i(\varphi_{1}+\frac{\pi}{2})}K_{12}e^{i(\varphi_{2}+\frac{\pi}{2})}K_{21}|^{2}\rangle$
$=$ $\langle|K_{11}K_{22}-K_{12}K_{21}|^{2}\rangle,$
Substituting Eq. (104) into Eq. (116), one-dimension temporal two-particle probability function is
$P^{(2)}(t_{1}-t_{2})\propto 1-\text{sinc}^{2}\frac{\Delta\omega}{2}(t_{1}-t_{2}),$ (117)
where point particle source is assumed. It is impossible to detect two boson particles when these two detectors are at symmetrical positions. HOM dip with visibility higher than 50% can be observed Lopes et al. (2015).
With the same method, one-dimension temporal two-particle probability function for fermions is
$P^{(2)}(t_{1}-t_{2})\propto 1+\text{sinc}^{2}\frac{\Delta\omega}{2}(t_{1}-t_{2}),$ (118)
in which two particles tends to leave BS from different output ports Bocquillon et al. (2013).
Then calculate the second-order interference of two independent and identical BECs in Fig. 18. There are three different ways to trigger a two-particle coincidence count. The first one is these two particles are both emitted by S1. The probability is 1/4 and there is one path, $A_{11,12}$. The second one is these two particles are both emitted by S2. The probability is 1/4 and there is one path, $A_{21,22}$. The third way is these two particles are emitted by S1 and S2 one each. The probability is 1/2 and there are two different paths, $A_{11,22}$ and $A_{12,21}$. When these four different paths are indistinguishable, the two-particle probability function is
$P^{(2)}(\vec{r}_{1},t_{1};\vec{r}_{2},t_{2})$ (119)
$=$ $\langle|\frac{1}{\sqrt{4}}A_{11,12}+\frac{1}{\sqrt{4}}A_{21,22}+\frac{1}{\sqrt{4}}(A_{11,22}+A_{12,21})|^{2}\rangle$
$\propto$ $\langle|e^{i(2\varphi_{1}+\frac{\pi}{2})}K_{11}K_{12}+e^{i((2\varphi_{2}+\frac{\pi}{2})}K_{21}K_{22}$
$+e^{i(\varphi_{1}+\varphi_{2})}(K_{11}K_{22}-K_{12}K_{21})|^{2}\rangle,$
where $\varphi_{1}$ and $\varphi_{2}$ are initial phases of particles emitted by S1 and S2, respectively. Since S1 and S2 are independent, Eq. (119) can be simplified as
$P^{(2)}(\vec{r}_{1},t_{1};\vec{r}_{2},t_{2})$
$\propto$ $\langle|K_{11}K_{12}|^{2}\rangle+\langle|K_{21}K_{22}|^{2}\rangle+\langle|(K_{11}K_{22}-K_{12}K_{21})|^{2}\rangle.$
Substituting Eq. (104) into Eq. (VI.2.2), it is straightforward to have temporal two-particle probability function
$P^{(2)}(t_{1}-t_{2})\propto 1-\frac{1}{2}\text{sinc}^{2}\frac{\Delta\omega}{2}(t_{1}-t_{2}).$ (121)
Compared to the second-order interference of two single boson sources, Eq. (117), the visibility of the second-order interference of two independent BECs can not exceed 50%, which is similar as the conclusion for the visibility of the second-order interference of two independent classical and nonclassical light beams Mandel (1983).
VI.3 Third-order interference of matter wave
Based on the results above, it is easy to predict that there is three-particle bunching when a cold boson particle beams in a third-order HBT interferometer and no three-particle bunching for a BEC Burt et al. (1997). This section will concentrate on the third-order interference of fermions. The scheme for the third-order interference of one fermion beam is the same as the one shown in Fig. 27, in which the meanings of all the symbols are similar as the one for photons. There are six different ways to trigger a three-particle coincidence count and the corresponding three-particle probability amplitudes are $A_{a1,b2,c3}$, $A_{a1,b3,c2}$, $A_{a2,b1,c3}$, $A_{a2,b3,c1}$, $A_{a3,b1,c2}$ and $A_{a3,b2,c1}$. When these six different paths are indistinguishable, the three-particle probability function is
$P^{(3)}(\vec{r}_{1},t_{1};\vec{r}_{2},t_{2};\vec{r}_{3},t_{3})$ (122)
$=$ $\langle|A_{a1,b2,c3}-A_{a1,b3,c2}-A_{a2,b1,c3}$
$+A_{a2,b3,c1}+A_{a3,b1,c2}-A_{a3,b2,c1}|^{2}\rangle,$
where the minus signs are due to exchange antisymmetry of fermions. With similar method above, temporal three-particle probability function is
$P^{(3)}(t_{1};t_{2};t_{3})$
$\propto$ $1-\text{sinc}^{2}\frac{\Delta\omega}{2}(t_{1}-t_{2})$
$-\text{sinc}^{2}\frac{\Delta\omega}{2}(t_{2}-t_{3})-\text{sinc}^{2}\frac{\Delta\omega}{2}(t_{3}-t_{1})$
$+2\text{sinc}\frac{\Delta\omega}{2}(t_{1}-t_{2})\text{sinc}\frac{\Delta\omega}{2}(t_{2}-t_{3})\text{sinc}\frac{\Delta\omega}{2}(t_{3}-t_{1}).$
Figure 35: Third-order temporal coherence function of fermions in a third-order HBT interferometer.
Figure 35 shows the normalized third-order temporal coherence function of fermions in a third-order HBT interferometer, which is proportional to Eq. (VI.3). The probability to detect three particles equals 0 when these three detectors are at the symmetrical positions, which is consistent with the results obtained by Liu et al. Liu (2016).
Similar method can be employed to calculate the third- and higher-order interference of multiple matter wave beams.
VII Summary
In summary, we have presented a new quantum optical coherence theory based on Feynman’s path integral, which provides a useful tool for studying the physics of coherence besides classical optical coherence theory based on Maxwell’s electromagnetic theory and Glauber’s quantum optical coherence theory based on wave mechanics formalism of quantum mechanics. Some interesting conclusions are drawn. For instance, unlike the well-accepted conclusion that the electric field models are the same for thermal and laser light within the coherence time, it is predicted by quantum optical coherence theory based on Feynman’s path integral that there may be some difference between these two. The physics behind two-photon bunching of thermal light in a HBT interferometer and HOM dip of entangled photon pairs in a HOM interferometer is the same, which can be interpreted by two-photon interference. In order to show how quantum optical coherence theory based on Feynman’s path integral can help us to understand the physics of optical coherence better, subwavelength interference is taken as an example to be illustrated and it is shown that there is no intrinsic difference between the observed second-order subwavelength interference for two detectors scanning in the same direction and opposite directions. Finally, quantum particle coherence theory based on Feynman’s path integral is also given by analogy of the one in optics.
We believe quantum optical coherence theory based on Feynman’s path integral will be helpful to understand the physics of optical and atomic coherence in other fields, and eventually lead us to the answer to the question: what is photon?
Acknowledgments
Jianbin Liu would like to thank the help from Guoquan Zhang and Yanhua Shih. This project is supported by National Science Foundation of China (No.11404255), Doctoral Fund of Ministry of Education of China (No.20130201120013), Shanxi Key Research and Development Project (No. 2019ZDLGY09-08), Xi’an Science and Technology Program Project (No. GX2331), Open fund of MOE Key Laboratory of Weak-Light Nonlinear Photonics (No. OS19-2), and Fundamental Research Funds for the Central Universities.

### v2

@@
 Abstract
-The first, second and higher-order interference of classical and nonclassical light is interpreted in quantum optical coherence theory based on Feynman’s path integral.
-Compared to the existed classical optical coherence theory based on Maxwell’s electromagnetic theory and Glauber’s quantum optical coherence theory based on wave mechanics formulation of quantum mechanics, quantum optical coherence theory based on Feynman’s path integral provides a new tool to study optical coherence and has the advantage of understanding the connection between the physics and mathematical calculations.
-For instance, based on the results of transient first-order interference of two independent light beams, it is predicted that the well-accepted electric field model for thermal light introduced by many optical textbooks may not be accurate.
+Compared to classical optical coherence theory based on Maxwell’s electromagnetic theory and Glauber’s quantum optical coherence theory based on matrix mechanics formulation of quantum mechanics, quantum optical coherence theory based on Feynman’s path integral formulation of quantum mechanics provides a novel tool to study optical coherence.
+It has the advantage of understanding the connection between mathematical calculations and physical interpretations better.
+Quantum optical coherence theory based on Feynman’s path integral is introduced and reviewed in this paper.
+Based on the results of transient first-order interference of two independent light beams, it is predicted that the classical model for electric field of thermal light introduced by classical optical textbooks may not be accurate.
 The physics of two-photon bunching of thermal light and Hong-Ou-Mandel dip of entangled photon pairs is the same, which can be interpreted by constructive and destructive two-photon interference, respectively.
-The second-order interference of two independent light beams are introduced in detail.
-Interpreting the interference of light in quantum optical coherence theory based on Feynman’s path integral is helpful to understand the physics of light, and may eventually lead to the answer of the question puzzling us for a long time: what is photon?
+Quantum optical coherence theory based on Feynman’s path integral is helpful to understand the coherence properties of light, which may eventually lead us to the answer of the question: what is a photon?
 I Introduction
-Ever since the interference of light introduced by Young in 1801 Young (1802), optical coherence has been one of the most important topics in the development of physics.
-With the contributions of generations of scientists, there are two main theories of optical coherence Born and Wolf (2013); Mandel and Wolf (1995).
+Ever since the interference of light was introduced by Young in 1801 Young (1802), optical coherence has been one of the most important topics in the development of physics.
+With contributions from generations of scientists, two main optical coherence theories have been established Born and Wolf (2013); Mandel and Wolf (1995).
 One is classical optical coherence theory based on Maxwell’s electromagnetic theory Born and Wolf (2013).
-The other one is Glauber’s quantum optical coherence theory based on wave mechanics formulation of quantum mechanics Glauber (1963a, c).
+The other one is Glauber’s quantum optical coherence theory based on matrix mechanics formulation of quantum mechanics Glauber (1963a, c).
 It was proved by Glauber and Sudarshan independently that classical and quantum optical coherence theories are mathematically equivalent for classical light Sudarshan (1963); Glauber (1963b).
@@
 (1977), Bell inequality violation Bell (1964, 1966) and so on.
-On the other hand, quantum optical coherence theory can interpret both classical and non-classical phenomena.
-Since Glauber’s quantum optical coherence theory can explain all the optical coherence phenomena and be a great success Glauber (2006), what is the reason to introduce another quantum optical coherence theory based on Feynman’s path integral?
-This is the question we asked ourselves since the first day we employed Feynman’s path integral to explain the subwavelength interference of light Liu and Zhang (2010).
-Similar question bothered Young and Feynman Young (1802); Feynman (1948), too.
-In the beginning of Young’s paper on the theory of light and colours Young (1802), he explained why he employed wave theory to interpret light and colours: Although the invention of plausible hypotheses, independent of any connection with experimental observations, can be of very little use in the promotion of natural knowledge; yet the discovery of simple and uniform principles, by which a great number of apparently heterogeneous phenomena are reduced to coherent and universal laws, must ever be allowed to be of considerable importance towards the improvement of the human intellect.
-When Feynman introduced path integral formulation of quantum mechanics, he also faced the same question.
-There had been wave mechanics and matrix mechanics formulations of quantum mechanics.
-What is the reason to introduce a third and equivalent formulation of quantum mechanics.
-Feynman’s answer to this question is Feynman (1948): The formulation is mathematically equivalent to the more usual formulations.
+On the other hand, Glauber’s quantum optical coherence theory can interpret both classical and non-classical phenomena.
+As Glauber’s quantum optical coherence theory can explain all the optical coherence phenomena Glauber (2006), why introduce another quantum optical coherence theory based on Feynman’s path integral?
+This is the question we asked ourselves when we employed Feynman’s path integral to explain the subwavelength interference of light Liu and Zhang (2010).
+Similar question bothered Young Young (1802) and Feynman Feynman (1948), too.
+At the beginning of Young’s paper on the theory of light and colours Young (1802), he explained why he employed wave theory: Although the invention of plausible hypotheses, independent of any connection with experimental observations, can be of very little use in the promotion of natural knowledge; yet the discovery of simple and uniform principles, by which a great number of apparently heterogeneous phenomena are reduced to coherent and universal laws, must ever be allowed to be of considerable importance towards the improvement of the human intellect.
+When Feynman introduced path integral formulation of quantum mechanics, he also faced similar question.
+There were wave mechanics and matrix mechanics formulations of quantum mechanics.
+Why introduce a third and equivalent formulation of quantum mechanics?
+Feynman’s answer to the question is Feynman (1948): The formulation is mathematically equivalent to the more usual formulations.
 There are, therefore, no fundamentally new results.
@@
 Also, there are problems for which the new point of view offers a distinct advantage.
-The answers offered by Young and Feynman are also the reasons why we introduced quantum optical coherence theory based on Feynman’s path integral besides Glauber’s quantum optical coherence theory.
+The answers given by Young and Feynman are also the reasons why we introduced quantum optical coherence theory based on Feynman’s path integral.
 As introduced in Sect.
-III, quantum optical coherence theory based on Feynman’s path integral can indeed offer deeper insight about the first- and second-order interference of two independent light beams, which is helpful to understand the physics of coherence and the concept of photon.
+III, quantum optical coherence theory based on Feynman’s path integral can indeed offer deeper insight about the physics of the first- and second-order interference of two independent light beams, which is essential to understand the nature of light.
 The remaining parts of the paper are organized as follows.
-Basic concepts in quantum optical coherence theory based on Feynman’s path integral are introduced in Sect.
+A short history of quantum optical coherence theory based on Feynman’s path integral and some basic concepts are introduced in Sect.
 II.
-The first- and second-order interference of light are introduced in Sects.
+Applying path integral to interpret the first- and second-order interference of light are in Sects.
 III and IV, respectively.
@@
 V.
-Section VI introduces the quantum atomic coherence theory based on Feynman’s path integral by analogy of the optical one.
-Section VII summarizes the conclusions.
+Section VI introduces quantum atomic coherence theory based on Feynman’s path integral by analogy of quantum optical coherence theory based on Feynman’s path integral.
+Our conclusions are summarized in Sect.
+VII.
 II Basic concepts in quantum optical coherence theory based on Feynman’s path integral
 II.1 History of quantum optical coherence theory based on Feynman’s path integral
-The history of optical coherence began in the 17th century when Grimaldi discovered diffraction of light Hecht (2012).
+The history of classical optical coherence theory began in the 17th century when Grimaldi discovered diffraction of light Hecht (2012).
 With the contributions from Young Young (1802, 1804), Fresnel Fresnel (1821), and many other scientists Hecht (2012); Born and Wolf (2013), classical optical coherence theory has been established based on Maxwell’s electromagnetic theory Maxwell (1865).
-We will not introduce the history of optical coherence since there are many good reviews about this topic Born and Wolf (2013); Hecht (2012); Mandel and Wolf (1965).
-The history of quantum optical coherence theory dates back to Dirac and Feynman.
-In his famous book on quantum mechanics, Dirac discussed the first-order interference of light with the concept of photon and concluded “Each photon then interferes only with itself.
-Interferences between two different photons never occurs Dirac (1930).” Feynman employed path integral to discuss Young’s double-slit interference of light in his lectures Feynman et al.
+The detail history of classical optical coherence can be found in Refs.
+Born and Wolf (2013); Hecht (2012); Mandel and Wolf (1965).
+The history of quantum optical coherence theory dates back to Dirac and Feynman Dirac (1930); Feynman et al.
 (2011).
-All the optical coherence discussions were limited to the first-order before 1956.
+In his famous book on quantum mechanics, Dirac discussed the first-order interference of light with the concept of photon and concluded that Each photon then interferes only with itself.
+Interferences between two different photons never occurs Dirac (1930).
+Feynman employed path integral to discuss the first-order interference of electrons in a Young’s double-slit interferometer in his lectures Feynman et al.
+(2011).
+All the discussions about optical coherence were limited to the first-order before 1956.
 Both quantum and classical optical coherence theories can explain the first-order interference of light.
-The need for quantum optical coherence theory was not so urgent.
+The need for quantum optical coherence theory was not so urgent then.
 In 1956, Hanbury Brown and Twiss found that photons emitted by thermal light source are not independent Brown and Twiss (1956a, b), which is known as two-photon bunching of thermal light.
-People began to realize that there are second- and higher-order interference of light besides the first-order.
+People began to realize that there are second- and higher-order interference of light besides the first-order interference.
 In 1963, Glauber published a series of papers explaining optical coherence in quantum theory Glauber (1963a, c, b), which laid the foundation of quantum optical coherence theory.
 Glauber was awarded 2005 Nobel physics prize for his contribution to quantum theory of optical coherence Glauber (2006).
-Glauber’s quantum optical coherence theory was employed by most scientists and introduced in many quantum optics books Mandel and Wolf (1995); Scully and Zubairy (1997); Loudon (2000); Shih (2020).
-Glauber’s quantum optical coherence theory is based on wave mechanics formulation of quantum mechanics Schrödinger (1926).
-There are two more equivalent formulations of quantum mechanics called matrix mechanics Heisenberg (1925, 1949) and Feynman’s path integral Feynman (1948); Feynman et al.
+Glauber’s quantum optical coherence theory was employed by most scientists in quantum optics and introduced in many quantum optics textbooks Mandel and Wolf (1995); Scully and Zubairy (1997); Loudon (2000); Shih (2020).
+Glauber’s quantum optical coherence theory is based on matrix mechanics formulation of quantum mechanics Heisenberg (1925, 1949).
+There are two more equivalent formulations of quantum mechanics called wave mechanics Schrödinger (1926) and Feynman’s path integral Feynman (1948); Feynman et al.
 (2010).
-In principle, one can establish quantum optical coherence theory based on matrix mechanics and Feynman’s path integral formulations, too.
-In the following parts of this paper, we will introduce quantum optical coherence theory based on Feynman’s path integral.
+In principle, one can establish quantum optical coherence theory based on wave mechanics and Feynman’s path integral formulations, too.
+In the following parts, we will introduce quantum optical coherence theory based on Feynman’s path integral.
 Quantum optical coherence theory based on Feynman’s path integral was first introduced by Feynman himself.
-In his lectures on physics, Feynman examined Young’s double-slit interference of electrons, which is similar as the Young’s double-slit interference of light Feynman et al.
+In his lectures on physics, Feynman examined Young’s double-slit interference of electrons, which is similar as the Young’s double-slit interference of photons Feynman et al.
 (2011).
@@
 As shown in Fig.
-1, there are two different ways for two photons emitted by atoms a and b and absorbed by atoms c and d.
-One is photon emitted by atom a is absorbed by atom c and photon emitted by atom b is absorbed by atom d.
-The second ways is emitted by atom a is absorbed by atom d and photon emitted by atom b is absorbed by atom c.
-Two-photon bunching of thermal light is interpreted as the superposition of these two different two-photon probability amplitudes corresponding these two different ways, which is later named as two-photon interference Shih (2020).
+1, there are two different and indistinguishable paths for two photons emitted by atoms a and b to be absorbed by atoms c and d.
+The first path is the photon emitted by atom a is absorbed by atom c and the photon emitted by atom b is absorbed by atom d.
+The second path is the photon emitted by atom a is absorbed by atom d and the photon emitted by atom b is absorbed by atom c.
+Two-photon bunching of thermal light is interpreted as the superposition of two different two-photon probability amplitudes corresponding to these two different paths, which is later named as two-photon interference Shih (2020).
 In his lecture on quantum electrodynamics, Feynman also employed similar idea to interpret two-photon bunching of thermal light Feynman (2006).
-Figure 1: Interpretation of two-photon bunching of thermal light in Feynman’s path integral. From Fano (1961)
-As mentioned before, Glauber’s quantum optical coherence theory was employed to study coherence in most researches in quantum optics.
-There are limited number of studies employing Feynman’s path integral to study optical coherence.
-Hillery and Zubairy employed Feynman’s path integral to discuss some problems in nonlinear optics and calculated the path-integral representation for the coherent-state propagator Hillery and Zubairy (1982).
-Yabuki employed Feynman’s path integral to discuss the nonlinear contribution in Young’s double-slit interferometer of light Yabuki (1986).
-Barut and Basri discussed single-photon interference in single- and double-slit interferometers with Feynman’s path integral Barut and Basri (1992).
+Figure 1: Interpretation of two-photon bunching of thermal light in Feynman’s path integral Fano (1961).
+Atoms a and b are two atoms that can emit a photon each.
+Atoms c and d are two atom that can absorb a photon each.
+$u$ and $v$ are two paths.
+As mentioned before, Glauber’s quantum optical coherence theory was employed to study optical coherence in most researches in quantum optics.
+There are limited number of researches employing Feynman’s path integral to study optical coherence.
+Hillery and Zubairy employed Feynman’s path integral to discuss some problems in nonlinear optics and calculated the path integral representation for coherent state propagator Hillery and Zubairy (1982).
+Yabuki employed Feynman’s path integral to discuss the nonlinear contribution in the first-order interference of light in a Young’s double-slit interferometer Yabuki (1986).
+Barut and Basri discussed single-photon interference in the single- and double-slit interferometers with Feynman’s path integral Barut and Basri (1992).
 Braidotti et.
 al.
-employed Feynman’s path integral to discuss non-paraxial optics problem Braidotti et al.
+employed Feynman’s path integral to discuss some non-paraxial optics problems Braidotti et al.
 (2018).
+Stöhr showed that Feynman’s path integral reduces to the conventional wave formalism of light diffraction in the lowest order Stöhr (2019).
 Wallner employed Feynman’s path integral to discuss the interference of entangled photon pairs Wallner (2021).
-We have employed employed Feynman’s path integral to discuss the first- and second-order interference of light since 2010 Liu and Zhang (2010, 2011); Liu et al.
+Wen et al.
+experimentally measured Feynman’s propagator for single photons Wen et al.
+(2023).
+We have employed Feynman’s path integral to discuss the first- and second-order interference of light since 2010 and obtained some interesting conclusions Liu and Zhang (2010, 2011); Liu et al.
 (2013, 2014b, 2014a, 2015b, 2015a); Bai et al.
@@
 (2021a, b).
-Some interesting conclusions have been drawn.
 For instance, it is predicted that there is no transient first-order interference pattern by superposing two independent thermal light beams Bai et al.
-(2016), which is different from the observed transient first-order interference pattern by superposing two independent laser light beams Javan et al.
+(2016), which is different from the one of superposing two independent laser light beams Javan et al.
 (1962); Magyar and Mandel (1963); Lipsett and Mandel (1963).
-This distinction leads to an interesting conclusion that electric field models for laser and classical light within the coherence time should be different, which is contrary to the usual electric model introduced in classical optics textbooks Born and Wolf (2013).
+The difference leads to an interesting conclusion that classical electric field models for laser and thermal light within the coherence time should be different, which is contrary to the well-accepted electric model introduced in classical optics textbooks Born and Wolf (2013).
 Detail discussions about this problem can be found in Sect.
 III.
-The researches on optical coherence with Feynman’s path integral did by other groups Fano (1961); Feynman et al.
+The researches on optical coherence with Feynman’s path integral performed by other groups Fano (1961); Feynman et al.
 (2011); Feynman (2006); Hillery and Zubairy (1982); Yabuki (1986); Barut and Basri (1992); Braidotti et al.
@@
 (2017, 2019a); Luo et al.
-(2021a, b) are all about special interference phenomena of light.
+(2021a, b) are all about certain specified interference phenomena of light.
 No systematic study about quantum optical coherence theory based on Feynman’s path integral is reported.
-This paper will present a comprehensive introduction about quantum optical coherence theory based on Feynman’s path integral by re-visiting the first-, second-, third- and higher-order interference of light.
+This paper presents a systematic introduction and review about quantum optical coherence theory based on Feynman’s path integral.
 II.2 Principles of quantum optical coherence theory based on Feynman’s path integral
@@
 The interference is lost: $P=P_{1}+P_{2}$
-In order to calculate the probability of an event with the above assumptions, there are two premises need to be figured out.
-The first one is how to calculate the probability amplitude $\phi$ for each alternative way.
-The second one is how to decide whether two different alternative ways for an event are distinguishable or not.
+There are two premises for calculating the probability of an event with the above assumptions.
+The first premise is how to calculate the probability amplitude, $\phi$, for each path.
+The second premise is how to decide two different alternatives for an event are distinguishable or not.
 II.2.1 Probability amplitude for a path
-The probability amplitude for a path is first given by Dirac by analogy of classical theory of dynamics Dirac (1930). Dirac assumed that the solution of schrödinger wave equation is as follows,
+The probability amplitude for a path is first given by Dirac by analogy of action in classical dynamics theory Dirac (1930). Dirac assumed that the solution of schrödinger wave equation is
 $\langle q_{t_{b}}|q_{t_{a}}\rangle=e^{iS/\hbar},$ (1)
@@
 Equation (1) means that the probability amplitude for a particle starting from $q_{t_{a}}$ at time $t_{a}$ and ending at $q_{t_{b}}$ at time $t_{b}$ equals $e^{iS/\hbar}$.
-Based on the above assumption, Dirac concluded that the probability amplitude can be expressed as a result of integral over all the possible paths
+Based on the above assumption, Dirac concluded that the probability amplitude of the path can be expressed as the integral of all the possible paths
 […]
@@
 where $t_{j}=t_{a}+(j-1)\times(t_{b}-t_{a})/m$ ($j=1,2,...,m$) and $q_{t_{j}}$ is the corresponding canonical coordinate at time $t_{j}$.
-Equation (2) is in fact path integral.
-The same equation can be found in Feynman’s book Feynman et al.
+Equation (2) is in fact the key result for path integral.
+The exactly same equation can be found in Feynman’s book Feynman et al.
 (2010).
-The difference between Dirac and Feynma’s work on path integral is that Dirac gave the probability amplitude by analogy of classical dynamics and he used the word, analogous to, while Feynman showed that the probability amplitude in quantum mechanics is equal to Eq.
+The difference between Dirac and Feynma’s work on path integral is that Dirac gave the probability amplitude by analogy of classical dynamics and he used the word, analogous to, while Feynman proved that the probability amplitude in quantum mechanics is equal to Eq.
 (1) Feynman et al.
 (2010).
-The probability amplitude for photon, usually called photon’s Feynman propagator, has simple form in Feynman’s Gauge in momentum space Greiner and Reinhardt (2013)
+The probability amplitude for photons, usually called photon’s Feynman propagator, has a simple form in momentum space in Feynman’s Gauge Greiner and Reinhardt (2013)
 $D_{F}^{\mu\nu}(k)=\frac{-g_{\mu\nu}}{k^{2}+i\epsilon},$ (3)
-where $g^{\mu\nu}$ is metric tensor in four-dimensional Minkowski spacetime, $k$ is four-dimension wave vector, $(k_{0},k_{1},k_{2},k_{3})$, $k_{0}=\omega$, $(k_{1},k_{2},k_{3})$ are three components of usual wave vector, $k^{2}=k^{\mu}k_{\mu}=k_{0}^{2}-k_{1}^{2}-k_{2}^{2}-k_{3}^{2}$, $\epsilon$ is an infinitesimal quantity introduced to ensure convergence of the integral.
+where $g^{\mu\nu}$ is the metric tensor in four-dimensional Minkowski spacetime, $k$ is a four-dimension wave vector, $(k_{0},k_{1},k_{2},k_{3})$, $k_{0}=\omega$, $(k_{1},k_{2},k_{3})$ are three components of usual wave vector, $k^{2}=k^{\mu}k_{\mu}=k_{0}^{2}-k_{1}^{2}-k_{2}^{2}-k_{3}^{2}$, $\epsilon$ is an infinitesimal quantity introduced to ensure convergence of the integral.
 Photon’s Feynman propagator in coordinate space can be obtained via Fourier transform of Eq. (3)
 $D_{F}^{\mu\nu}(x;y)=\int\frac{d^{4}k}{(2\pi)^{4}}e^{-ik\cdot(y-x)}D_{F}^{\mu\nu}(k),$ (4)
-where $x$ and $y$ are four-dimensional Minkowski coordinate vector, $x=(x_{0},x_{1},x_{2},x_{3})$, $x_{0}$ is time $t$, $(x_{1},x_{2},x_{3})$ is the three components of usual coordinate vector $\vec{x}$.
+where $x$ and $y$ are four-dimensional Minkowski coordinate vectors, $x=(x_{0},x_{1},x_{2},x_{3})$, $x_{0}$ is time $t$, $(x_{1},x_{2},x_{3})$ is the three components of the usual coordinate vector $\vec{x}$.
 Equation (4) can be further simplified as Greiner and Reinhardt (2013)
-$D_{F}(x;y)=\int\frac{d^{3}k}{(2\pi)^{3}}\frac{i}{(2\omega_{k})^{2}}e^{-i[\omega(t_{y}-t_{x})-\vec{k}\cdot(\vec{y}-\vec{x})]},$ (5)
+$D_{F}(x;y)=\int\frac{d^{3}k}{(2\pi)^{3}}\frac{i}{(2\omega)^{2}}e^{-i[\omega(t_{y}-t_{x})-\vec{k}\cdot(\vec{y}-\vec{x})]},$ (5)
 in which $t_{y}>t_{x}$ is assumed.
-Comparing photon’s Feynman propagator Eq. (5) with Green function in classical optics Born and Wolf (2013)
+Comparing photon’s Feynman propagator, Eq. (5), with Green function in classical optics Born and Wolf (2013)
 […]
-these two functions have similar expression with different contribution for each momentum component.
+it is easy to find out that these two functions have similar expressions with different contributions for each momentum component.
 The reason is that all the paths are taken into account in Eq.
@@
 $=$ $a(t)\dot{x}^{2}+b(t)\dot{x}x+c(t)x^{2}+d(t)\dot{x}+e(t)x+f(t),$
-where one dimension $x$ is assumed for simplicity. The particle’s Feynman propagator can be expressed as
+the particle’s Feynman propagator can be expressed as
 $K(x_{a},x_{b})=e^{iS_{cl}(x_{a},x_{b})/\hbar}F(t_{a},t_{b}),$ (8)
-where $x_{a}$ and $x_{b}$ are the starting and ending positions of the particle, respectively, $S_{cl}(a,b)$ is the classical action of the particle between $x_{a}$ and $x_{b}$, $F(t_{a},t_{b})$ is a function of the starting and ending time.
-Equation (8) means that the probability amplitude of particle going from $x_{a}$ to $x_{b}$ can be related to the classical propagation function with an extra function only dependent on the starting and ending time.
-The Lagrangian of free photon has the form of Eq.
+where one dimension case is assumed for simplicity, $x_{a}$ and $x_{b}$ are the starting and ending positions of the particle, respectively, $S_{cl}(a,b)$ is the classical action of the particle, $F(t_{a},t_{b})$ is a function of the starting and ending time.
+Equation (8) means that the particle’s Feynman propagator going from $x_{a}$ to $x_{b}$ can be related to the classical propagation function with an extra function which is only dependent on the starting and ending time.
+The Lagrangian of free photons has the form of Eq.
 (7) Srednicki (2007).
@@
 II.2.2 Distinguishable and indistinguishable paths
-Indistinguishable paths is a unique concept in quantum theory.
-All the paths in classical theory are distinguishable.
-One reason is that measurement are different in quantum and classical theories Wheeler and Zurek (2014).
+Indistinguishable paths are a unique concept in quantum theory.
+All the paths are distinguishable in classical theory.
+One reason is that measurements are different in quantum and classical theories Wheeler and Zurek (2014).
 In classical theory, measurement precision can be arbitrarily high (at least in principle) without introducing unpredictable influence to the measured system.
-In quantum theory, measurement precision can not be arbitrarily high due to the measurement will introduce unpredictable influence on the measured system.
-Another reason is that there is no identical particles in classical theory.
-As Griffiths and Schroeter mentioned in their book that all electrons are utterly identical, in a way that no two classical objects can never be, identical particles is a unique concept in quantum theory Griffiths and Schroeter (2018).
-Figure 2: Distinguishable and indistinguishable paths. S1 and S2 are two sources can emit quantum or classical particles. D is a single-particle detector.
-Taking Fig.
+In quantum theory, Measurement precision cannot be arbitrarily high because measurements will introduce unpredictable influence and the uncertainty of the measurement is limited by Heisenberg uncertainty principle.
+Another reason is that there are no identical particles in classical theory.
+As Griffiths and Schroeter mention in their book that all electrons are utterly identical, in a way that no two classical objects can never be, identical particle is a unique concept in quantum theory Griffiths and Schroeter (2018).
+Figure 2: Distinguishable and indistinguishable paths. S1 and S2 are two sources which can emit quantum or classical particles. D is a single-particle detector.
+Let us take the scheme shown in Fig.
 2 as an example to illustrate distinguishable and indistinguishable paths.
@@
 The particle comes from either S1 or S2.
-In classical theory, the particles emitted by these two sources are always distinguishable.
+In classical theory, the particles emitted by these two sources are distinguishable.
 Even if all the properties such as mass, volume, color and so on, are identical [^1] This assumption is impossible in classical theory.
 For instance, two particles having the same weight in classical theory is only valid in some precision.
-Since classical particles are infinitely divisible, one can always find higher precision in which these two particles having different weights., the momentum difference between these two cases can be employed to distinguish these two paths.
+Classical particles are infinitely divisible.
+One can always find higher precision in which these two particles having different weights., the momentum difference between these two cases can be employed to distinguish these two paths.
 Different paths are always distinguishable in classical theory.
 Things become different in quantum theory.
-Assuming the particles emitted by S1 and S2 are identical.
-The only difference between these two paths is the momentum difference.
+Assuming the particles emitted by S1 and S2 are identical, the only difference between these two paths is the momentum difference.
 The measurement precision of momentum in quantum theory is limited by Heisenberg uncertainty principle Heisenberg (1949)
@@
 The uncertainty of position is usually determined by the detection mechanism of the employed detector.
-For instance, the position uncertainty of photon detection is usually at the size of atom absorbing the photons via photoelectric effect, which is usually at $10^{-10}$ m range for metal atoms Shackelford (2016).
+Photon is usually detected by photoelectric effect.
+The position uncertainty of photon detection equals the size of the atom absorbing the photon, which is usually at $10^{-10}$ m range for metal atoms Shackelford (2016).
 The paths with momentum difference less than $h/10^{-10}$ are indistinguishable.
 Different paths can be indistinguishable in quantum theory.
-Heisenberg uncertainty principle should be employed to analyze the path difference when detail measuring apparatus is taken into account.
-II.3 What is photon?
+Heisenberg uncertainty principle is employed to analyze the path difference when detail measuring apparatus is taken into account.
+II.3 What is a photon?
 The concept of photon can trace back to Planck Plank (1901) and Einstein Einstein (1905).
-The answer to the question of what is photon is still not satisfactory.
+The answer to the question of what is a photon is still not satisfactory.
 However, it does not stop photons appear everywhere even when there is no need for them Schleich (2003).
@@
 In quantum optical coherence theory based on Feynman’s path integral, the concept of photon is unavoidable.
-It is impossible for us to answer the question of what is photon right now.
-However, it is helpful to clarify what we are talking about when the concept photon is employed.
-We think photon is a particle, which is described by photon wave function $\psi(\vec{r},t)$.
+It is impossible for us to answer the question of what is a photon right now.
+However, it is helpful to clarify what we are talking about when the concept of photon is employed.
+In short, we picture photon as a particle, which is described by photon wave function $\psi(\vec{r},t)$.
 The probability to find a photon at $(\vec{r},t)$ in a volume $dV$ is given by $|\psi(\vec{r},t)|^{2}dV$.
@@
 Wave-particle duality is well-accepted for photon.
-However, when we talk about photon in this paper, we try to picture it as a particle like electron, proton and so on.
+When we talk about photon in this paper, we picture it as a particle like electron, proton and so on.
 The particle is described by a photon wave function $\psi(\vec{r},t)$, which is different from the concept of particle in classical theory.
-The evidences for that photon is a particle are photoelectric effect Hertz (1887) and photon antibunching effect Paul (1982).
-Hertz observed the following phenomena when shining light on the surface of a metal:
+The evidence that photon is a particle includes photoelectric effect Hertz (1887) and photon antibunching effect Paul (1982).
+Hertz observed the following phenomena when shining a light beam on the surface of a metal:
 (I) The kinetic energy $E$ of the ejected electron is equal to $\hbar\nu-\Phi$, where $\nu$ is the frequency of incidental light, and $\Phi$ is the work function of the metal.
 (II) The rate of electron ejection is proportional to the square of the incident electric filed.
-(III) The time delay $T$ between the light shining on the metal and the ejected photoelectron is very short.
+(III) The time delay $T$ between the light shining on the metal and the ejected photo-electron is very short.
 It is experimentally proved by Lawrence and Beams that $T$ is shorter than 3 ns Lawrence and Beams (1928).
@@
 The first two phenomena can be interpreted by wave theory.
-However, the last phenomenon can only be interpreted by treating photon is a particle Muthukrishnan et al.
+The last phenomenon can only be interpreted by treating photon is a particle Muthukrishnan et al.
 (2017).
-The second evidence for that photon is a particle is photon antibuching effect Paul (1982).
-The physics of photon antibunching is that a photon can never trigger two separated single-photon detectors simultaneously.
+The second evidence for that photon is particle is photon antibuching effect Paul (1982).
+The physics of photon antibunching is that one photon cannot trigger two separated single-photon detectors simultaneously.
 The key to observe photon antibunching is to generate single-photon state.
@@
 There are also other evidence for photon is a particle.
-For instance, when a beam of light with low intensity is incident to a interferometer, the first-order interference pattern is accumulated point by point on the observation plane Dimitrova and Weis (2008).
+For instance, when a light beam with low intensity is incident into an interferometer, the first-order interference pattern builds up point by point on the observation plane Dimitrova and Weis (2008).
 II.3.2 Photon’s wave function
-Strictly speaking, there is no wave function since for photon there is no position operator for photon Bohm (2012).
-However, one can define photon’s position operator from the point view of detection.
+Strictly speaking, there is no wave function for photon since there is no position operator for photon Bohm (2012).
+However, one can define photon’s position operator from the point of view of detection.
 Photon is usually detected by photoelectric effect.
 The position resolution is limited by the size of atom.
-As long as the position accuracy of photon’s wave function is not arbitrary high, one can define and use photon’s wave function by analogy of the wave function of massive particle.
+As long as the position accuracy of photon’s wave function is lower than the size of atom, one can define and use photon’s wave function by analogy of the wave function of a massive particle.
 The wave function of a single photon emitted by a two-level atom is Scully and Zubairy (1997)
 $\psi(\vec{r},t)=K\frac{\sin\eta}{r}\theta(t-\frac{r}{c})\exp{[-i(\omega+i\frac{\Gamma}{2})(t-\frac{r}{c})]},$ (10)
-where $K$ is a normalization constant, $r=|\vec{r}-\vec{r}_{0}|$ is the distance between the position of the atom $\vec{r}_{0}$ and the position of photon $\vec{r}$, $\eta$ is the angle of $\vec{r}$ comparing to the electric dipole of atom, $\theta(t-\frac{r}{c})$ is a step function, which means that photon can only be detected when photon propagates at position $\vec{r}$, $\omega$ is the frequency of photon, $\Gamma$ is the width of up-level energy of two-level system, $\exp{[-i(\omega+i\frac{\Gamma}{2})(t-\frac{r}{c})]}$ means that the probability to detect the photon is decayed by $e^{-\Gamma t}$.
-Equation (10) only gives the temporal part of photon’s wave function.
+where $K$ is a normalization constant, $r=|\vec{r}-\vec{r}_{0}|$ is the distance between the position of the atom, $\vec{r}_{0}$, and the position of photon, $\vec{r}$, $\eta$ is the angle of $\vec{r}$ related to the electric dipole of atom, $\theta(t-\frac{r}{c})$ is a step function, which indicates that a photon can be detected at $\vec{r}$ only when it propagates to position $\vec{r}$, $\omega$ is the frequency of photon, $\Gamma$ is the width of the up-level energy of two-level system, $\exp{[-i(\omega+i\frac{\Gamma}{2})(t-\frac{r}{c})]}$ indicates that the probability of detecting the photon decays by $e^{-\Gamma t}$.
+Equation (10) gives the temporal part of photon’s wave function.
 The spatial part of photon’s wave function can be treated the same as the spatial structure of classical electric field Scully and Zubairy (1997).
@@
 II.3.3 Photon’s phase
-Photon’s phase is defined by analogy of the phase of electric field in classical electromagnetic theory Dirac (1930). Classical electromagnetic field can be expressed as Born and Wolf (2013)
+Photon’s phase is defined by analogy of the phase of electric field in classical electromagnetic theory Dirac (1930). Classical electric field can be expressed as Born and Wolf (2013)
 $\vec{E}(\vec{r},t)=\vec{E}_{0}\exp{[-i(\omega t-\vec{k}\cdot\vec{r})+i\varphi_{0}]},$ (11)
@@
 When $r$ goes to infinity, the discrete phase becomes continuous.
-The uncertainties of phase ($\Delta\varphi$) and number of photons ($\Delta N$) satisfies the following uncertainty principle Pegg and Barnett (1997)
+The uncertainties of phase ($\Delta\varphi$) and number of photons ($\Delta N$) satisfies the following uncertainty principle Pegg and Barnett (1997),
 $\Delta N\Delta\varphi\geq\frac{1}{2}.$ (13)
-For number state $|N\rangle$, the uncertainty of photon number is zero, which means that the phase of photon in number state is complete uncertain.
+For number state, $|N\rangle$, the uncertainty of photon number is zero, which means that the phases of photons in a number state are completely uncertain.
 It can be proved that the phase distribution function,
 $P(\varphi)=\frac{1}{2\pi}|\sum_{0}^{\infty}c_{n}\exp{(-in\varphi)}|^{2},$ (14)
-of single-photon state equals $1/2\pi$, which means the phase of photon in single-photon state is equally distributed between 0 and $2\pi$ Pegg and Barnett (1997).
-Feynman assumed that the initial phase of photon is determined by the emission time Feynman (2006).
-Based on the phase relationship between different photons, light can be categorized into two groups.
+of single-photon state equals $1/2\pi$, which means that the phase of the photon in a single-photon state is equally distributed between 0 and $2\pi$ Pegg and Barnett (1997).
+Feynman assumed that the initial phase of a photon is determined by the emission time Feynman (2006).
+Based on the phase relationship between different photons in a light beam, light can be categorized into two groups.
 The first group is that the initial phases of photons in a light beam are random.
@@
 The photons in this type of light are usually emitted by stimulated emission.
-Single-mode laser, amplified spontaneous emission belongs to the second type.
+Single-mode laser, amplified spontaneous emission belongs to the second type of light.
 II.4 Coherence volume and degeneracy factor of light
-The key to apply superposition principle in Feynman’s path integral to calculate optical coherence is distinguishing different paths for an event are distinguishable or not.
-Heisenberg’s uncertainty principle should be employed to judge whether different paths are distinguishable or not.
-Indistinguishable path is close related to indistinguishable photons.
+The key to applying superposition principle in Feynman’s path integral is to determine whether different paths for an event are distinguishable or not.
+Heisenberg uncertainty principle is employed to judge whether different paths are distinguishable or not.
+Indistinguishable paths are close related to indistinguishable particles.
 Photons within the same coherence volume are indistinguishable Martienssen and Spiller (1964).
-The uncertainties of coordinates ($\Delta x$, $\Delta y$, and $\Delta z$) and momentum ($\Delta p_{x}$, $\Delta p_{y}$, and $\Delta p_{z}$) of photons within the same coherence volume satisfies the following equation,
-$\Delta x\Delta p_{x}=h$
-$\Delta y\Delta p_{y}=h$ (15)
+The uncertainties of coordinates ($\Delta x$, $\Delta y$, and $\Delta z$) and momentum ($\Delta p_{x}$, $\Delta p_{y}$, and $\Delta p_{z}$) of photons within the same coherence volume satisfies the following equations,
+$\Delta x\Delta p_{x}=h,$
+$\Delta y\Delta p_{y}=h,$ (15)
 $\Delta z\Delta p_{z}=h.$
@@
 $\delta=\frac{1}{e^{h\nu/k_{B}T}-1},$ (16)
-where $h$ is Planck constant, $v$ is the frequency of radiation, $k_{B}$ is Boltzmann constant, $T$ is temperature of blackbody.
+where $h$ is the Planck constant, $\nu$ is the frequency of radiation, $k_{B}$ is the Boltzmann constant, $T$ is the temperature of blackbody.
+Thermal light can be treated as blackbody radiation.
 The degeneracy factor of usual thermal light is much less than 1.
@@
 On the other hand, the degeneracy factor of laser light can be much larger than 1.
-For instance, the degeneracy factor a 1-mW single-mode continuous-wave He-Ne laser light equals $3\times 10^{9}$ Mandel and Wolf (1995).
+For instance, the degeneracy factor a 1-mW single-mode continuous-wave He-Ne laser light beam equals $3\times 10^{9}$ Mandel and Wolf (1995).
 III First-order interference of light
 The first-order interference of light can be well interpreted in classical optical coherence theory based on Maxwell’s electromagnetic theory Born and Wolf (2013).
-Dirac interpreted the first-order interference of light in the language of photon and gave the well-known statement about the first-order interference of light: “Each photon then interferes only with itself.
+Dirac firstly interpreted the first-order interference of light in the language of photons and gave the well-known statement about the first-order interference of light: “Each photon then interferes only with itself.
 Interferences between two different photons never occurs Dirac (1930).” The contents in this section are mainly based on Dirac’s statement.
@@
 III.1 First-order interference of one light beam
-The first-order interference of one light beam means that light emitted by one source is split into two or more components and then combined together to observe interference.
-It corresponds to that each photon interferes only with itself in Dirac’s statement.
-Taking Mach-Zehnder interferometer shown in Fig.
+The first-order interference of one light beam means that light emitted by one source is split into two or more components and then combined together to observe the first-order interference pattern.
+It corresponds to each photon interferes only with itself in Dirac’s statement.
+Take Mach-Zehnder interferometer shown in Fig.
 3 as an example to show how to employ Feynman’s path integral to discuss the first-order interference of one light beam.
-Figure 3: Mach-Zehnder interferometer. S is light source. BS1 and BS2 are two non-polarizing 1:1 beam splitter. M1 and M2 are two mirrors. D is a single-photon detector.
-The event for the single-photon detector D detecting a photon in Fig.
-3 has two different paths.
-The first one is the detected photon taking the upper path (transmitting through beam splitter BS1, reflected by mirror M1, and reflected by beam splitter BS2) and recorded as path 1 for simplicity.
-The probability amplitude for photon taking path 1 is written as $A_{1}(S,D)$.
+Figure 3: Mach-Zehnder interferometer. S is a light source. BS1 and BS2 are two non-polarizing 1:1 beam splitter. M1 and M2 are two mirrors. D is a single-photon detector.
+There are two different paths for a single-photon detector, D, detecting a photon in Fig.
+3.
+The first path is the detected photon taking the upper path (transmitting through beam splitter BS1, reflected by mirror M1, and reflected by beam splitter BS2) and recorded as path 1 for simplicity.
+The probability for the photon taking the path is $1/2$ and the probability amplitude for the photon taking the path is $A_{1}(S,D)$.
 The second path is the photon taking the lower path (reflected by BS1 and M2, transmitting through BS2) and recorded as path 2.
-The probability amplitude for photon taking path 2 is written as $A_{2}(S,D)$.
+The probability for the photon taking the path is $1/2$ and the probability amplitude for photon taking the path is $A_{2}(S,D)$.
 If these two different paths to trigger a photon detection event are indistinguishable, the probability function for the $j$th detected photon is
-$P_{j}(t)$ $=$ $|A_{1}(S,D)+A_{2}(S,D)|^{2}$ (17)
-$=$ $|e^{i\varphi_{j}}K_{1}(S,D)+e^{i\varphi_{j}}K_{2}(S,D)|^{2},$
-where $j$ is a positive integer, $\varphi_{j}$ is the initial phase of the $j$th detected photon, $K_{1}(S,D)$ is photon’s Feynman propagator from the source to detector via path 1, $A_{1}(S,D)=e^{i\varphi_{j}}K_{1}(S,D)$, $\varphi_{j}$ is the initial phase of the $j$th detected photon.
+$P_{j}(t)$ $=$ $|\frac{1}{\sqrt{2}}A_{1}(S,D)+\frac{1}{\sqrt{2}}A_{2}(S,D)|^{2}$ (17)
+$\propto$ $|e^{i\varphi_{j}}K_{1}(S,D)+e^{i\varphi_{j}}K_{2}(S,D)|^{2},$
+where $j$ is a positive integer, $\varphi_{j}$ is the initial phase of the $j$th detected photon, $K_{1}(S,D)$ is photon’s Feynman propagator from the source to the detector via path 1, $A_{1}(S,D)=e^{i\varphi_{j}}K_{1}(S,D)$.
 Similar definition holds for $K_{2}(S,D)$.
 The first-order interference of one light beam is due to photon interferes with itself.
-The initial phases are the same for two different probability amplitudes.
-The first-order interference pattern is proportional to the sum of the probability functions of all the detected photon,
+The initial phase is identical for the same photon taking two different paths.
+The first-order interference pattern is proportional to the sum of the probability functions of all the detected photons,
 $I(t)\propto\sum_{j=1}^{N}P_{j}(t),$ (18)
-where $N$ is the number of detected photon. For a large enough number of photons,
+where $N$ is the number of the detected photons. For a large enough number of photons,
 $I(t)\propto\langle P_{j}(t)\rangle,$ (19)
-where delimited-⟨$\langle$..$\rangle$ … ⟩ is ensemble average by taking all possible realizations Mandel and Wolf (1995).
+where delimited-⟨$\langle$..$\rangle$ … ⟩ is ensemble average by taking all possible realizations Mandel and Wolf (1995); Shih (2020).
 Substituting Eq.
@@
 (19), the first-order interference pattern in Fig.
-3 is proportional to
+3 is
 $I(t)\propto\langle|K_{1}(S,D)+K_{2}(S,D)|^{2}\rangle.$ (20)
-The initial phase of photon, $\varphi_{j}$, does not contribute to the first-order interference pattern in Fig.
+The initial phase of the photon, $\varphi_{j}$, does not contribute to the first-order interference pattern in Fig.
 3.
-No matter whether thermal or laser light is employed, as long as the frequency bandwidth and spatial distribution are the same, the observed first-order interference pattern will be identical.
+Regardless of whether thermal or laser light is used, as long as the frequency bandwidth and spatial distribution are identical, the observed first-order interference patterns will be the same.
 The reason is that the first-order interference is a result of photon interferes with itself.
-The initial phases of different paths for the $j$th detected photon are identical.
-For a point light source, photon’s Feynman propagator can be approximated as
+For a point light source, the photon’s Feynman propagator can be approximated as
 $K(\vec{r}_{0},t_{0};\vec{r},t)\propto\frac{1}{|\vec{r}-\vec{r}_{0}|}e^{-i[\omega_{0}(t-t_{0})-\vec{k}\cdot(\vec{r}-\vec{r}_{0})]},$ (21)
-where $(\vec{r}_{0},t_{0})$ is the space-time coordinate of the photon being emitted, $(\vec{r},t)$ is the space-time coordinate of the photon being detected, $\omega_{0}$ is the angular frequency, $\vec{k}$ is wave vector.
-For simplicity, single-frequency is assumed in the following calculations.
+where $(\vec{r}_{0},t_{0})$ is the space-time coordinate of the photon being emitted, $(\vec{r},t)$ is the space-time coordinate of the photon being detected, $\omega_{0}$ is angular frequency, $\vec{k}$ is wave vector.
+For simplicity, single-frequency light is assumed firstly.
 Substituting Eq.
@@
 $\propto$ $1+\cos[\omega_{0}(t_{1}-t_{2})],$
-where $t_{1}$ and $t_{2}$ are the travailing time of photon from S to D via path 1 and 2, respectively.
+where $t_{1}$ and $t_{2}$ are the travailing time of the photon from S to D via path 1 and 2, respectively.
 The first-order interference pattern is proportional to the probability function of the $j$th detected photon in Fig.
@@
 where $f(\omega)$ is the frequency distribution function, $\hat{a}^{\dagger}(\omega)$ is the creation operator, $|0\rangle$ is vacuum state.
-When the time difference between the photon taking two paths, $|t_{1}-t_{2}|$, is larger than the coherence time of the employed single-photon state, the two different paths to trigger a photon detection event at D are distinguishable.
+When the time difference between the photon taking two paths, $|t_{1}-t_{2}|$, is longer than the coherence time of the employed single-photon state, the two different paths to trigger a photon detection event at D are distinguishable.
 The probability function for the $j$th detected photon equals
 $P_{j}(t)$
-$=$ $|e^{i\varphi_{j}}K_{1}(S,D)|^{2}+|e^{i\varphi_{j}}K_{2}(S,D)|^{2}$
+$=$ $|\frac{1}{\sqrt{2}}e^{i\varphi_{j}}K_{1}(S,D)|^{2}+|\frac{1}{\sqrt{2}}e^{i\varphi_{j}}K_{2}(S,D)|^{2}$
 […]
-where $t$ is the time of photon being detected by detector, $t_{1}=t-t_{01}$, $t_{01}$ is the emitting time of photon detected by D via path 1, $t_{2}=t-t_{02}$ and $t_{02}$ is the emitting time of photon detected by D via path 2.
-The probability of the $j$th detected photon is the sum of the probabilities of the photon taking two different paths.
+where $t$ is the time of photon being detected, $t_{1}=t-t_{01}$, $t_{01}$ is the emitting time of the photon detected by D via path 1, $t_{2}=t-t_{02}$ and $t_{02}$ is the emitting time of the photon detected by D via path 2.
+The probability function of the $j$th detected photon is the sum of the probability functions of the photon taking two different paths.
 There is no first-order interference.
-When the time difference, $|t_{1}-t_{2}|$, is less than the coherence time of the employed single-photon state, these two different paths to trigger a photon detection event at D are indistinguishable.
+When the time difference, $|t_{1}-t_{2}|$, is shorter than the coherence time of the employed single-photon state, these two different paths to trigger a photon detection event at D are indistinguishable.
 The probability function of the $j$th detected photon is
-$P_{j}(t)$ $=$ $|e^{i\varphi_{j}}K_{1}(S,D)+e^{i\varphi_{j}}K_{2}(S,D)|^{2}$
+$P_{j}(t)$ $=$ $|\frac{1}{\sqrt{2}}e^{i\varphi_{j}}K_{1}(S,D)+\frac{1}{\sqrt{2}}e^{i\varphi_{j}}K_{2}(S,D)|^{2}$
 […]
@@
 $f(\omega)=A\exp{[-\frac{(\omega-\omega_{0})^{2}}{2\sigma^{2}}]},$ (26)
-where $A$ is a normalization constant, $\sigma^{2}$ is variance of Gaussian function, $\omega_{0}$ is the central frequency of photon. Equation (III.1) can be simplified as Shih (2020)
+where $A$ is a normalization constant, $\sigma^{2}$ is variance of Gaussian function, $\omega_{0}$ is the central frequency. Equation (III.1) can be simplified as Shih (2020)
 $P_{j}(t_{1}-t_{2})=1+\frac{e^{-\sigma^{2}(t_{1}^{2}+t_{2}^{2})/2}}{e^{-\sigma^{2}t_{1}^{2}}+e^{-\sigma^{2}t_{2}^{2}}}\cos{\omega_{0}(t_{1}-t_{2})}.$ (27)
-The first-order interference pattern can not be observed with only one photon.
-The first-order interference pattern of single-photon state can be observed by preparing many identical single-photon states and put these single photon states into the interferometer one by one.
-By collecting large enough number of photons, the first-order interference pattern with single-photon source in Fig.
+The first-order interference pattern of light can not be observed with only one photon.
+The first-order interference pattern of a single-photon state can be observed by preparing many identical single-photon states and put these single-photon states into the interferometer one by one.
+By collecting large enough number of photons, the first-order interference pattern with single-photon states in Fig.
 3 is
@@
 $=$ $1+\frac{e^{-\sigma^{2}(t_{1}^{2}+t_{2}^{2})/2}}{e^{-\sigma^{2}t_{1}^{2}}+e^{-\sigma^{2}t_{2}^{2}}}\cos{\omega_{0}(t_{1}-t_{2})},$
-which is the same the probability of the $j$th detected photon.
-Figure 4: First-order interference pattern of single-photon states in Mach-Zehnder interferometer
-Figure 4 shows the first-order interference pattern of single-photon state with Gaussian spectrum in a Mach-Zehnder interferometer.
-$\tau=t_{1}-t_{2}$ is the time difference between path 1 and 2, I is the intensity of light.
+which is the same the probability function of the $j$th detected photon.
+Figure 4: First-order temporal interference pattern of single-photon states in a Mach-Zehnder interferometer.
+I is intensity of light and $\tau$ is time difference between photon taking these two paths.
+Figure 4 shows the simulated first-order interference pattern of single-photon states with Gaussian spectrum in a Mach-Zehnder interferometer.
+$\tau=t_{1}-t_{2}$ is the time difference between paths 1 and 2, I is the intensity of light.
 When the time difference equals 0, the visibility of the first-order interference pattern can reach 100%.
-As the time difference becomes larger, the visibility of first-order interference pattern decreases.
-When the time difference is larger than the coherence length of light, the visibility goes to zero, which is consistent with Eq.
+As the time difference becomes larger, the visibility of the first-order interference pattern decreases.
+When the time difference is larger than the coherence time of light, the visibility goes to zero, which is consistent with Eq.
 (III.1).
 The simulated result in Fig.
-4 is consistent with experimental results with single photons Braig et al.
+4 is consistent with the experimental results with single photons Braig et al.
 (2003)
-Similar method can be employed to calculate the spatial first-order interference pattern of one light beam in Feynman’s path integral.
+Similar method can be employed to calculate the first-order spatial interference pattern of one light beam in Feynman’s path integral.
 III.2 First-order interference of two independent light beams
 The first-order interference of one light beam confirms Dirac’s statement that photon only interferes with itself.
-However, based on Popper’s philosophy that a scientific conclusion can not be proved to be true Popper (2005).
-The first-order interference of light does not prove Dirac’s statement is correct.
+Based on Popper’s philosophy that a scientific conclusion can not be proved to be true Popper (2005).
+The first-order interference of one light beam does not prove Dirac’s statement is correct.
 On the contrary, any scientific conclusion should be possible to be disproved.
-Dirac’s statement can be disproved if one can observe the first-order interference pattern with two independent light beams.
-When Young first discussed interference of light, he pointed out that there is first-order interference pattern of two light beams from different origins by analogy of sound and water waves Young (1802): When two Undulations, from different Origins, coincide either perfectly or very nearly in Direction, their joint effect is a Combination of the Motions belonging to each.
-Later, Young realized that it is impossible to observe first-order interference pattern of light emitted by two different origins.
+Dirac’s statement may be disproved if one can observe the first-order interference pattern with two independent light beams.
+When Young first discussed the interference of light, he pointed out that there is first-order interference pattern of two light beams from different origins by analogy of sound waves Young (1802): When two Undulations, from different Origins, coincide either perfectly or very nearly in Direction, their joint effect is a Combination of the Motions belonging to each.
+Later, Young realized that it is impossible to observe first-order interference pattern of light from two different origins.
 He observed the first-order interference pattern of sunlight by splitting one beam into two and then combining these two beams together Young (1804).
-It is now well-accepted that in order to observe the first-order interference pattern with two light beams, these two light beams must have the polarization and frequency, and fixed relative phase Hecht (2012).
-The easiest way to meet these condition is these two light beams are splitting from the same one.
+It is now well-accepted that in order to observe the first-order interference pattern with two light beams, these two light beams must have the same polarization and frequency, and fixed relative phase Hecht (2012).
+The easiest way to meet these conditions is that these two light beams are splitting from the same one.
 Most of the first-order interferometers are designed in this way Born and Wolf (2013).
@@
 (1962); Magyar and Mandel (1963); Lipsett and Mandel (1963), which is usually called transient first-order interference pattern.
-Does the observed transient first-order interference pattern with two independent laser light beams mean that the second part of Dirac’s statement is disproved?
+Does the observed transient first-order interference pattern with two independent laser light beams means that the second part of Dirac’s statement is disproved?
 Further experiments shows that the first-order interference pattern exists when the intensities of these two laser light beams are so low that there is only one photon in the interferometer in most time Pfleegor and Mandel (1968); Radloff (1968); Hariharan et al.
 (1993).
-It is thought by Pfleegor and Mandel that the observed transient first-order interference pattern does not conflict with Dirac’s statement Pfleegor and Mandel (1967).
-It was the detection of the photon makes the photon into a superposition state and give interference Pfleegor and Mandel (1967).
+It is interpreted by Pfleegor and Mandel that the observed transient first-order interference pattern does not conflict with Dirac’s statement Pfleegor and Mandel (1967).
+It is the detection of a photon that puts the photon into a superposition state and causes interference Pfleegor and Mandel (1967).
 What Glauber said about these experiments is easier to understand Glauber (1995): The things that interfere in quantum mechanics are not particles.
@@
 III.2.1 First-order interference of two independent laser light beams
-In the first-order interference of two independent laser light beams shown in Fig.
-5, there are two different paths for a single-photon detector D detecting a photon.
-The first one is that the detected photon comes from laser S1 and the corresponding probability amplitude is $A_{1}(x)$.
-The second one is that the detected photon comes from laser S2 and the corresponding probability amplitude is $A_{2}(x)$.
-If these two laser light beams have the same polarization, frequency bandwidth, intensity, these two different ways can only be distinguished by momentum difference.
+In the scheme of the first-order interference of two independent laser light beams shown in Fig.
+5, there are two different paths for the single-photon detector, D, detecting a photon.
+The first path is that the detected photon comes from laser S1.
+The probability for the detected photon comes from S1 is $I_{1}/(I_{1}+I_{2})$, where $I_{1}$ and $I_{2}$ are the intensities of light emitted by S1 and S2, respectively.
+The corresponding probability amplitude is $A_{1}(x)$.
+The second one is that the detected photon comes from laser S2.
+The probability for the detected photon comes from S2 is $I_{2}/(I_{1}+I_{2})$ and the corresponding probability amplitude is $A_{2}(x)$.
+If these two laser light beams have the same polarization and frequency bandwidth, these two different paths can only be distinguished by momentum difference.
 As introduced in Sect.
@@
 $d\leq\frac{\lambda L}{\Delta x},$ (29)
-these two different paths are indistinguishable, where $L$ is the distance between source and observation planes, $\lambda$ is the wavelength of photon, and $\Delta x$ is the position uncertainty of photon detection.
-The probability function for the $j$th detected photon is
-$P_{j}(x)$ $=$ $|A_{j1}(x)+A_{j2}(x)|^{2}$ (30)
-$=$ $|e^{i\varphi_{j1}}K_{j1}(x)+e^{i\varphi_{j2}}K_{j2}(x)|^{2},$
+these two different paths are indistinguishable, where $L$ is the distance between source and observation planes, $\lambda$ is the wavelength of the photon, and $\Delta x$ is the position uncertainty of photon detection.
+The probability function of the $j$th detected photon is
+$P_{j}(x)$ $=$ $|\sqrt{\frac{I_{1}}{I_{1}+I_{2}}}A_{j1}(x)+\sqrt{\frac{I_{2}}{I_{1}+I_{2}}}A_{j2}(x)|^{2}$ (30)
+$\propto$ $|\sqrt{I_{1}}e^{i\varphi_{j1}}K_{j1}(x)+\sqrt{I_{2}}e^{i\varphi_{j2}}K_{j2}(x)|^{2},$
 where $\varphi_{j1}$ and $\varphi_{j2}$ are the initial phases of the $j$th detected photon emitted by S1 and S2, respectively, $K_{j1}(x)$ and $K_{j2}(x)$ are the corresponding photon’s Feynman propagators.
@@
 The distance between two sources is $d$.
-The distance between the source plane and the observation plane, O, equals $L$.
+The distance between the source and observation planes equals $L$.
+O is the observation plane.
 D is a single-photon detector.
-For a point light source, Eq. (30) can be simplified as
+For two point light sources with equal intensity, Eq. (30) can be simplified as
 $P_{j}(x)\propto 1+\cos[\frac{2\pi d}{\lambda L}x+(\varphi_{j1}-\varphi_{j2})],$ (31)
-where paraxial approximation and one-dimension are employed to simplify the calculation.
-The initial phases of photons in single-mode continuous-wave laser light are identical within the coherence volume.
+where paraxial approximation and one-dimension are employed to simplify the calculation Born and Wolf (2013).
+The initial phases of photons in a single-mode continuous-wave laser light beam are identical within one coherence volume.
 The initial phases of photons in different coherence volumes are random.
@@
 However, the relative phase, $\varphi_{j1}-\varphi_{j2}$, is fixed within one coherence volume for two independent lasers.
-The observed intensity is proportional to the sum of the probabilities of all the detected photons
+The observed intensity is proportional to the sum of the probability functions of all the detected photons
 $I(x)$ $\propto$ $\sum_{j}^{N}P_{j}(x)$ (32)
@@
 $I(x)$ $\propto$ $1,$ (34)
-where the first-order interference pattern can not be observed for a long observation time.
+where the first-order interference pattern of two independent laser light beams disappears for a long observation time.
 The results above are consistent with Magyar and Mandel’s experiments Magyar and Mandel (1963).
 Even through the physics of the interference of two independent lasers is in question for a long time Mandel (1964); Louradour et al.
-(1993); Wallace (1994); Davis and Parigger (1994); Paul (1986), the experimental results are certain and the predictions in quantum optical coherence theory is consistent with the one obtained in classical theory.
+(1993); Wallace (1994); Davis and Parigger (1994); Paul (1986), the experimental results are certain and the predictions in quantum optical coherence theory is consistent with the ones obtained in classical optical coherence theory.
 III.2.2 First-order interference of two independent thermal light beams
-Things become interesting for the first-order interference of two independent thermal light beams.
-As mentioned in Sect.
-II.3, there are two major differences between photons in laser and thermal light.
-The first one is that the initial phases of photons in thermal light are random within the coherence volume and the initial phases of photons in laser light are identical within one coherence volume.
-The second difference is that the degeneracy factor of thermal light is usually much less than 1 and the degeneracy factor of laser light can be much larger than one.
-The latter difference indicates that it may be impossible to detect large enough number of photons to see whether there is transient first-order interference pattern of two independent thermal light beams within the coherence time.
-No transient first-order interference pattern of two independent thermal light beams has been reported as the one for two independent lasers Magyar and Mandel (1963).
-Theoretically, we can assume the degeneracy factor of thermal light is much larger than one and discuss whether there exists transient first-order interference pattern or not.
-The scheme for the first-order interference of two independent thermal light beams is also represented by Fig.
+Things become different for the first-order interference of two independent thermal light beams.
+There are two major differences between photons in laser and thermal light.
+The first difference is that the initial phases of photons in thermal light are random within one coherence volume and the initial phases of photons in laser light are identical within one coherence volume.
+The second difference is that the degeneracy factor of thermal light is usually much less than one and the degeneracy factor of laser light can be much larger than one.
+The latter difference indicates that it may be impossible to detect large enough number of photons to observe whether there is transient first-order interference pattern of two independent thermal light beams within the coherence time or not.
+No experimental observation of the transient first-order interference pattern of two independent thermal light beams was reported as the one for two independent lasers Magyar and Mandel (1963).
+Theoretically, one can assume that the degeneracy factor of thermal light is much larger than one and discuss whether there exists transient first-order interference pattern or not.
+The scheme for the first-order interference of two independent thermal light beams is shown in Fig.
 5, where S1 and S2 are two identical and independent thermal light sources.
 There are two different paths for the single-photon detector D detecting a photon.
+Assume the intensities of these two light beams are equal and the same assumption is always employed in the following situations unless specified otherwise.
 With the same process as the one in Sect.
@@
 $P_{j}(x)\propto 1+\cos[\frac{2\pi d}{\lambda L}x+(\varphi_{j1}-\varphi_{j2})],$ (35)
-where paraxial approximation and one-dimension are assumed. The probability function for $N$ detected photon is
+where paraxial approximation and one-dimension are assumed. The probability function for $N$ detected photons is
 $P_{N}(x)$ $=$ $\sum_{j}^{N}1+\cos[\frac{2\pi d}{\lambda L}x+(\varphi_{j1}-\varphi_{j2})].$ (36)
@@
 $P_{N}(x)\propto 1+\frac{1}{\sqrt{N}}\cos(\frac{2\pi d}{\lambda L}x+\varphi),$ (37)
-where $\varphi$ is the result of the sum of large random phases. The visibility of the first-order interference pattern after detecting $N$ photon is
+where $\varphi$ is the result of the sum of large number of random phases. The visibility of the first-order interference pattern after detecting $N$ photon is
 $V_{N}=\frac{P_{Nmax}-P_{Nmin}}{P_{Nmax}+P_{Nmin}}=\frac{1}{\sqrt{N}},$ (38)
 which decreases as the number of detected photons increases.
-Figure 6 shows the relationship between the simulated visibility of first-order interference pattern of two independent thermal light beams and the number of detected photons.
+Figure 6 shows the relationship between the visibility of the simulated first-order interference pattern of two independent thermal light beams and the number of detected photons.
 $V$ is the visibility and $N$ is the number of detected photons.
-The red line is theoretical value given by Eq.
+The red solid line is theoretical value given by Eq.
 (38) and square dots are numerical results.
@@
 (2016).
-Figure 6: The visibility of first-order interference pattern of two independent thermal light beams vs. the number of detected photons Bai et al. (2016).
-The first-order interference pattern can not be observed with only one or several photons due to one photon only triggers a photo-electron Dimitrova and Weis (2008).
+Figure 6: The visibility of first-order interference pattern of two independent thermal light beams vs.
+the number of detected photons Bai et al.
+(2016).
+$V$ is the visibility of first-order interference pattern and $N$ is the number of detected photons.
+The first-order interference pattern of light can not be observed with only one or several photons due to one photon only triggers one photo-electron Dimitrova and Weis (2008).
 There must be large enough number of photons to retrieve the first-order interference pattern.
-However, the visibility decreases as the number of detected photon increases via Eq.
+However, the visibility decreases as the number of detected photons increases via Eq.
 (38).
-The first-order interference pattern of two independent thermal light beams may not be observable, which is similar as the first-order interference pattern of two independent laser light beams can not be observed for a long observation time.
+The first-order interference pattern of two independent thermal light beams may not be observable.
+The reason is similar as the one for the first-order interference pattern of two independent laser light beams can not be observed for a long observation time.
 III.2.3 First-order interference of two independent non-classical light beams
@@
 5 are two identical and independent single-photon sources.
-There are two different paths for single-photon detector D detecting a photon, which are emitted by S1 and S2.
+There are two different paths for single-photon detector detecting a photon.
+The detected photon is emitted by either S1 or S2.
 The probability amplitudes for these two paths are $A_{1}(x)$ and $A_{2}(x)$, respectively.
 When the distance between these two sources is less than $\lambda L/\Delta x$, these two difference paths can not be distinguished by momentum difference.
-However, this condition alone does not guarantee that these two paths are indistinguishable as the one in laser and thermal light cases.
+However, these conditions do not guarantee that these two paths are indistinguishable as the one in laser and thermal light cases.
 When there is only one photon in the interferometer, these two difference paths are always distinguishable.
@@
 The probability function for the $j$th detected photon is
-$P_{j}(x)$ $=$ $|A_{j1}(x)|^{2}+|A_{j2}(x)|^{2}$ (39)
-$=$ $|e^{i\varphi_{j1}}K_{j1}(x)|^{2}+|e^{i\varphi_{j2}}K_{j2}(x)|^{2}.$
-The probability equals the sum of the probabilities for different paths. There is no interference.
-When these two single-photon sources emit photons simultaneously, there are two photons in the interferometer and these two different paths are distinguishable when all the other parameters are indistinguishable.
+$P_{j}(x)$ $=$ $|\frac{1}{\sqrt{2}}A_{j1}(x)|^{2}+|\frac{1}{\sqrt{2}}A_{j2}(x)|^{2}$ (39)
+$\propto$ $|e^{i\varphi_{j1}}K_{j1}(x)|^{2}+|e^{i\varphi_{j2}}K_{j2}(x)|^{2}.$
+The probability function equals the sum of the probability functions for these two paths. There is no interference.
+When these two single-photon sources emit photons simultaneously, there are two photons in the interferometer and these two different paths can be indistinguishable when all the other conditions are satisfied.
 The probability function for the $j$th detected photon is
-$P_{j}^{\prime}{}(x)=|e^{i\varphi_{j1}}K_{j1}(x)+e^{i\varphi_{j2}}K_{j2}(x)|^{2}.$ (40)
-The probability function after detecting $N$ photons is the sum of the probabilities correspond to these two cases,
+$P_{j}^{\prime}{}(x)=|\frac{1}{\sqrt{2}}e^{i\varphi_{j1}}K_{j1}(x)+\frac{1}{\sqrt{2}}e^{i\varphi_{j2}}K_{j2}(x)|^{2}.$ (40)
+The probability function after detecting $N$ photons is the sum of the probability function corresponding to these two cases,
 $P_{N}(x)=\sum_{j=1}^{N_{1}}P_{j}(x)+\sum_{j=1}^{N_{2}}P_{j}^{\prime}{}(x),$ (41)
@@
 $P_{N}(x)\propto 1+\frac{N_{2}}{N}\frac{1}{\sqrt{N_{2}}}\cos(\frac{2\pi d}{\lambda L}x+\varphi).$ (42)
-$N_{2}$ is much less than $N$ for two independent single-photon sources.
+$N_{2}$ is much less than $N$ for the probability of two independent single-photon sources simultaneously emitting one photon each is much less than the probability of these two sources emitting one photon only.
 The visibility of the first-order interference pattern of two independent single-photon states is much less than the one of two-independent thermal light beams.
 It is impossible to observe the first-order interference pattern of two independent single-photon states for either short or long time comparing to the coherence time.
-With the same method above, one can calculate the first-order interference of two independent number states with different numbers, one thermal and one laser light beams, and other conditions.
+With the same method above, one can calculate the first-order interference of two independent number states with different numbers of photons, one thermal and one laser light beams, and so on.
 No transient first-order interference pattern can be observed except the one of two independent laser light beams.
-No first-order interference pattern can be observed for two independent light beams with long observing time.
-III.2.4 Remakes about the classical model of laser and thermal light field
-With the results above, it is interesting to re-visit the electric model for laser and thermal light in classical electromagnetic theory.
+No first-order interference pattern exists for two independent light beams with long observing time.
+III.2.4 Remakes about the classical models of laser and thermal light field
+With the results above, it is interesting to re-visit the electric field models for laser and thermal light in classical electromagnetic theory.
 For simplicity, quasi-monochromatic light and plane wave are assumed.
@@
 The same model as Eq.
-(43) is employed to describe thermal light field within one coherence time.
-For instance, in Born and Wolf’s book, they wrote: In a monochromatic wave field the amplitude of the vibrations at any point P is constant, while the phase varies linearly with time.
+(43) is employed to describe electric field of thermal light within one coherence time.
+For instance, in Born and Wolf’s book Born and Wolf (2013), they wrote: In a monochromatic wave field the amplitude of the vibrations at any point P is constant, while the phase varies linearly with time.
 This is no longer the case in a wave field produced by a real source: the amplitude and phase undergo irregular fluctuation, the rapidity of which depends essentially on the effective width $\Delta\nu$ of the spectrum.
@@
 The characteristic time $\Delta t=1/\Delta v$ is of the order of the coherence time introduced in §7.5.8.
-They thought the amplitude and phase of electric field produced by a real source undergoes irregular fluctuations for a long time interval.
-The amplitude and phase of electric field can be seen as constant within the coherence time.
+It means that the amplitude and phase of electric field produced by a real source undergo irregular fluctuations for a long time interval.
+The amplitude and phase of electric field can be seen as constant within one coherence time.
 In other words, the electric field produced by real source, i.e., laser or thermal light sources, can be described by Eq.
@@
 The same electric field model is employed to describe laser and thermal light field within the coherence time.
-This model is employed by Forrester to calculate the first-order interference of two independent thermal light beams, too Forrester (1956).
+This model was employed by Forrester to calculate the first-order interference of two independent thermal light beams, too Forrester (1956).
 In classical theory, it is straightforward to obtain the conclusion that there is transient first-order interference pattern by superposing two independent thermal light beams as the one of two independent laser light beams.
@@
 III.2.2.
-Experiments should be employed to verify the different conclusions.
-However, due to the very low degeneracy factor of thermal light, it is very difficult to observe transient first-order interference pattern of two independent thermal light beams directly.
+Experiments should be employed to verify these two different conclusions.
+However, due to the very low degeneracy factor of thermal light, it is difficult to observe transient first-order interference pattern of two independent thermal light beams directly.
 Forrester et al.
-claimed they observed interference between independently generated light waves Forrester et al.
+claimed that they observed interference between independently generated light waves Forrester et al.
 (1955).
-Detail analysis of their experiments indicates that their experiments is a type of one beam interference including nonlinear effect as the slit in a Young’s double-slit interferometer, which is similar as the ones with single atom Grangier et al.
+Detail analysis of their experiments indicates that their experiment belongs to one beam interference including nonlinear effect as the slits in a Young’s double-slit interferometer, which is similar as the ones with single atom Grangier et al.
 (1985); Eichmann et al.
 (1993) and atom ensembles Afzelius et al.
-(2007) as the slit in a Young’s double-slit interferometer.
-Forrester et al.’s experiments can not be interpreted as the first-order interference of two independent thermal light beams.
-If there is no transient first-order interference pattern with two independent thermal light beams, the classical model for thermal light within the coherence time may not be the same as the one of laser light.
-This result is helpful for us to understand the physics of thermal light and worthy of further experimental studies.
+(2007) as the slits in a Young’s double-slit interferometer.
+Forrester et al.’s experiments can not be interpreted as the results of first-order interference of two independent thermal light beams.
+If there is no transient first-order interference pattern with two independent thermal light beams, the classical model for electric field of thermal light within the coherence time may not be the same as the one of laser light.
+This result is helpful to understand the physics of thermal light and be worthy of further experimental studies.
 III.3 First-order interference of multiple independent light beams
 The same method can be employed to calculate the first-order interference of multiple independent light beams.
-Figure 7 shows the first-order interference of three light beams emitted by three independent light sources, S1, S2, and S3.
+Figure 7 shows the scheme for the first-order interference of three light beams emitted by three independent light sources, S1, S2, and S3.
 Figure 7: First-order interference of three independent light beams.
 S1, S2, and S3 are three independent light sources.
-D is a single-photon detector in the observation plane.
+D is a single-photon detector in the observation plane, O.
 $L$ is the distance between the source and observation planes.
-There are three different paths for the single-photon detector D detecting a photon.
-The first one is that the detected photon is emitted by S1 and the corresponding probability amplitude is $A_{1}(S,D))$.
-The second and third paths can be defined similarly and the corresponding probability amplitudes are $A_{2}(S,D)$ and $A_{3}(S,D)$, respectively.
+There are three different paths for the single-photon detector, D, detecting a photon.
+The first path is that the detected photon is emitted by S1 and the corresponding probability amplitude is $A_{1}(S,D)$.
+The second and third paths are defined similarly and the corresponding probability amplitudes are $A_{2}(S,D)$ and $A_{3}(S,D)$, respectively.
 For simplicity, these three light sources are assumed to be identical single-mode continuous-wave lasers with equal intensities.
-The polarizations and frequency distributions of the photons emitted by these three lasers are assumed to be the same.
+The polarizations and frequency bandwidths of the photons emitted by these three lasers are the same.
 When the distance between $S_{1}$ and S3 is less than $\lambda L/\Delta x$, these three different paths to trigger a photon detection event are indistinguishable.
@@
 $P_{j}(x)$
-$=$ $|A_{1}(S,D)+A_{2}(S,D)+A_{3}(S,D)|^{2}$
-$=$ $|e^{i\varphi_{j1}}K_{1}(S,D)+e^{i\varphi_{j2}}K_{2}(S,D)+e^{i\varphi_{j3}}K_{3}(S,D)|^{2},$
+$=$ $|\frac{1}{\sqrt{3}}[A_{1}(S,D)+A_{2}(S,D)+A_{3}(S,D)]|^{2}$
+$\propto$ $|e^{i\varphi_{j1}}K_{1}(S,D)+e^{i\varphi_{j2}}K_{2}(S,D)+e^{i\varphi_{j3}}K_{3}(S,D)|^{2},$
 where $\varphi_{j1}$, $\varphi_{j2}$ and $\varphi_{j3}$ are the initial phases of the $j$th detected photon emitted by S1, S2 and S3, respectively.
-$K_{1}(S,D)$, $K_{2}(S,D)$ and $K_{3}(S,D)$ are the corresponding photon’s Feyman propagators for these three paths.
-Equation (7) can be simplified as
+$K_{1}(S,D)$, $K_{2}(S,D)$ and $K_{3}(S,D)$ are the corresponding photon’s Feynman propagators for the three paths.
+Equation (III.3) can be simplified as
 $P_{j}(x)$ $=$ $\frac{3}{2}+\cos[\frac{2\pi d_{12}}{\lambda L}(x-\frac{d_{23}}{2})+(\varphi_{1j}-\varphi_{2j})]$ (45)
@@
 $+\cos[\frac{2\pi d_{13}}{\lambda L}x+(\varphi_{1j}-\varphi_{3j})],$
-where paraxial approximation and one-dimension are assumed to simplify the calculation, $d_{12}$ is the distance between S1 and S2, similar definitions hold for $d_{13}$ and $d_{23}$.
-The relative position for these three sources is shown in Fig.
+where paraxial approximation and one-dimension case are assumed to simplify the calculation, $d_{12}$ is the distance between S1 and S2, similar definitions hold for $d_{13}$ and $d_{23}$.
+The relative positions for the three sources are shown in Fig.
 8.
-Figure 8: Relative positions of three sources.
-Since these three light sources are independent lasers, $\varphi_{j1}$, $\varphi_{j2}$ and $\varphi_{j3}$ are random.
+Figure 8: Relative positions of three sources S1, S2 and S3.
+Since the three lasers are independent, $\varphi_{j1}$, $\varphi_{j2}$ and $\varphi_{j3}$ are random.
 The relative phases are fixed if all the detected photons are within one coherence volume.
-The probability for $N$ detected photon is
+The probability function for $N$ detected photons is
 $P_{N}(x)\propto P_{j}(x),$ (46)
-where transient first-order interference pattern between three independent laser light beams can be observed.
-Figure 9 shows transient first-order interference pattern of three laser light beams with equal initial phases.
-When $d_{12}\neq d_{23}$, the visibility of the transient first-order interference pattern can not reach 100%.
+where transient first-order interference pattern of three independent laser light beams can be observed.
+Figure 9 shows the transient first-order interference pattern of three laser light beams with equal initial phases.
+When $d_{12}\neq d_{23}$, the visibility of the transient first-order interference pattern can not reach 100% as shown in Fig.
+9(a).
 The visibility can reach 100% when $d_{12}=d_{23}$ as shown in Fig.
 9(b).
-Figure 9: Transient first-order interference pattern for (a) $d_{12}\neq d_{23}$ and (b) $d_{12}=d_{23}$.
-When the observation time is much longer than the coherence time, the probability for $N$ detected photons can be approximated as a constant due to the random initial phases.
+Figure 9: Transient first-order interference patterns (a) $d_{12}\neq d_{23}$ and (b) $d_{12}=d_{23}$.
+When the observation time is much longer than the coherence time, the probability function for $N$ detected photons is approximated to be a constant due to the sum of large number of random initial phases.
 The same method can be employed to calculate the first-order interference of three and more different types of independent light beams.
-It can be concluded that only when more than two laser light beams are employed, transient first-order interference pattern can be observed.
+It can be predicted that only when two or more laser light beams are employed, transient first-order interference pattern can be observed.
 IV Second-order interference of light
 As mentioned above, nearly all the studies of optical coherence are about the first-order interference of light before 1956 Born and Wolf (2013).
-In 1956, Hanbury Brow and Twiss discovered two-photon bunching of thermal light Brown and Twiss (1956b, a), which is one of the most famous second-order interference phenomena in quantum optics.
-The history of second-order coherence of light can trace back to 1940s when several groups measured the correlation of photon pair generated in the annihilation of a positron-electron pair Hanna (1948); Bleuler and Bradt (1948); Wu and Shaknov (1950).
-The second-order interference of light is measured with coincidence count detection system, which is proposed by Bothe in 1925 Bothe and Geiger (1925).
-The second-order interference of light has an unique role in the development of optical coherence theory.
+In 1956, Hanbury Brown and Twiss discovered two-photon bunching of thermal light Brown and Twiss (1956b, a), which is regarded as one of the cornerstones of modern quantum optics.
+The history of second-order coherence of light can trace back to 1940s when several groups measured the correlation of photon pair generated in the annihilation of positron-electron pairs Hanna (1948); Bleuler and Bradt (1948); Wu and Shaknov (1950).
+The key element to measure the second-order interference of light is coincidence count detection system, which is proposed by Bothe in 1925 Bothe and Geiger (1925).
+The second-order interference of light plays an unique role in the history of the development of optical coherence.
 The first-order interference of light exists in nature.
 For instance, soap bubbles, transparent wings of insects, some butterfly wings, etc.
-appear colored streaks when exposed to sunlight is due to the first-order interference of light.
+appearing colored streaks when exposed to sunlight is due to the first-order interference of light.
 The second-order interference of light does not exist in nature.
-It is a human invented phenomenon.
+It is a human-invented phenomenon.
 Special detection scheme is required to observe the second-order interference of light.
-The second-order coherence is qualitatively different from the first-order interference.
-It may be reason why people took more than 200 years since the first observed phenomenon of first-order interference of light to the first observed phenomenon of second-order interference of light.
-Extending the second-oder coherence to the third- and higher-order coherence is natural.
-The third- and higher-order coherence is quantitative different from the second-order coherence.
-Figure 10: (a) First-, (b) second- and (c) third-order interference.
-S is light source.
-IT1, IT2, IT3 are first-, second- and third-order interferometers, respectively.
-D is photon detector.
-CC is coincidence count detection scheme.
+The second-order interference of light is qualitatively different from the first-order interference of light.
+It may be the reason why people took more than 200 years to observe the first phenomenon of the second-order interference of light since the first observation of the first-order interference of light.
+On the other hand, generalizing the second-oder interference of light to the third- and higher-order interference of light is straightforward.
+The third- and higher-order interference of light is quantitatively different from the second-order interference of light.
+Figure 10: (a) First-, (b) second- and (c) third-order interference of light.
+S is a light source.
+IT1, IT2, and IT3 are first-, second- and third-order interferometers, respectively.
+D is a single-photon detector.
+CC is a coincidence count detection scheme.
 Figure 10 shows the schemes for the first-, second- and third-order interference of light.
@@
 10(a) is illustrated first.
-In a typical experiment of first-order interference of light, light generated by the light source S is split into more than one components in the first-order interferometer IT1.
-Then the split light beams are recombined to observe the intensity distribution over space or time by a single photo-detector.
-Young’s double-slit interferometer, Mach-Zehnder interferometer, Michelson interferometer are the typical first-order interferometers Born and Wolf (2013).
+In a typical experiment of first-order interference of light, light generated by the light source, S, is split into more than one components in the first-order interferometer, IT1.
+Then the split light beams are recombined to observe the intensity distribution over space or time by one detector.
+Young’s double-slit interferometer, Mach-Zehnder interferometer, Michelson interferometer are some typical first-order interferometers Born and Wolf (2013).
 The first-order interference of two and multiple independent light beams introduced in Sects.
@@
 10(b).
-Light emitted by source S is split into two or more different components in the second-order interferometer (IT2).
-Then the output light beams of IT2 are detected by two detectors (D1 and D2).
+Light emitted by S is split into two or more different components in the second-order interferometer, IT2.
+Then the output light beams of IT2 are detected by two detectors, D1 and D2.
 Hanbury Brown-Twiss (HBT) interferometer Brown and Twiss (1956b, a), Hong-Ou-Mandel (HOM) interferometer Hong et al.
 (1987); Shih and Alley (1988); Alley and Shih (1986), Franson interferometer Franson (1989) are some typical second-order interferometers.
-The output signals of these two detectors are sent into coincidence count detection system (CC) to measure the correlation of these two signals.
-Based on what types of photo-detector is employed, there are two work principles for CC.
-The first one is for normal photo-electric detector.
+The output signals of these two detectors are sent into a coincidence count detection system, CC, to measure the correlation of these two signals.
+Based on what types of detector is employed, there are two work principles for CC.
+The first one is for normal detectors.
 The output of the detector is electric current or voltage proportional to the intensity of the incidental light.
-The measured correlation is
+The measured correlation is given by the product of these two output signals,
 $\Gamma^{(2)}(\vec{r}_{1},t_{1};\vec{r}_{2},t_{2})=\langle I_{1}(\vec{r}_{1},t_{1})I_{2}(\vec{r}_{2},t_{2})\rangle,$ (47)
@@
 Equation (47) measures the correlation of intensity fluctuations detected by these two detectors.
-The CC system based on normal photo-electric detector was employed by Hanbury Brown and Twiss in their correlation experiments Brown and Twiss (1956b, a).
-The second one is for single-photon detector.
-The output of single-photon detector is a voltage pulse.
-Equation (47) can not be employed to calculate the correlation of the outputs from single-photon detectors.
+The CC system based on normal photo-electric detectors was employed by Hanbury Brown and Twiss in their experiments Brown and Twiss (1956b, a).
+The second principle is for single-photon detectors.
+The output of a single-photon detector is a voltage pulse.
+Equation (47) can not be employed to calculate the correlation of the outputs from two single-photon detectors.
 A different coincidence count principle should be employed.
-When the time difference between these two photon detection events is less than a time window $T$, these two photon detection events is a two-photon coincidence count.
-Otherwise, these two photon detection is not a two-photon coincidence count.
+When the time difference between these two photon detection events is less than a time window $T$, these two photon detection events are a two-photon coincidence count.
+Otherwise, these two photon detection events are not a two-photon coincidence count.
 The time window $T$ should be less than the coherence time due to these two photons are indistinguishable within the same coherence volume.
@@
 (2009), sum frequency generation Pe’er et al.
-(2007), which are also based on the same principle as the one for single-photon detector.
-The second-order interference shown in Fig.
-10(b) can be extended to the third- and higher-order interference.
+(2007), which are also based on the same principle as the one for single-photon detectors.
+The second-order interference of light shown in Fig.
+10(b) can be extended to the third- and higher-order interference of light.
 Figure 10(c) shows the scheme for the third-order interference of light.
-The light beam emitted by source S is split into multiple components and detected by three detectors, D1, D2 and D3.
-The output of these three detectors are sent into a three-photon coincidence count system (CC) to measure the third-order correlation.
-Based on different types of detectors are employed, there are also two different types of principle for three-photon coincidence count system.
-When normal photo-detectors are employed, the third-order correlation is given by
+The light beam emitted by S is split into multiple components and detected by three detectors, D1, D2 and D3.
+The output of these three detectors are sent into a three-photon coincidence count system, CC, to measure the third-order correlation.
+Based on different types of the employed detectors, there are also two different principles for three-photon coincidence count system.
+When normal detectors are employed, the third-order correlation is given by
 $\Gamma^{(3)}(\vec{r}_{1},t_{1};\vec{r}_{2},t_{2};\vec{r}_{3},t_{3})=\langle I_{1}(\vec{r}_{1},t_{1})I_{2}(\vec{r}_{2},t_{2})I_{3}(\vec{r}_{3},t_{3})\rangle,$ (48)
@@
 (47).
-When single-photon detectors are employed, the time differences between these three photon detection events are less than a time window $T$, the three photon detection events is a three-photon coincidence count.
-Otherwise, the three photon detection events is not a three-photon coincidence count.
+When single-photon detectors are employed, the time differences between these three photon detection events are less than a time window $T$, the three photon detection events are a three-photon coincidence count.
+Otherwise, the three photon detection events are not a three-photon coincidence count.
 Similar method can be employed to define the fourth- and higher-order interference of light.
 IV.1 Second-order interference of one light beam
-Figure 11 shows the second-order interference one light beam in a HBT interferometer.
-The light beam emitted by source S is split into two by a 1:1 non-polarizing beam splitter BS.
+Figure 11 shows the scheme for the second-order interference one light beam in a HBT interferometer.
+The light beam emitted by source S is split into two by a 1:1 non-polarizing beam splitter, BS.
 The output light beams of BS are detected by two single-photon detectors, D1 and D2.
-The outputs of these two detectors are sent into two-photon coincidence count system CC to measure the second-order correlation of light.
-Figure 11: One light beam in a HBT interferometer. S is light source. BS is 1:1 nonpolarizing beam splitter. D1 and D2 are two single-photon detectors. CC is two-photon coincidence count system.
+The outputs of these two detectors are sent into a two-photon coincidence count detection system, CC, to measure the second-order correlation of light.
+Figure 11: Second-order interference of one light beam in a HBT interferometer.
+S is a light source.
+BS is a 1:1 non-polarizing beam splitter.
+D1 and D2 are two single-photon detectors.
+CC is a two-photon coincidence count detection system.
 IV.1.1 Second-order interference of one thermal light beam
@@
 11, the scheme shown in Fig.
-11 is equivalent to the scheme employed by Hanbury Brown and Twiss discovering two-photon bunching of thermal light in 1956 Brown and Twiss (1956b, a).
-There are two different paths for D1 and D2 detecting a two-photon coincidence count in Fig.
-11.
-The first path is photon a is detected by D1 and photon b is detected by D2.
+11 is equivalent to the scheme employed by Hanbury Brown and Twiss in 1956 Brown and Twiss (1956b, a).
+There are two different paths for D1 and D2 detecting a two-photon coincidence count.
+The first path is that photon a is detected by D1 and photon b is detected by D2.
 The probability for the path is $1/2$ and the corresponding two-photon probability amplitude is $A(a,D_{1};b,D_{2})$.
-The second path is photon a is detected by D2 and photon b is detected by D1.
+The second path is that photon a is detected by D2 and photon b is detected by D1.
 The probability for the path is $1/2$ and the corresponding two-photon probability amplitude is $A(a,D_{2};b,D_{1})$.
@@
 $P^{(2)}_{j}(\vec{r}_{1},t_{1};\vec{r}_{2},t_{2})$ (49)
-$=$ $|A_{j}(a,D_{1};b,D_{2})+A_{j}(a,D_{2};b,D_{1})|^{2}$
-$=$ $|e^{i(\varphi_{aj}+\varphi_{bj})}K_{j}(a,D_{1};b,D_{2})$
+$=$ $|\frac{1}{\sqrt{2}}[A_{j}(a,D_{1};b,D_{2})+A_{j}(a,D_{2};b,D_{1})]|^{2}$
+$\propto$ $|e^{i(\varphi_{aj}+\varphi_{bj})}K_{j}(a,D_{1};b,D_{2})$
 $+e^{i(\varphi_{aj}+\varphi_{bj})}K_{j}(a,D_{2};b,D_{1})|^{2}$
@@
 The meaning of $K_{j}(a,D_{2};b,D_{1})$ is defined similarly.
-Based on the principle for the probability amplitude of independent events introduced by Feynman, $K_{j}(a,D_{1};b,D_{2})=K_{j}(a,D_{1})K_{j}(b,D_{2})$.
+Based on the principle for calculating the probability amplitude of independent events introduced by Feynman Feynman et al.
+(2011), $K_{j}(a,D_{1};b,D_{2})$ equals $K_{j}(a,D_{1})K_{j}(b,D_{2})$.
 Equation (49) can be simplified as
@@
 $=$ $|K_{j}(a,D_{1})K_{j}(b,D_{2})+K_{j}(a,D_{2})K_{j}(b,D_{1})|^{2}.$
-The two-photon probability function of $N$ detected photon pairs is the sum of the probabilities of $N$ detected photon pairs,
+The two-photon probability function of $N$ detected photon pairs is the sum of the probability functions of $N$ detected photon pairs,
 $P^{(2)}(\vec{r}_{1},t_{1};\vec{r}_{2},t_{2})$
@@
 $p^{(2)}(\vec{r}_{1},t_{1};\vec{r}_{2},t_{2})=\frac{P^{(2)}(\vec{r}_{1},t_{1};\vec{r}_{2},t_{2})}{P^{(1)}(\vec{r}_{1},t_{1})P^{(1)}(\vec{r}_{2},t_{2})},$ (52)
-is employed to eliminate the influence of the first-order interference on the second-order interference, where $P^{(1)}(\vec{r}_{1},t_{1})$ and $P^{(1)}(\vec{r}_{2},t_{2})$ are single-photon probability function at space-time coordinates $(\vec{r}_{1},t_{1})$ and $(\vec{r}_{2},t_{2})$, respectively.
-$P^{(1)}(\vec{r}_{1},t_{1})$ is written as $P(\vec{r}_{1},t_{1})$ if no misunderstanding exists.
-Two-photon probability function $P^{(2)}(\vec{r}_{1},t_{1};\vec{r}_{2},t_{2})$ and normalized two-photon probability function $p^{(2)}(\vec{r}_{1},t_{1};\vec{r}_{2},t_{2})$ correspond to the second-order coherence function $G^{(2)}(\vec{r}_{1},t_{1};\vec{r}_{2},t_{2})$ and normalized second-order coherence function $g^{(2)}(\vec{r}_{1},t_{1};\vec{r}_{2},t_{2})$ in Glauber’s quantum optical coherence theory, respectively Glauber (1963a, c).
+is employed to eliminate the influence of the first-order interference pattern on the second-order interference pattern, where $P^{(1)}(\vec{r}_{1},t_{1})$ and $P^{(1)}(\vec{r}_{2},t_{2})$ are single-photon probability functions at space-time coordinates $(\vec{r}_{1},t_{1})$ and $(\vec{r}_{2},t_{2})$, respectively.
+$P^{(1)}(\vec{r}_{1},t_{1})$ is usually written as $P(\vec{r}_{1},t_{1})$ for short if no possible misunderstanding exists.
+The two-photon probability function, $P^{(2)}(\vec{r}_{1},t_{1};\vec{r}_{2},t_{2})$, and the normalized two-photon probability function, $p^{(2)}(\vec{r}_{1},t_{1};\vec{r}_{2},t_{2})$, correspond to the second-order coherence function, $G^{(2)}(\vec{r}_{1},t_{1};\vec{r}_{2},t_{2})$, and the normalized second-order coherence function, $g^{(2)}(\vec{r}_{1},t_{1};\vec{r}_{2},t_{2})$, in Glauber’s quantum optical coherence theory, respectively Glauber (1963a, c).
 Let us first calculate temporal two-photon bunching of thermal light.
-Assuming S is a point light source, D1 and D2 are at symmetrical positions of BS to simplify the calculation.
+Assume S is a point thermal light source.
+D1 and D2 are at symmetrical positions of BS to simplify the calculation.
 In this case, photon’s Feynman propagator is
 […]
-where the spatial part is ignored, $\omega_{0}$ is the central angular frequency, $\Delta\omega$ is angular frequency distribution function of thermal light, $t_{1}$ is the time of photon detection event at D1, $t_{a}$ is the emitting time of photon a.
+where the spatial part is ignored, $\omega_{0}$ is the central angular frequency, $\Delta\omega$ is the angular frequency distribution function of thermal light, $t_{1}$ is the time of photon detection event at D1, $t_{a}$ is the emitting time of photon a.
 Substituting Eq.
 (53) into Eq.
-(IV.1.1) and considering $f(\omega)$ equals 1 within $[\omega_{0}-\Delta\omega/2,\omega_{0}+\Delta\omega/2]$ and equals 0 in other domains, the temporal two-photon probability function is Shih (2020)
+(IV.1.1) and assuming $f(\omega)$ equals 1 within $[\omega_{0}-\Delta\omega/2,\omega_{0}+\Delta\omega/2]$ and equals 0 in other domains, the temporal two-photon probability function is Shih (2020)
 $P^{(2)}(t_{1}-t_{2})\propto 1+\text{sinc}^{2}\frac{\Delta\omega(t_{1}-t_{2})}{2}.$ (54)
-It is easy to prove that $P^{(1)}(t_{1})$ and $P^{(2)}(t_{1})$ are constant for thermal light.
+It is easy to prove that $P^{(1)}(t_{1})$ and $P^{(1)}(t_{2})$ are constant for thermal light.
 The normalized temporal two-photon probability function, $p^{(2)}(t_{1}-t_{2})$ is proportional to $P^{(2)}(t_{1}-t_{2})$.
-In the following, if the normalized probability function is proportional to the corresponding probability function, only probability function is calculated.
+In the following parts, if the normalized probability function is proportional to the corresponding probability function, only probability function is calculated.
 Figure 12: Temporal two-photon bunching of thermal light.
-$t_{1}-t_{2}$ is the relative position difference between D1 and D2.
-Lines 1, 2, 3 correspond to the frequency bandwidths equal $\Delta\omega$, $2\Delta\omega$, and $4\Delta\omega$, respectively.
+$t_{1}-t_{2}$ is the time difference between photon detection events at D1 and D2.
+Lines 1, 2, 3 correspond to the frequency bandwidths equaling $\Delta\omega$, $2\Delta\omega$, and $4\Delta\omega$, respectively.
 Figure 12 shows the temporal two-photon probability functions of thermal light with different frequency bandwidths.
-When two detectors are at symmetrical positions, the probability of detecting two photons equal 2.
+When $t_{1}$ equals $t_{2}$, the probability of detecting two photons equals 2.
 As the time difference, $|t_{1}-t_{2}|$, increases, the probability of detecting two photons decreases.
-When the time difference between two photon detection events is much longer than the coherence time, the probability of detecting two photons equal 1, which means that these two photon detection events are independent.
-The spatial two-photon bunching of thermal light can be calculated in the similar way.
-Assuming one dimension light source is employed, the light emitted by S is single frequency light, and the distance between the source plane and D1 is equal the one of source plane and D2.
-photon’s Feynman propagator in this case can be approximated by Green function in classical optics Goodman (2005)
+When the time difference between two photon detection events is much longer than the coherence time, the probability of detecting two photons equals 1, which indicates that these two photon detection events are independent.
+The spatial two-photon bunching of thermal light can be calculated in the same way.
+Assume one dimension thermal light source is employed.
+The light emitted by S is single-frequency light.
+The distance between the source plane and D1 equals the one between source plane and D2.
+The photon’s Feynman propagator in this case can be approximated by Green function in classical optics Goodman (2005)
 […]
-where $d$ is the length of thermal light source, $s(x_{s})$ is the distribution function of S, $\vec{r}_{1}$ is position vector of D1, $\vec{r}_{s}$ is the position vector of S emitting photon a, $\vec{k}_{s1}$ is wave vector of light.
+where $d$ is the length of thermal light source, $s(x_{s})$ is the intensity distribution function of S, $\vec{r}_{1}$ is the position vector of D1, $\vec{r}_{s}$ is the position vector of S emitting photon a, $\vec{k}_{s1}$ is the wave vector of light.
 Substituting Eq.
 (55) into Eq.
-(IV.1.1) and considering $s(x_{s})$ equals 1 within $[-d/2,d/2]$ and equals 0 in other domains, spatial two-photon probability function of thermal light in a HBT interferometer is Shih (2020)
+(IV.1.1) and assuming $s(x_{s})$ equals 1 within $[-d/2,d/2]$ and equals 0 in other domains, spatial two-photon probability function of thermal light in a HBT interferometer is Shih (2020)
 $P^{(2)}(x_{1}-x_{2})\propto 1+\text{sinc}^{2}\frac{d}{L\lambda}(x_{1}-x_{2}).$ (56)
-Based on the results above, two-photon bunching of thermal light can be interpreted as the result of two-photon interference, which is the superposition of two-photon probability amplitudes for two different and indistinguishable paths to trigger a two-photon coincidence count Fano (1961); Feynman (2006); Glauber (1995).
-Two-photon interference in the second-order interference of light can not be used to disprove Dirac’s statement that different photons never interferes with each other in the first-order interference of light Paul (1986).
-By analogy, Dirac’s statement about the first-order interference of light can be extended to the second-order interference of light: Each photon pair only interferes with itself and interferes between two different photon pairs never occurs Shih (2020).
+Based on the results above, two-photon bunching of thermal light can be interpreted as the result of two-photon interference.
+It is the result of superposing two-photon probability amplitudes for two different and indistinguishable paths to trigger a two-photon coincidence count Fano (1961); Feynman (2006); Glauber (1995).
+Two-photon interference in the second-order interference of light can not be used to disprove Dirac’s statement that different photons never interferes with each other, which was meant for the first-order interference of light Paul (1986).
+By analogy, Dirac’s statement about the first-order interference of light can be generalized to the second-order interference of light: Each photon pair only interferes with itself and interferes between two different photon pairs never occurs Shih (2020).
 IV.1.2 Second-order interference of one superbunching light beam
-$g^{(2)}(0)$ equals 2 for thermal light.
-$g^{(2)}(0)$ measured in the HBT interferometer is the degree of second-order coherence of light, which is usually employed to categorize different types of light Loudon (2000).
+$g^{(2)}(0)$ measured in a HBT interferometer is the degree of second-order coherence of light, which is usually employed to categorize different types of light Loudon (2000).
 $g^{(2)}(0)$ equals 1 for single-mode continuous-wave laser light.
 When $g^{(2)}(0)$ is less than 1, the light belongs to non-classical light.
-When $g^{(2)}(0)$ is larger than 2, it is superbunching light Ficek and Swain (2005).
-Superbunching light can be either classical or nonclassical.
+$g^{(2)}(0)$ equals 2 for thermal light.
+When $g^{(2)}(0)$ is larger than 2, it belongs to superbunching light Ficek and Swain (2005).
+Superbunching light can be either classical or non-classical.
 For instance, entangled photon pairs is non-classical superbunching light Klyshko (1996) and superbunching pseudothermal light is classical light Zhou et al.
 (2017).
-This section will concentrate on the second-order interference of a classical superbunching light beam in HBT interferometer.
+This section will concentrate on the second-order interference of a classical superbunching light beam in a HBT interferometer.
 There are two different methods to increase the degree of second-order coherence of classical light.
 The first one is adding more two-photon paths.
-The second one is the two-photon probability amplitudes changes with time.
+The second one is the two-photon probability amplitudes changing with time.
 Figure 13: Increase the degree of second-order coherence by adding more two-photon paths.
+a and b are two photons.
+D1 and D2 are two single-photon detectors.
+$\alpha$ and $\beta$ are two nodes for photons.
 Figure 13 shows how to increase the degree of second-order coherence by adding more two-photon paths.
-Figure 13(a) shows there are two different ways for photon a and b to trigger a two-photon coincidence count in D1 and D2.
-The first one is photon a is detected by D1 and photon b is detected by D2.
+Figure 13(a) shows there are two different paths for photons a and b to trigger a two-photon coincidence count in D1 and D2.
+The first path is that photon a is detected by D1 and photon b is detected by D2.
 The corresponding two-photon probability amplitude is $A(a,D_{1};b,D_{2})$.
-The second one is photon a is detected by D2 and photon b is detected by D1.
+The second path is that photon a is detected by D2 and photon b is detected by D1.
 The corresponding two-photon probability amplitude is $A(a,D_{2};b,D_{1})$.
@@
 $P^{(2)}(\vec{r}_{1},t_{1};\vec{r}_{2},t_{2})$ (57)
-$=$ $\langle|A(a,D_{1};b,D_{2})+A(a,D_{2};b,D_{1})|^{2}\rangle.$
+$=$ $\langle|\frac{1}{\sqrt{2}}[A(a,D_{1};b,D_{2})+A(a,D_{2};b,D_{1})]|^{2}\rangle.$
 When these two different paths are distinguishable, the two-photon probability function is
 $P^{(2)}(\vec{r}_{1},t_{1};\vec{r}_{2},t_{2})$ (58)
-$=$ $\langle|A(a,D_{1};b,D_{2})|^{2}\rangle+\langle|A(a,D_{2};b,D_{1})|^{2}\rangle,$
-which will contribute to the background of two-photon probability function.
+$=$ $\frac{1}{2}[\langle|A(a,D_{1};b,D_{2})|^{2}\rangle+\langle|A(a,D_{2};b,D_{1})|^{2}\rangle],$
+which consists of the background of two-photon probability function.
 Figure 13(a) and Eqs.
-(57) are valid for two-photon bunching of thermal light.
+(57) are valid for thermal light in a HBT interferometer.
 $g^{(2)}(0)$ equals 2 for thermal light.
@@
 The are four terms in Eq.
-(57) and $P^{(2)}(0)$ equals $4|A|^{2}$, where $A$ is short for $|A(a,D_{1};b,D_{2})|$.
+(57) and $P^{(2)}(0)$ equals $2|A|^{2}$, where $A$ is short for $|A(a,D_{1};b,D_{2})|$.
 There are 2 terms in Eq.
-(58) and $P^{(2)}(0)$ equals $2|A|^{2}$.
+(58) and $P^{(2)}(0)$ equals $|A|^{2}$.
 The degree of second-order coherence equals the normalized two-photon probability function when these two detectors are at the same space-time coordinate,
-$g^{(2)}(0)=p^{(2)}(0)=\frac{4|A|^{2}}{2|A|^{2}}=2.$ (59)
+$g^{(2)}(0)=p^{(2)}(0)=\frac{2|A|^{2}}{|A|^{2}}=2.$ (59)
 The reason why Eq.
 (58) can be treated as the background is that these two paths are distinguishable.
-The two photon detection events are independent and $P^{(2)}(\vec{r}_{1},t_{1};\vec{r}_{2},t_{2})$ equals $P^{(1)}(\vec{r}_{1},t_{1})P^{(1)}(\vec{r}_{2},t_{2})$.
+The two photon detection events are independent and $P^{(2)}(\vec{r}_{1},t_{1};\vec{r}_{2},t_{2})$ equals $P^{(1)}(\vec{r}_{1},t_{1})P^{(1)}(\vec{r}_{2},t_{2})$ in this case.
 Similar method can be employed to calculate the degree of second-order coherence of light in Fig.
 13(b).
-There are four different paths for photons a and b triggering a two-photon coincidence count.
-The first one is photon a transmitting through $\alpha$ is detected by D1 and photon b transmitting through $\beta$ is detected by D2.
+There are four different paths for photons a and b to trigger a two-photon coincidence count.
+The first path is that photon a transmitting through $\alpha$ is detected by D1 and photon b transmitting through $\beta$ is detected by D2.
 The corresponding two-photon probability amplitude is $A(a,\alpha,D_{1};b,\beta,D_{2})$.
 The other three paths can be defined similarly and the corresponding two-photon probability amplitudes are $A(a,\alpha,D_{2};b,\beta,D_{1})$, $A(a,\beta,D_{1};b,\alpha,D_{2})$ and $A(a,\beta,D_{2};b,\alpha,D_{1})$.
-When these four different paths to trigger a two-photon coincidence count are indistinguishable, the two-photon probability function
+When these four different paths to trigger a two-photon coincidence count are indistinguishable, the two-photon probability function is
 $P^{(2)}(\vec{r}_{1},t_{1};\vec{r}_{2},t_{2})$
-$=$ $\langle|A(a,\alpha,D_{1};b,\beta,D_{2})+A(a,\alpha,D_{2};b,\beta,D_{1})$
-$+A(a,\beta,D_{1};b,\alpha,D_{2})+A(a,\beta,D_{2};b,\alpha,D_{1})|^{2}\rangle.$
-When these four different paths are distinguishable, the two-photon probability function
+$=$ $\langle|\frac{1}{\sqrt{4}}[A(a,\alpha,D_{1};b,\beta,D_{2})+A(a,\alpha,D_{2};b,\beta,D_{1})$
+$+A(a,\beta,D_{1};b,\alpha,D_{2})+A(a,\beta,D_{2};b,\alpha,D_{1})]|^{2}\rangle.$
+When these four different paths are distinguishable, the two-photon probability function equals
 $P^{(2)}(\vec{r}_{1},t_{1};\vec{r}_{2},t_{2})$
-$=$ $\langle|A(a,\alpha,D_{1};b,\beta,D_{2})|^{2}\rangle+\langle|A(a,\alpha,D_{2};b,\beta,D_{1})|^{2}\rangle$
-$+\langle|A(a,\beta,D_{1};b,\alpha,D_{2})|^{2}\rangle+\langle|A(a,\beta,D_{2};b,\alpha,D_{1})|^{2}\rangle$
-Again, assuming photons have equal probability being detected by D1 and D2. The ratio between Eqs. (IV.1.2) and (IV.1.2) at the same space-time coordinate equals the degree of second-order coherence,
-$g^{(2)}(0)=p^{(2)}(0)=\frac{16|A^{\prime}{}|^{2}}{4|A^{\prime}{}|^{2}}=4,$ (62)
+$=$ $\frac{1}{4}[\langle|A(a,\alpha,D_{1};b,\beta,D_{2})|^{2}\rangle+\langle|A(a,\alpha,D_{2};b,\beta,D_{1})|^{2}\rangle$
+$+\langle|A(a,\beta,D_{1};b,\alpha,D_{2})|^{2}\rangle+\langle|A(a,\beta,D_{2};b,\alpha,D_{1})|^{2}\rangle]$
+Again, assume photons have equal probability being detected by D1 and D2. The ratio between Eqs. (IV.1.2) and (IV.1.2) at the same space-time coordinate equals the degree of second-order coherence,
+$g^{(2)}(0)=p^{(2)}(0)=\frac{4|A^{\prime}{}|^{2}}{|A^{\prime}{}|^{2}}=4,$ (62)
 where $|A^{\prime}{}|$ is short for $|A(a,\alpha,D_{1};b,\beta,D_{2})|$.
@@
 (2017).
-Figure 14: Experimental scheme to observe two-photon superbunching with $2N$ different paths Zhou et al. (2017). P1, P2, …, PN are $N$ pinholes. RG1, RG2, …, RGN are $N$ rotating groundglasses.
+Figure 14: Scheme to observe two-photon superbunching with $2N$ different paths Zhou et al. (2017). P1, P2, …, PN are $N$ pinholes. RG1, RG2, …, RGN are $N$ rotating groundglasses.
 Figure 14 shows the experimental scheme to observe two-photon superbunching with $2N$ different paths.
-Single-mode continuous-wave laser is employed as the light source.
+A single-mode continuous-wave laser is employed as the light source.
 After passing through pinhole P1, the light is scattered by a rotating groundglass, RG1.
-Then the scattered light is filtered with another pinhole, P2.
-The size of P2 should be less than the coherence area of scattered light at P2 plane.
+Then the scattered light is filtered by another pinhole, P2.
+The size of P2 should be less than the coherence area of the scattered light at P2 plane.
 Then the filtered light is scattered by RG2.
@@
 $t_{1}-t_{2}$ is the time difference between two photon detection events at D1 and D2.
-$g^{(2)}(t_{1}-t_{2})$ is the normalized second-order coherence function in Glauber’s quantum optical coherence theory, which is equivalent to the normalized two-photon probability function.
-Figure 15(a)-(c) shows the measured temporal second-order coherence function of light with one, two, and three rotating groundglasses, respectively Bai et al.
+$g^{(2)}(t_{1}-t_{2})$ is the normalized second-order coherence function in Glauber’s quantum optical coherence theory, which is equivalent to the normalized two-photon probability function in quantum optical coherence theory based on Feynman’s path integral.
+Figure 15(a)-(c) shows the measured temporal second-order coherence functions of light with one, two, and three rotating groundglasses, respectively Bai et al.
 (2017).
@@
 15(a), in which there is one rotating groundglass and two-photon bunching is observed.
-The reason why it is less than 2 is due to part of the un-scattered laser light contribute to the two-photon coincidence count.
+The reason why it is less than 2 is due to that part of the un-scattered laser light contributes to the two-photon coincidence count.
 The source with one rotating groundglass is pseudothermal light source invented by Martienssen and Spiller Martienssen and Spiller (1964).
 $g^{(2)}(0)$ equals $3.80\pm 0.04$ in Fig.
-15(b) when there are two rotating groundglass.
-When there are three rotating groundglass, $g^{(2)}(0)$ equals $7.10\pm 0.07$ in Fig.
+15(b) when there are two rotating groundglasses.
+When there are three rotating groundglasses, $g^{(2)}(0)$ equals $7.10\pm 0.07$ in Fig.
 15(c).
@@
 15(b) and 15(c).
-Another method to increase the degree of second-order coherence of light is the two-photon probability amplitudes changes with time.
+Another method to increase the degree of second-order coherence of light is that the two-photon probability amplitudes changes with time.
 Figure 16 shows the scheme for observing two-photon superbunching by changing the two-photon probability amplitudes.
 The light source shown in Fig.
-16 as the light source in a HBT interferometer shown in Fig.
+16 is the light source in a HBT interferometer shown in Fig.
 11.
@@
 The two-photon probability amplitudes are $A(a,D_{1};b,D_{2})$ and $A(a,D_{2};b,D_{1})$.
-Figure 16: Simplified superbunching pseudothermal light sourc Zhou et al. (2019b). IM is intensity modulator and RG is rotating groudglass.
-The difference between usual pseudothermal light source and the one shown in Fig.
-16in the HBT interferometer is $A(a,D_{1};b,D_{2})$ and $A(a,D_{2};b,D_{1})$ will change with time in the latter due to that intensity modulator (IM) is employed to modulate the intensity of incidental laser light.
-When these two different paths to trigger a two-photon coincidence count are indistinguishable, the temporal two-photon probability function
+Figure 16: Simplified superbunching pseudothermal light source Zhou et al. (2019b). IM is intensity modulator and RG is rotating groundglass.
+The difference between usual pseudothermal light source and the light source shown in Fig.
+16 in a HBT interferometer is that $A(a,D_{1};b,D_{2})$ and $A(a,D_{2};b,D_{1})$ will change with time in the latter case.
+An intensity modulator (IM) is employed to modulate the intensity of incidental laser light.
+When these two different paths to trigger a two-photon coincidence count are indistinguishable, the temporal two-photon probability function is
 $P^{(2)}(t_{1};t_{2})$ $=$ $\langle|\sqrt{P_{a1}(t_{a1})P_{b2}(t_{b2})}K_{a1}K_{b2}$ (63)
-$+\sqrt{P_{a2}(t_{a2})P_{b1}(t_{b1})}K_{a2}K_{b1}|^{2}\rangle$
+$+\sqrt{P_{a2}(t_{a2})P_{b1}(t_{b1})}K_{a2}K_{b1}|^{2}\rangle,$
 where $P_{a1}(t_{a1})$ is the probability that photon a is the detected by D1 at time $t_{a1}$ and the meanings of other symbols are defined similarly.
@@
 $\propto$ $\gamma(t_{1}-t_{2})[1+\text{sinc}^{2}\frac{\Delta\omega(t_{2}-t_{1})}{2}],$
-where $\gamma(t_{1}-t_{2})=\langle I(t_{1})I(t_{2})\rangle/(\langle I(t_{1})\rangle\langle I(t_{2})\rangle)$ is the correlation function of laser light introduced by IM, $I(t_{1})$ and $I(t_{2})$ are the intensities of light before RG, respectively.
+where $\gamma(t_{1}-t_{2})=\langle I(t_{1})I(t_{2})\rangle/(\langle I(t_{1})\rangle\langle I(t_{2})\rangle)$ is the correlation function of laser light introduced by IM, $I(t_{1})$ and $I(t_{2})$ are the intensities of light at time $t_{1}$ and $t_{2}$ before RG, respectively.
 When $\gamma(t_{1}-t_{2})$ equals 1, Eq.
 (IV.1.2) describes normal two-photon bunching of thermal light.
-When $\gamma(t_{1}-t_{2})$ is larger than 1, two-photon suerpbunching can be observed.
-Figure 17: Observed two-photon superbunching with changing probability amplitudes Zhou et al. (2019b).
-Figure 17 shows the experimental measured temporal second-order coherence function of superbunching pseudotheral light with the scheme shown in Fig.
+When $\gamma(t_{1}-t_{2})$ is larger than 1, two-photon superbunching can be observed.
+Figure 17: Two-photon superbunching with changing probability amplitudes Zhou et al. (2019b).
+Figure 17 shows the experimental measured temporal second-order coherence function of superbunching pseudothermal light with the scheme shown in Fig.
 16.
-Two-photon superbuching with $g^{(2)}(0)$ equaling 20.45 is observed Zhou et al.
+Two-photon superbunching with $g^{(2)}(0)$ equaling 20.45 is observed Zhou et al.
 (2019b).
@@
 IV.1.3 Second-order interference of one laser light beam
-Assuming S is a single-mode continuous-wave laser in Fig.
+Assume S is a single-mode continuous-wave laser in Fig.
 11.
@@
 The probability amplitudes for these two paths are $A(a,D_{1};b,D_{2})$ and $A(a,D_{2};b,D_{1})$.
-If these two different ways are indistinguishable, the two-photon probability is
+If these two different paths are indistinguishable, the two-photon probability function is
 $P^{(2)}(\vec{r}_{1},t_{1};\vec{r}_{2},t_{2})$ (65)
-$=$ $\langle|A(a,D_{1};b,D_{2})+A(a,D_{2};b,D_{1})|^{2}\rangle.$
-It seems that two-photon bunching should also be observed for single-mode continuous-wave laser light in a HBT interferometer.
-However, it was proved that photon detection events are independent when a single-mode continuous-wave laser light beam in a HBT interferometer Arecchi (1965); Freed and Haus (1965).
-It is predicted in Glauber’s quantum optical coherence theory that the degree of second-order coherence of coherent state equals 1 Glauber (1963a), which indicates that photon detection events are independent in single-mode continuous-wave laser light.
+$=$ $\langle|\frac{1}{\sqrt{2}}[A(a,D_{1};b,D_{2})+A(a,D_{2};b,D_{1})]|^{2}\rangle.$
+It seems that two-photon bunching should be observed for single-mode continuous-wave laser light in a HBT interferometer.
+However, it was proved that photon detection events are independent for a single-mode continuous-wave laser light beam in a HBT interferometer Arecchi (1965); Freed and Haus (1965).
+It is predicted that the degree of second-order coherence of coherent state equals 1 in Glauber’s quantum optical coherence theory Glauber (1963a), which indicates that the photon detection events are independent in single-mode continuous-wave laser light.
 The reason why there is no two-photon bunching of single-mode continuous-wave laser light lies at the difference between the superposition principles in classical and quantum theories.
-In classical theory, assuming there are two fields at space-time coordinate $(\vec{r},t)$, the final field is the superposition of these two fields,
+In classical theory, if there are two fields at space-time coordinate $(\vec{r},t)$, the final field is the superposition of these two fields,
 $\vec{E}(\vec{r},t)=\vec{E}_{1}(\vec{r},t)+\vec{E}_{2}(\vec{r},t).$ (66)
-If $\vec{E}_{1}(\vec{r},t)$ is identical to $\vec{E}_{2}(\vec{r},t)$, the superposition will give $2\vec{E}_{1}(\vec{r},t)$.
-The intensity at that point will be four times of the one of $\vec{E}_{1}(\vec{r},t)$.
-This is not true for the superposition in quantum theory. Assuming there are two wave functions at space-time coordinate $(\vec{r},t)$, the final wave function is the superposition of these two,
+If $\vec{E}_{1}(\vec{r},t)$ is identical to $\vec{E}_{2}(\vec{r},t)$, the superposition will result in $2\vec{E}_{1}(\vec{r},t)$.
+The intensity at the point will be four times of the one of $\vec{E}_{1}(\vec{r},t)$.
+This is not the case for the superposition principle in quantum theory.
+Assuming there are two wave functions at space-time coordinate $(\vec{r},t)$, the final wave function is the superposition of these two wave functions,
 $\Psi(\vec{r},t)=\Psi_{1}(\vec{r},t)+\Psi_{2}(\vec{r},t).$ (67)
 If $\Psi_{1}(\vec{r},t)$ is identical to $\Psi_{2}(\vec{r},t)$, the superposition will not give $2\Psi_{1}(\vec{r},t)$.
-The results of superposition of two identical states equals one of the state, not two times of the individual state Bohm (2012).
+The result of superposition of two identical states equals the individual state, not two times of the individual state Bohm (2012).
 The wave function can be related to the probability amplitude Feynman et al.
@@
 […]
-The superposition of two probability amplitudes for two identical paths will not give two times of each probability amplitude, but only one probability amplitude.
+The superposition of two probability amplitudes for two identical paths will not get two times of each probability amplitude, but only one probability amplitude.
 The two different paths in Eq.
@@
 $P^{(2)}(\vec{r}_{1},t_{1};\vec{r}_{2},t_{2})=\langle|A(S,D_{1};S,D_{2})|^{2}\rangle,$ (69)
-no second-order interference exists and $A(S,D_{1};S,D_{2})$ is the probability amplitude for two photons emitted by S going to D1 and D2, respectively.
-Comparing the two-photon probability functions of thermal and laser light in a HBT interferometer, the reason why these two cases are different is due to the initial phases of photons in thermal light are random and the initial phases of photons in laser light are identical within one coherence volume.
-This difference makes two different paths, $A(a,D_{1};b,D_{2})$ and $A(a,D_{2};b,D_{1})$ different for thermal light and identical for laser light.
-Hence two-photon bunching can be observed with thermal light and can not be observed for single-mode continuous-wave laser light.
+no second-order interference pattern exists and $A(S,D_{1};S,D_{2})$ is the probability amplitude for two photons emitted by S going to D1 and D2, respectively.
+Comparing the two-photon probability functions of thermal and laser light in a HBT interferometer, the reason why these two cases are different is due to the initial phases of photons are random in thermal light and the initial phases of photons are identical in laser light within one coherence volume.
+This difference makes two different paths, $A(a,D_{1};b,D_{2})$ and $A(a,D_{2};b,D_{1})$, different for thermal light and identical for laser light.
+There is two-photon bunching for thermal light and no two-photon bunching for single-mode continuous-wave laser light.
 IV.2 Second-order interference of two light beams
-The first second-order interference of two light beams experiment are correlation measurement of photon pairs generated in the annihilation of positron-electron pairs Hanna (1948); Bleuler and Bradt (1948); Wu and Shaknov (1950).
-Pfleegor and Mandel employed the second-order correlation of light to prove there exist transient first-order interference pattern for two independent lasers at low light level Pfleegor and Mandel (1967).
+The first second-order interference experiment with two light beams is correlation measurement of photon pairs generated in the annihilation of positron-electron pairs Hanna (1948); Bleuler and Bradt (1948); Wu and Shaknov (1950).
+Pfleegor and Mandel employed the second-order correlation of light to prove that there exists transient first-order interference pattern for two independent lasers at low light level Pfleegor and Mandel (1967).
 The most famous second-order interference of two light beams is the second-order interference of entangled photon pairs in a HOM interferometer Hong et al.
 (1987); Shih and Alley (1988); Alley and Shih (1986).
-Many other second-order interference of light experiments were performed, please see Refs.
+Many other second-order interference experiments with light were performed, please see Refs.
 Belinski and Klyshko (1993); Paul (1986); Ou (2007); Shih (2020) for details.
@@
 S1 and S2 are two light sources.
-BS is a 1:1 nonpolarizing beam splitter.
-D1 and D2 are two photo detectors.
-CC is coincidence count detection system.
-The second-order interference of two light beams in a HOM interferometer is shown in Fig.
+BS is a 1:1 non-polarizing beam splitter.
+D1 and D2 are two single-photon detectors.
+CC is a two-photon coincidence count detection system.
+The scheme for the second-order interference of two light beams in a HOM interferometer is shown in Fig.
 18.
-Two light beams emitted by source S1 and S2 are incident to two input ports of a 1:1 beam splitter (BS).
+Two light beams emitted by two sources, S1 and S2, are incident into two input ports of a 1:1 non-polarizing beam splitter, BS.
 The output light beams from BS are detected by two single-photon detectors, D1 and D2.
@@
 IV.2.1 Second-order interference of entangled photon pairs
-Assume two photons in the same entangled photon pair are in the scheme shown in Fig.
-18, which is similar as the original second-order interference with entangled photon pairs Hong et al.
+Assume two photons in one entangled photon pair are in the scheme shown in Fig.
+18, which is similar as the original second-order interference of entangled photon pairs Hong et al.
 (1987); Shih and Alley (1988); Alley and Shih (1986).
@@
 For simplicity, assume photon a is emitted by S1 and photon b is emitted by S2.
-The first path is photon a is detected by D1 and photon b is detected by D2.
+The first path is that photon a is detected by D1 and photon b is detected by D2.
 The probability for the path is 1/2 and the corresponding two-photon probability amplitude is $A_{a,D_{1};b,D_{2}}$.
-The second path is photon a is detected by D2 and photon b is detected by D1.
+The second path is that photon a is detected by D2 and photon b is detected by D1.
 The probability for the path is 1/2 and the corresponding two-photon probability amplitude is $A_{a,D_{2};b,D_{1}}$.
 If these two different paths are indistinguishable, the two-photon probability function is
-$P^{(2)}(\vec{r}_{1},t_{1};\vec{r}_{2},t_{2})$ $=$ $\langle|A_{a,D_{1};b,D_{2}}+A_{a,D_{2};b,D_{1}}|^{2}\rangle.$ (70)
-There is an extra $\pi/2$ phase of the reflected photon compared to the transmitted photon introduced by BS Loudon (2000), which can not be ignored in HOM interferometer as the one in HBT interferometer.
+$P^{(2)}(\vec{r}_{1},t_{1};\vec{r}_{2},t_{2})$ (70)
+$=$ $\langle|\frac{1}{\sqrt{2}}[A_{a,D_{1};b,D_{2}}+A_{a,D_{2};b,D_{1}}]|^{2}\rangle.$
+There is an extra $\pi/2$ phase for the reflected photon compared to the transmitted photon introduced by BS Loudon (2000), which can not be ignored in HOM interferometer.
 Equation (70) can be simplified as
 $P^{(2)}(\vec{r}_{1},t_{1};\vec{r}_{2},t_{2})$ (71)
-$=$ $\langle|e^{i\varphi_{a}}K_{aD_{1}}e^{i\varphi_{b}}K_{bD_{2}}+e^{i\varphi_{a}}e^{i\pi/2}K_{aD_{2}}e^{i\varphi_{b}}e^{i\pi/2}K_{bD_{1}}|^{2}\rangle$
+$\propto$ $\langle|e^{i\varphi_{a}}K_{aD_{1}}e^{i\varphi_{b}}K_{bD_{2}}+e^{i\varphi_{a}}e^{i\pi/2}K_{aD_{2}}e^{i\varphi_{b}}e^{i\pi/2}K_{bD_{1}}|^{2}\rangle$
 $=$ $\langle|K_{aD_{1}}K_{bD_{2}}-K_{aD_{2}}K_{bD_{1}}|^{2}\rangle,$
-where $\varphi_{a}$ and $\varphi_{b}$ are the initial phases of photon a and b, respectively, $K_{aD_{1}}$ is the photon’s Feynman propagator for photon a emitted by S1 goes to D1 and the meanings of other symbols are defined similarly.
-Since $\varphi_{a}$ and $\varphi_{b}$ are random for entangled photon pairs, the probability to detect one photon at D1 or D2 is a constant.
+where $\varphi_{a}$ and $\varphi_{b}$ are the initial phases of photons a and b, respectively, $K_{aD_{1}}$ is the photon’s Feynman propagator for photon a emitted by S1 goes to D1 and the meanings of other symbols are defined similarly.
+$\varphi_{a}$ and $\varphi_{b}$ are random for entangled photon pairs, the probability to detect one photon at D1 or D2 is a constant.
 The normalized two-photon probability function is proportional to the two-photon probability function, $P^{(2)}(\vec{r}_{1},t_{1};\vec{r}_{2},t_{2})$.
-Both temporal and spatial second-order interference can be calculated.
-First calculate the temporal case.
-Assume D1 and D2 are in the symmetrical positions about BS, all the properties such as polarization, frequency are the same for photons a and b.
+Both temporal and spatial second-order interference patterns can be calculated.
+Let us first calculate the temporal case.
+Assume D1 and D2 are in the symmetrical positions of the BS.
+All the properties such as polarization, frequency, are the same for photons a and b.
 With the same method as the one in Sect.
@@
 Point light sources are assumed to simplify the calculation.
-Figure 19: HOM dip for entangled photon pairs with different frequency bandwidths. 1, 2, and 3 correspond to $\Delta\omega$, $2\Delta\omega$, and $4\Delta\omega$, respectively.
-When the time difference between these two photon detection events is zero, $P^{(2)}(t_{1}-t_{2})$ equals zero, which means it is impossible to detect two photons at D1 and D2 simultaneously.
+Figure 19: Simulated HOM dip for entangled photon pairs with different frequency bandwidths. 1, 2, and 3 correspond to $\Delta\omega$, $2\Delta\omega$, and $4\Delta\omega$, respectively.
+When the time difference between these two photon detection events is zero, $P^{(2)}(t_{1}-t_{2})$ equals zero, which means that it is impossible to detect two photons at D1 and D2 simultaneously.
 These two photons are in the same output port of BS.
 When $|t_{1}-t_{2}|$ increases, $P^{(2)}(t_{1}-t_{2})$ increases from zero to one.
-$P^{(2)}(t_{1}-t_{2})$ equals 1 when $|t_{1}-t_{2}|$ is much longer than the coherence time, which indicates the detections of these two photons are independent.
-Similar method can be employed to calculate spatial two-photon probability function
+$P^{(2)}(t_{1}-t_{2})$ equals 1 when $|t_{1}-t_{2}|$ is much longer than the coherence time, which indicates that the detections of these two photons are independent.
+Similar method can be employed to calculate the spatial two-photon probability function,
 $P^{(2)}(x_{1}-x_{2})\propto 1-\text{sinc}^{2}\frac{d}{\lambda L}(x_{1}-x_{2}),$ (73)
 where $x_{1}$ and $x_{2}$ are the transverse coordinates of D1 and D2, respectively, $d$ is the size of one dimension source.
-Paraxial approximation and one-dimension are assumed to simplify the results.
+Paraxial approximation and one-dimension case are assumed to simplify the results.
 Experiments with entangled photon pairs confirmed the result in Eq.
@@
 (2019).
-It is interesting to comparing two-photon bunching of thermal light in a HBT interferometer and HOM dip of entangled photon pair in a HOM interferometer.
+It is interesting to compare two-photon bunching of thermal light in a HBT interferometer and HOM dip of entangled photon pair in a HOM interferometer.
 Two-photon bunching of thermal light shown in Fig.
-12 looks like the contrary effect of HOM dip shown in Fig.
+12 looks like a contrary effect of HOM dip shown in Fig.
 19.
 Photons have the tendency to bunch together when these two detectors are at symmetrical positions in Fig.
-12 while photons have the tendency to avoid each other when two detectors are at symmetrical positions in Fig.
+12.
+Photons have the tendency to leave the BS from the same output port when these two detectors are at symmetrical positions in Fig.
 19.
-Even through different phenomena are observed with types of light in different interferometers, the physics of these two phenomena are the same.
+Even through different phenomena are observed with different types of light in different interferometers, the physics of these two phenomena are the same.
 As shown in Eqs.
@@
 19.
-The same HOM dip can be observed with two independent and identical single-photon states as the input of HOM interferometer.
-The theoretical calculations are exactly the same as the one for entangled photon pairs and the results are experimentally verified by several groups Legero et al.
+The same HOM dip with 100% visibility can be observed with two independent and identical single-photon states.
+Theoretical calculations are exactly the same as the one for entangled photon pairs and the results are experimentally verified by several groups Legero et al.
 (2004); Beugnon et al.
@@
 There are three different ways to trigger a two-photon coincidence count in this case.
-The first one is these two photons are both emitted by S1.
+The first way is these two photons are both emitted by S1.
 The probability is 1/4.
@@
 The third way is the detected two photons are emitted by S1 and S2 one each.
-The probability is 1/2 and there are two different paths in this way.
+The probability is 1/2 and there are two different paths.
 The corresponding two-photon probability amplitudes are $A_{11,22}$ and $A_{12,21}$.
-If these four different paths are indistinguishable, the two-photon probability function is
+If these four different paths to trigger a two-photon coincidence count are indistinguishable, the two-photon probability function is
 $P^{(2)}(\vec{r}_{1},t_{1};\vec{r}_{2},t_{2})$ (74)
@@
 The last $1/\sqrt{4}$ in Eq.
-(74) is the product of $1/\sqrt{2}$ and $1/\sqrt{2}$, where the first $1/\sqrt{2}$ is due to the probability of these two paths equals 1/2 and the second $1/\sqrt{2}$ is the normalization parameter.
+(74) is the product of $1/\sqrt{2}$ times $1/\sqrt{2}$, where the first $1/\sqrt{2}$ is due to the probability of these two paths equals 1/2 and the second $1/\sqrt{2}$ is the normalization parameter.
 Table 1 summarizes the phases for these four different paths, where $\varphi_{1}$ and $\varphi_{2}$ are the initial phases of photons emitted by S1 and S2, respectively.
@@
 $=$ $\langle|K_{11,12}|^{2}\rangle+\langle|K_{21,22}|^{2}\rangle+\langle|K_{11,22}-K_{12,21}|^{2}\rangle.$
-Assuming S1 and S2 are point light sources, temporal two-photon probability function can be obtained via simplifying Eq. (IV.2.2)
+Assuming S1 and S2 are point laser light sources, temporal two-photon probability function can be obtained by simplifying Eq. (IV.2.2),
 $P^{(2)}(t_{1}-t_{2})$ $\propto$ $1+1+2[1-\cos\Delta\omega(t_{1}-t_{2})]$ (77)
@@
 There is second-order interference pattern by superposing two independent laser light beams.
-Unlike the transient first-order interference pattern of two independent laser light beams, the second-order interference pattern in Eq.
+Unlike the transient first-order interference pattern of two independent laser light beams is observed within the coherence time, the second-order interference pattern in Eq.
 (77) is observed over long time average.
-Figure 20: Temporal second-order interference pattern of two independent laser light beams in a HOM interferometer Liu et al. (2016)
-Figure 20 shows the temporal second-order interference pattern with two tunable single-mode continuous-wave lasers.
+Figure 20: Temporal second-order interference patterns of two independent laser light beams in a HOM interferometer Liu et al. (2016)
+Figure 20 shows the temporal second-order interference patterns with two tunable single-mode continuous-wave lasers.
 The frequency bandwidth of laser light is 100 kHz.
@@
 20(a)-(d), respectively.
-Periodic interference pattern is observed in Figs.
+Periodic interference patterns are observed in Figs.
 20(a)-(b).
@@
 20(c)-(d).
-The reason why it can not be observed is the period of the interference pattern is too small to observe in the present time scale Liu et al.
+The reason why it can not be observed is that the period of the interference pattern is too small to be observed in the time scale Liu et al.
 (2016).
 The same method can be employed to calculate the second-order interference of two independent thermal and laser light beams.
-Assume S1 is single-mode continuous-wave laser and S2 is thermal light source in Fig.
+Assume S1 is a single-mode continuous-wave laser and S2 is a thermal light source in Fig.
 18.
-The intensities, polarizations, frequencies of the light beams emitted by these two sources are the same.
+The intensities, polarizations, and frequencies of the light beams emitted by these two sources are the same.
 There are three different ways for D1 and D2 detecting a two-photon coincidence count.
-The first way is these two photons are emitted by laser S1.
+The first way is these two photons are emitted by laser, S1.
 The probability is 1/4 and there is one path, $A_{11,12}$.
-The second way is these two photons are emitted by thermal light source S2.
+The second way is these two photons are emitted by thermal light source, S2.
 The probability is 1/4 and there are two paths, $A_{2a1,2b2}$ and $A_{2a2,2b1}$.
@@
 The probability is 1/2 and there are two different paths, $A_{11,22}$ and $A_{12,21}$.
-If the five different paths are indistinguishable, the two-photon probability function equals
+If these five different paths are indistinguishable, the two-photon probability function equals
 $P^{(2)}(\vec{r}_{1},t_{1};\vec{r}_{2},t_{2})$ (78)
@@
 where $l$ is the length of thermal light source, $d$ is the distance between the central points of thermal light source and the virtual source of laser relative to BS Liu et al. (2014b).
-Figure 21: Second-order interference pattern of thermal and laser light (b) and the second-order coherence function of laser and thermal light beams in the interferometer alone (a) Liu et al. (2014b).
-Figure 21(b) shows the second-order interference pattern of thermal and laser light beams in a HOM interferometer, which is consistent with Eq.
+Figure 21: Second-order interference pattern of thermal and laser light (b) and the second-order coherence functions of laser and thermal light beams in the interferometer alone (a) Liu et al.
+(2014b).
+Figure 21(b) shows the spatial second-order interference pattern of thermal and laser light beams in a HOM interferometer, which is consistent with Eq.
 (IV.2.2).
@@
 (2014b).
-The second-order interference between two independent thermal light beams Liu et al.
-(2013), classical and nonclassical light beams Hong et al.
+The second-order interference of two independent thermal light beams Liu et al.
+(2013), classical and non-classical light beams Hong et al.
 (1988) and other two independent light beams can also be calculated in the same method above.
-The main difference between the second-order of two independent classical and nonclassical light beams is that the visibility of the second-order interference pattern of two independent nonclassicla light beams can reach 100%, while the maximal visibility of the second-order interference pattern between two independent classical light beam is 50% Mandel (1983).
-The reason is clearly shown in Eqs.
+The main difference between the second-order of two independent classical and non-classical light beams is that the visibility of the second-order interference pattern of two independent non-classical light beams can reach 100%, while the maximal visibility of the second-order interference pattern of two independent classical light beam is 50% Mandel (1983).
+The reason is clearly illustrated by Eqs.
 (70), (74), and (78).
-In the second-order interference of two independent classical light beams, it is impossible to distinguish one source emitting two photons from two sources emitting a photon each, which will contribute to the background coincidence counts.
-If special two-photon coincidence count system can distinguish these two types of coincidence count, the visibility of second-order interference pattern of two independent classical light beams can exceed 50%.
+In the second-order interference of two independent classical light beams, it is impossible to distinguish two sources emitting a photon each from one source emitting two photons, which will contribute to the background coincidence counts.
+If special two-photon coincidence count detection system can distinguish these two types of coincidence count, the visibility of the second-order interference pattern of two independent classical light beams can exceed 50%.
 For instance, Kaltenbaek et al.
-observed the visibility of second-order interference pattern with two independent laser light beams exceeding 85% by employing sum frequency generation to distinguish these two different types of coincidence count.
+observed the visibility of second-order interference pattern with two independent laser light beams exceeding 85% by employing sum frequency generation to distinguish these two different types of coincidence count Kaltenbaek et al.
+(2008).
 With two non-independent classical light beams and normal two-photon coincidence count system, Sadana et al.
-observed near 100% visibility second-order interference pattern Sadana et al.
+observed the second-order interference pattern with near 100% visibility Sadana et al.
 (2019).
-The results do not conflict with Mandel’s conclusion due to Mandel’s conclusion is valid for two independent light beams and normal two-photon coincidence count detection system Mandel (1983).
+Their results Sadana et al.
+(2019); Kaltenbaek et al.
+(2008) do not conflict with Mandel’s conclusion due to Mandel’s conclusion is valid for two independent light beams and normal two-photon coincidence count detection system Mandel (1983).
 IV.3 Second-order interference of multiple light beams
 Feyman’s path integral can be employed to calculate the second-order interference of more than two independent light beams.
-Figure 22 shows the second-order interference of three independent light beams.
+Figure 22 shows the scheme for the second-order interference of three independent light beams.
 S1, S2 and S3 are three independent light sources.
 D1 and D2 are two single-photon detectors.
-The output of two detectors are sent into a two-photon coincidence count system to measure the second-order correlation, which is not shown in the figure.
-Figure 22: Second-order interference of three independent light beams
+The output of these two detectors are sent into a two-photon coincidence count detection system to measure the second-order correlation, which is not shown in the figure.
+Figure 22: Second-order interference of three independent light beams.
+S1, S2 and S3 are three light sources.
+D1 and D2 are two single-photon detectors.
+The output of two detectors are sent into a two-photon coincidence count system, which is not shown.
 Assume S1, S2 and S3 are three independent and identical single-photon sources.
 There are three different ways to trigger a two-photon coincidence count.
-The first one is the two detected photons are emitted by S1 and S2.
+The first way is that the two detected photons are emitted by S1 and S2.
 The probability is 1/3.
 There are two different paths and the corresponding two-photon probability amplitudes are $A_{11,22}$ and $A_{12,21}$.
-The second way is these two photons are emitted by S1 and S3.
+The second way is that these two photons are emitted by S1 and S3.
 The probability is 1/3.
 There are two paths and the corresponding two-photon probability amplitudes are $A_{11,32}$ and $A_{12,31}$.
-The third way is these two photons are emitted by S2 and S3.
+The third way is that these two photons are emitted by S2 and S3.
 The probability is 1/3.
 There are two paths and the corresponding two-photon probability amplitudes are $A_{21,32}$ and $A_{22,31}$.
-If only one single-photon source emit a photon, it is impossible to trigger a two-photon coincidence count.
-If two single-photon sources emit photons simultaneously, these three different ways are distinguishable.
-One can measure the status of the single-photon source to ensure which two sources emitted photons.
-The probability for three independent single-photon sources simultaneously emitting photons is much smaller than the one of two single-photon sources simultaneously emitting photons, which will be ignored in the calculation.
+If only one single-photon source emits a photon, it is impossible to trigger a two-photon coincidence count.
+If two single-photon sources emit a photon simultaneously, these three different ways are distinguishable.
+One can measure the status of the single-photon sources to ensure which two sources emitted photons.
+The probability for three independent single-photon sources simultaneously emitting a photon is much smaller than the probability of two single-photon sources simultaneously emitting a photon.
+Hence three single-photon sources simultaneously emitting a photon can be ignored in the calculation.
 If every two different paths to trigger a two-photon coincidence count are indistinguishable, the two-photon probability function is
@@
 $+|\frac{1}{\sqrt{6}}(A_{21,32}+A_{22,31})|^{2}\rangle.$
-With the same method above, Eq. (81) can be simplified by considering the three single-photon sources are independent,
+With the same method above, Eq. (81) can be simplified as,
 $P^{(2)}(\vec{r}_{1},t_{1};\vec{r}_{2},t_{2})$ (82)
-$=$ $\langle|K_{11}K_{22}+K_{12}K_{21}|^{2}\rangle+\langle|K_{11}K_{32}+K_{12}K_{31}|^{2}\rangle$
+$\propto$ $\langle|K_{11}K_{22}+K_{12}K_{21}|^{2}\rangle+\langle|K_{11}K_{32}+K_{12}K_{31}|^{2}\rangle$
 $+\langle|K_{21}K_{32}+K_{22}K_{31}|^{2}\rangle,$
 where $K_{11}$ is the photon’s Feynman propagator from S1 to D1 and the meanings of other symbols are defined similarly.
-Assuming S1, S2 and S3 are point source and in the same source plane, the distance between the source and the observation planes is much larger than one between sources, the spatial two-photon probability function can be obtained via Eq.
-(82)
+Assuming S1, S2 and S3 are point single-photon sources in the same plane, the distance between the source and the observation planes is much larger than one between sources, the spatial two-photon probability function can be obtained via Eq.
+(82),
 $P^{(2)}(x_{1}-x_{2})$
@@
 $+\cos\frac{2\pi d_{13}}{L\lambda}(x_{1}-x_{2})+\cos\frac{2\pi d_{23}}{L\lambda}(x_{1}-x_{2}+\frac{d_{12}}{2}),$
-where paraxial approximation and one-dimension have been assumed, $\lambda$ is the wavelength of light and single-frequency is employed to simplify the calculation, $d_{12}$ is the distance between S1 and S2 and relative positions of these three sources are shown in Fig.
+where paraxial approximation and one-dimension case have been assumed, $\lambda$ is the wavelength of light and single-frequency light is employed to simplify the calculation, $d_{12}$ is the distance between S1 and S2 and the relative positions of these three sources are shown in Fig.
 8.
 The meanings of other symbols are defined similarly.
-Figure 23: Second-order interference pattern of three independent light beams. $x_{1}-x_{2}$ is the relative position of D1 and D2.
-Figure 23 shows the second-order interference pattern of three single-photon sources shown in Fig.
-22.
+Figure 23: Second-order interference patterns of three independent light beams. $x_{1}-x_{2}$ is the transverse position difference of D1 and D2.
+Figure 23 shows the second-order interference patterns of three single-photon sources.
 The patterns will change as the distances between these three sources change.
 Compared to the first-order interference pattern of three light beams shown in Fig.
-9, the visibility of the second-order interference pattern with three independent light beams can not reach 100% even when single-photon states are employed.
-The second-order interference of three independent laser light beams can also be calculated in the same method.
+9, the visibility of the second-order interference pattern with three independent light beams can not reach 100% even when non-classical light are employed.
+The second-order interference of three independent laser light beams can be calculated in the same method.
 Assume S1, S2 and S3 are three independent and single-mode continuous-wave lasers and the polarizations of these three laser light beams are identical.
@@
 22.
-The first way is these two photons are both emitted by S1.
+The first way is that these two photons are both emitted by S1.
 The probability is $I_{1}^{2}/(I_{1}+I_{2}+I_{3})^{2}$, where $I_{1}$, $I_{2}$, and $I_{3}$ are the intensities of laser light beams emitted by S1, S2 and S3, respectively.
 There is one path and the corresponding two-photon probability amplitude is $A_{11,12}$.
-The second ways is these two photons are both emitted by S2.
+The second ways is that these two photons are both emitted by S2.
 The probability is $I_{2}^{2}/(I_{1}+I_{2}+I_{3})^{2}$.
 There is one path and the two-photon probability amplitude is $A_{21,22}$.
-The third way is these two photons are both emitted by S3.
+The third way is that these two photons are both emitted by S3.
 The probability is $I_{3}^{2}/(I_{1}+I_{2}+I_{3})^{2}$.
 There is one path and the probability amplitude is $A_{31,32}$.
-The fourth way is these two photons are emitted by S1 and S2 one each.
+The fourth way is that these two photons are emitted by S1 and S2 one each.
 The probability is $2I_{1}I_{2}/(I_{1}+I_{2}+I_{3})^{2}$.
 There are two paths and the probability amplitudes are $A_{11,22}$ and $A_{12,21}$.
-The fifth way is these two photons are emitted by S1 and S3 one each.
+The fifth way is that these two photons are emitted by S1 and S3 one each.
 The probability is $2I_{1}I_{3}/(I_{1}+I_{2}+I_{3})^{2}$ and the two paths are $A_{11,32}$ and $A_{12,31}$.
-The sixth way is these two photons are emitted by S2 and S3 one each.
+The sixth way is that these two photons are emitted by S2 and S3 one each.
 The probability is $2I_{2}I_{3}/(I_{1}+I_{2}+I_{3})^{2}$ and the two paths are $A_{21,32}$ and $A_{22,31}$.
@@
 $+\sqrt{I_{1}I_{3}}(A_{11,32}+A_{12,31})+\sqrt{I_{2}I_{3}}(A_{21,32}+A_{22,31})|^{2}\rangle,$
-where the common factor $1/(I_{1}+I_{2}+I_{3})$ has been ignored, $\sqrt{I_{1}I_{2}}=\sqrt{2I_{1}I_{2}}\times 1/\sqrt{2}$.
+where the common factor $1/(I_{1}+I_{2}+I_{3})$ has been ignored, $\sqrt{I_{1}I_{2}}$ equals $\sqrt{2I_{1}I_{2}}\times 1/\sqrt{2}$.
 Considering three lasers are independent, Eq. (IV.3) can be simplified as
-$P^{(2)}(\vec{r}_{1},t_{1};\vec{r}_{2},t_{2})$
+$P^{(2)}(\vec{r}_{1},t_{1};\vec{r}_{2},t_{2})$ (85)
 $=$ $I_{1}^{2}\langle|K_{11}K_{12}|^{2}\rangle+I_{2}^{2}\langle|K_{21}K_{22}|^{2}\rangle+I_{3}^{2}\langle|K_{31}K_{32}|^{2}\rangle$
 $+I_{1}I_{2}\langle|K_{11}K_{22}+K_{12}K_{21}|^{2}\rangle$
-$+I_{1}I_{3}\langle|K_{11}K_{32}+K_{12}K_{31}|^{2}\rangle+I_{2}I_{3}\langle|K_{21}K_{32}+K_{22}K_{31}|^{2}\rangle.$
-Assuming the intensities of these three light beams are equal and single frequency light is emitted, Eq. (IV.3) can be simplified as
-$P^{(2)}(x_{1}-x_{2})$
+$+I_{1}I_{3}\langle|K_{11}K_{32}+K_{12}K_{31}|^{2}\rangle$
+$+I_{2}I_{3}\langle|K_{21}K_{32}+K_{22}K_{31}|^{2}\rangle.$
+Assuming the intensities of these three light beams are equal and single-frequency light is emitted, Eq. (85) can be simplified as
+$P^{(2)}(x_{1}-x_{2})$ (86)
 $\propto$ $4.5+\cos\frac{2\pi d_{12}}{L\lambda}(x_{1}-x_{2}-\frac{d_{23}}{2})$
-$+\cos\frac{2\pi d_{13}}{L\lambda}(x_{1}-x_{2})+\cos\frac{2\pi d_{23}}{L\lambda}(x_{1}-x_{2}+\frac{d_{12}}{2}),$
-where paraxial approximation and one-dimension are assumed to simplify the calculations.
-Comparing the second-order interference pattern of three independent single-photon states, Eq.
-(IV.3) and the second-order interference pattern of three independent single-mode continuous-wave lasers, the interference patterns are the same except the constant changes from 3 to 4.5.
-The reason for this difference is that laser light can not avoid two photons emitted from the same source with normal two-photon coincidence count detection system, which is exactly the same as the one for the difference between the visibility of the second-order interference pattern of two independent classical and non-classical light beams Mandel (1983).
+$+\cos\frac{2\pi d_{13}}{L\lambda}(x_{1}-x_{2})$
+$+\cos\frac{2\pi d_{23}}{L\lambda}(x_{1}-x_{2}+\frac{d_{12}}{2}),$
+where paraxial approximation and one-dimension case are assumed to simplify the calculations.
+The second-order interference patterns of three independent single-photon states and three independent single-mode continuous-wave lasers are the same except the constant changes from 3 to 4.5.
+The reason for this difference is that laser light can not avoid two photons emitted by the same source with normal two-photon coincidence count detection system, which is the same as the one for the difference between the visibilities of the second-order interference patterns of two independent classical and non-classical light beams Mandel (1983).
 IV.4 Subwavelength interference
@@
 IV, we have introduced how to employ Feynman’s path integral to calculate the second-order interference of one, two, and multiple light beams.
-As an example of how quantum optical coherence theory based on Feynman’s path integral will help us to understand the physics of optical coherence, subwavelength interference is analyzed in Feynman’s path integral.
-Taking the first- and second-order interference of light emitted by two independent sources shown in Fig.
+As an example of how quantum optical coherence theory based on Feynman’s path integral can be helpful to understand the physics of optical coherence, subwavelength interference of light is analyzed in Feynman’s path integral.
+Take the first- and second-order interference of light emitted by two independent sources shown in Fig.
 5 as an example to illustrate subwavelength interference.
@@
 $P^{(1)}(x)\propto 1+\cos(\frac{2\pi d}{\lambda L}x+\varphi).$ (87)
-Putting two detectors in the detection plane in Fig. 5, the second-order interference pattern of two independent single-mode continuous-wave lasers is proportional to
+Putting two detectors in the detection plane in Fig. 5, the second-order interference pattern of two independent single-mode continuous-wave lasers is
 $p^{(2)}(x_{1}-x_{2})\propto 1+\frac{1}{2}\cos\frac{2\pi d}{\lambda L}(x_{1}-x_{2}).$ (88)
@@
 (88) with Eq.
-(87) , if these two detectors are moving in the opposite direction with the same speed, the period of the second-order interference pattern will be half of the one of the first-order interference pattern, which is called subwavelength interference.
-Subwavelength interference can be understood better with the concept of de Broglie wavelength.
+(87) , if these two detectors are moving in the opposite directions with the same speed, the period of the second-order interference pattern will be half of the period of the first-order interference pattern, which is called subwavelength interference.
+Subwavelength interference can be understood with the concept of de Broglie wavelength.
 The de Broglie wavelength of a photon equals $h/p$, where $h$ is Planck constant and $p$ is the momentum of the photon.
 If two photons can be treated as a whole, the corresponding de Broglie wavelength equals $h/(2p)$, which is only half of the wavelength of one photon.
-The same goes for the de Broglie wavelength of $N$ photon equals $h/(Np)$, where $N$ is a positive integer larger than 1 Jacobson et al.
+The de Broglie wavelength of $N$ photon equals $h/(Np)$, where $N$ is a positive integer larger than 1 Jacobson et al.
 (1995).
@@
 (2001).
-Then it was found that subwavelength interference can be observed with thermal light Scarcelli et al.
+Then it was found that two-photon subwavelength interference can also be observed with thermal light Scarcelli et al.
 (2004); Xiong et al.
@@
 (2005) and laser light Liu and Zhang (2010).
-Figure 24: Scheme for subwavelength interference with two laser light beams Liu and Zhang (2010)
-Figure 24 shows the scheme for observing two-photon subwavelength with two independent laser light beams Liu and Zhang (2010), which is equivalent to a two point source interferometer.
-Two independent laser light beams was obtained by splitting a single-mode continuous-wave laser light beam by a beam splitter (BS1) and randomize the phase of one beam by a mirror mounted on a piezoelectric ceramic (PM).
+Figure 24: Scheme for two-photon subwavelength interference with two laser light beams Liu and Zhang (2010).
+Laser is a single-mode continuous-wave laser.
+M is a mirror.
+PM is mirror mounted on a piezoelectric ceramic.
+SA and SB are two virtual sources of source S of M and PM, respectively.
+The meanings of other symbols are similar as the ones above.
+Figure 24 shows the scheme for observing two-photon subwavelength interference with two independent laser light beams Liu and Zhang (2010), which is equivalent to a two point-source interferometer.
+Two independent laser light beams was obtained by splitting a single-mode continuous-wave laser light beam with a beam splitter (BS1) and randomizing the phase of one beam by a mirror mounted on a piezoelectric ceramic (PM).
 One beam splitter (BS2), two single-photon detectors (D1 and D2), and a two-photon coincidence count detection system (CC) are employed to measure the second-order interference pattern.
-Figure 25: The observed normal second-order interference pattern Liu and Zhang (2010).
+Figure 25: The observed normal second-order interference pattern (b) when the position of D1 is scanning why the position of D2 is fixed Liu and Zhang (2010).
+R1 and R2 in (a) are single-photon counting rates of D1 and D2, respectively.
 Figure 25(b) shows the normal second-order interference pattern by fixing the position of D2 and scanning the position of D1.
@@
 Figure 25(a) shows the single-photon counting rates observed by two single-photon detectors, respectively.
-There is no first-order interference pattern observed.
-Figure 26: Two-photon subwavelength interference pattern Liu and Zhang (2010).
-Figure 26(b) shows the observed two-photon subwavelength interference pattern by by scanning the positions of D1 and D2 in the opposite directions.
+There is no first-order interference pattern.
+Figure 26: Two-photon subwavelength interference pattern (b) when the positions of D1 and D2 are scanned in the opposite directions Liu and Zhang (2010).
+R1 and R2 in (a) are single-photon counting rates of D1 and D2, respectively.
+Figure 26(b) shows the observed two-photon subwavelength interference pattern by scanning D1 and D2 in the opposite directions.
 The period of the second-order interference pattern is 400 $\mu$m, which is exactly half of the one shown in Fig.
-25(b) and indicates subwavelength interference is observed.
+25(b) and indicates two-photon subwavelength interference is observed.
 The visibility of the second-order interference pattern equals $(48\pm 4)\%$.
-One may argue that it is meaningless to observe subwavelength by scanning two detectors in the opposite directions Liu et al.
+One may argue that it is meaningless to observe subwavelength interference by scanning two detectors in the opposite directions Liu et al.
 (2014c).
-However, one can obtain subwavelength interference with two detectors scanning in the same direction for classical light with suitable changes on the experimental scheme Cao et al.
+However, two-photon subwavelength interference with two detectors scanning in the same direction for classical light can be observed with suitable changes on the experimental scheme Cao et al.
 (2010), which is similar as the one with entangled photon pairs Fonseca et al.
@@
 (2001).
-As shown below, there is no such big difference between the observed subwavelength interference patterns for detectors scanning in the same direction and opposite directions.
-In order to see the physics of subwavelength interference, take the second-order interference of two single-mode continuous-wave laser with the same phase as an example.
-Assume the intensities and polarizations of these two laser light beams are the same.
+As shown below, there is no intrinsic difference between the observed subwavelength interference patterns for detectors scanning in the same direction and opposite directions.
+In order to understand the physics of subwavelength interference, take the second-order interference of two single-mode continuous-wave laser light beams with the same phase as an example.
+Assume the intensities, frequency spectra and polarizations of these two laser light beams are the same.
 There are four different paths to trigger a two-photon coincidence count in Fig.
@@
 $P^{(2)}(x_{1};x_{2})$ (89)
-$=$ $\langle|A_{11,12}+A_{21,22}+A_{11,22}+A_{12,21}|^{2}\rangle$
-$=$ $|e^{i2\varphi}(K_{11}K_{12}+K_{21}K_{22}+K_{11}K_{22}+K_{12}K_{21}|^{2}\rangle$
+$=$ $\langle|\frac{1}{\sqrt{4}}[A_{11,12}+A_{21,22}+A_{11,22}+A_{12,21}]|^{2}\rangle$
+$\propto$ $|e^{i2\varphi}(K_{11}K_{12}+K_{21}K_{22}+K_{11}K_{22}+K_{12}K_{21}|^{2}\rangle$
 $=$ $|K_{11}K_{12}+K_{21}K_{22}+K_{11}K_{22}+K_{12}K_{21}|^{2},$
@@
 where $I_{1}=|K_{11}|^{2}=|K_{21}|^{2}$ and $I_{2}=|K_{12}|^{2}=|K_{22}|^{2}$, which equals half of the intensities detected by D1 and D2, respectively.
-Assume point light sources and single-frequency light beams are employed, Eq.
+Assume point laser light sources are employed, Eq.
 (90) can be simplified as Liu and Zhang (2010)
@@
 The last two lines on the right hand side of Eq.
-(91) corresponds to two-photon subwavelength when two detectors scanning in the same direction and opposition directions, respectively.
+(91) correspond to two-photon subwavelength when two detectors scanning in the same direction and opposition directions, respectively.
 $2I_{1}I_{2}\cos\frac{2\pi d}{\lambda L}(x_{1}+x_{2})$ corresponds to $K_{11}K_{12}K_{21}^{*}K_{22}^{*}+K_{21}K_{22}K_{11}^{*}K_{12}^{*}$ in Eq.
@@
 (90), which is a result of two-photon interference for these two photons emitted by one source each.
-If one can eliminate other terms and only keep one of the last two terms in Eq.
+If one can eliminate all the other terms and keep only one of the last two terms in Eq.
 (91), one can observe subwavelength interference.
@@
 (91).
-Subwavelength interference is observed when two detectors are scanning in the opposite directions.
-Equations (90) and (91) shows that there is no intrinsic difference between the subwavelength interference when these two detectors scanning in the same direction and opposite directions.
+Subwavelength interference is observed when these two detectors are scanning in the opposite directions.
+Equations (90) and (91) shows that there is no intrinsic difference between the subwavelength interference when these two detectors are scanning in the same direction and opposite directions.
 They are results of two-photon interference with different two-photon paths.
 V Third- and higher-order interference of light
-As mentioned above, the third- and higher-order interference of light is an natural extension of the second-order interference of light.
+The third- and higher-order interference of light is a natural generalization of the second-order interference of light.
 The method introduced above can also be employed to calculate the third- and higher-order interference of light.
-The studies of third- and higher-order interference dates back to 1963 when Glauber gave a unified definition of $N$th-order coherence function of light Glauber (1963a, c).
+The history of the third- and higher-order interference dates back to 1963 when Glauber gave a unified definition of $N$th-order coherence function of light Glauber (1963a, c).
 Greenberg et al.
@@
 (1989, 1990).
-It is difficult to generate entangled three-photon directly.
+It was difficult to generate entangled three-photon directly then.
 Bouwmeester et al.
@@
 (2005) via the second-order interference of light.
-Ghost imaging via the third- and higher-order interference of light was discussed for entangled three-photon Wen et al.
+Ghost imaging with the third- and higher-order interference of light was discussed for entangled three-photon Wen et al.
 (2007) and thermal light Ou and Kuang (2007); Bai and Han (2007); Agafonov et al.
@@
 (2009).
-In the following, third-order interference of light is taken as an example to show how to employ Feyman’s path integral theory to calculate the third- and high-order interference of light.
+In the following, third-order interference of light is taken as an example to show how to employ Feyman’s path integral to calculate the third- and high-order interference of light.
 V.1 Third-order interference of one light beam
-Figure 27 shows the third-order HBT interferometer by analogy of HBT interferometer shown in Fig.
+Figure 27 shows the scheme for the third-order HBT interferometer by analogy of the HBT interferometer shown in Fig.
 11.
-The light emitted by light source (S) is split into two via a 1:2 non-polarizing beam splitter BS1.
-The transmitted light beam is split into two via a 1:1 non-polarizing beam splitter.
+The light emitted by a light source, S, is split into two via a 1:2 non-polarizing beam splitter, BS1.
+The transmitted light beam is split into two via a 1:1 non-polarizing beam splitter, BS2.
 The intensities detected by D1, D2, and D3 are equal.
-TCC is short for three-photon coincidence count and is usually written as CC for convenience.
+TCC is short for three-photon coincidence count detection system and usually written as CC for convenience.
 The distance between the source and detector planes are equal.
@@
 TCC is short for three-photon coincidence count and is usually written as CC for convenience.
-The meanings of other symbols are the same as the one in HBT interferometer.
-Assume S is thermal light source in Fig.
+BS1 is a 1:2 non-polarizing beam splitter.
+The meanings of other symbols are the same as the ones in a usual HBT interferometer.
+Assume S is a thermal light source in Fig.
 27.
-There are six different paths to trigger a three-photon coincidence count.
-Figure 28 shows the six different ways.
-For instance, Fig.
-28(a) show the path for photon a is detected by D1, photon b is detected by D2 and photon c is detected by D3.
-If all the six different ways are indistinguishable, the three-photon probability function is
+There are six different paths to trigger a three-photon coincidence count in D1, D2, and D3.
+Figure 28 shows the schematic diagram of the six different paths.
+Fig.
+28(a) shows the path for that photon a is detected by D1, photon b is detected by D2 and photon c is detected by D3.
+The corresponding three photon probability amplitude is $A_{a1,b2,c3}$.
+Other five paths are defined similarly.
+If all the six different paths are indistinguishable, the three-photon probability function is
 $P^{(3)}(\vec{r}_{1},t_{1};\vec{r}_{2},t_{2};\vec{r}_{3},t_{3})$ (92)
-$=$ $\langle|A_{a1,b2,c3}+A_{a1,b3,c2}+A_{a2,b1,c3}$
-$+A_{a2,b3,c1}+A_{a3,b1,c2}+A_{a3,b2,c1}|^{2}\rangle.$
+$=$ $\langle|\frac{1}{\sqrt{6}}(A_{a1,b2,c3}+A_{a1,b3,c2}+A_{a2,b1,c3}$
+$+A_{a2,b3,c1}+A_{a3,b1,c2}+A_{a3,b2,c1})|^{2}\rangle.$
 The initial phases of photons in thermal light are random, Eq. (92) can be simplified as
 $P^{(3)}(\vec{r}_{1},t_{1};\vec{r}_{2},t_{2};\vec{r}_{3},t_{3})$ (93)
-$=$ $\langle|K_{a1}K_{b2}K_{c3}+K_{a1}K_{b3}K_{c2}+K_{a2}K_{b1}K_{c3}$
+$\propto$ $\langle|K_{a1}K_{b2}K_{c3}+K_{a1}K_{b3}K_{c2}+K_{a2}K_{b1}K_{c3}$
 $+K_{a2}K_{b3}K_{c1}+K_{a3}K_{b1}K_{c2}+K_{a3}K_{b2}K_{c1}|^{2}\rangle,$
 where $K_{a1}$ is the photon’s Feynman propagator from S to D1.
-Figure 28: Third-order HBT interferometer.
-TCC is short for three-photon coincidence count and is usually written as CC for convenience.
-The meanings of other symbols are the same as the one in HBT interferometer.
-With the same method as the one for thermal light in a HBT interferometer introduced in Sect. IV.1.1, the spatial three-photon probability function is Liu and Shih (2009)
+Figure 28: Six different paths to trigger a three-photon coincidence count.
+(a) shows the path for photon a is detected by D1, photon b is detected by D2 and photon c is detected by D3.
+Other five paths are defined similarly.
+With the same method as the one for thermal light in a HBT interferometer in Sect. IV.1.1, the spatial three-photon probability function is Liu and Shih (2009)
 $P^{(3)}(x_{1};x_{2};x_{3})$
@@
 $+2\text{sinc}\frac{\pi d}{L\lambda}(x_{1}-x_{2})\text{sinc}\frac{\pi d}{L\lambda}(x_{2}-x_{3})\text{sinc}\frac{\pi d}{L\lambda}(x_{3}-x_{1}),$
-where paraxial approximation and one-dimension have been employed, $\lambda$ is the wavelength of thermal light and single frequency light is assumed, $d$ is the length of one-dimension thermal light source, $L$ is the distance between the source and detector planes.
+where paraxial approximation and one-dimension case were assumed, $\lambda$ is the wavelength of thermal light and single-frequency light is employed, $d$ is the length of the one-dimension thermal light source, $L$ is the distance between the source and detector planes.
 Figure 29(a) shows the simulated spatial third-order coherence function of thermal light in a third-order HBT interferometer.
@@
 The theoretical and experimental results are consistent with each other.
-Figure 29: Simulated Liu and Shih (2009) and measured Zhou et al. (2010) spatial third-order coherence functions of thermal light in the third-order interferometer.
-Temporal three-photon probability function can be obtained in the same way Zhou et al. (2010)
+Figure 29: Simulated spatial third-order coherence function Liu and Shih (2009) and measured three-photon coincidence counts Zhou et al.
+(2010) of pseudothermal light in a third-order HBT interferometer.
+$x1-x3$ in (a) and $x_{13}$ in (b) represent transverse position difference of D1 and D3.
+Similar definition holds for $x1-x2$ in (a) and $x_{23}$ in (b).
+Triple C.C.
+in (b) is three-photon coincidence counts, which is proportional to the normalized third-order coherence function in (a).
+Temporal three-photon probability function can be calculated in the same way Zhou et al. (2010),
 $P^{(3)}(t_{1};t_{2};t_{3})$
@@
 where $\Delta\omega$ is the frequency bandwidth of thermal light and point thermal light source is assumed.
-The degree of third order coherence, $g^{(3)}(0)$, of thermal light equals 6, which means that the probability to detect three photons when these three detectors are in the same space-time coordinate is 6 times of the one when these three detectors are in different coherence volumes of thermal light.
-It is three-photon bunching of thermal light.
+The degree of third order coherence, $g^{(3)}(0)$, of thermal light equals 6, which means that the probability of detecting three photons when these three detectors are in the same space-time coordinate is 6 times of the one when these three detectors are in different coherence volumes.
+It is defined as three-photon bunching of thermal light.
+The degree of $N$th-oder coherence of thermal light equals $N!$ Liu and Shih (2009).
 Three-photon superbunching can be observed with superbunching pseudothermal light Zhou et al.
-(2019a) and the degree of $N$th-oder coherence of thermal light equals $N!$ Liu and Shih (2009).
-The third-order interference of single-mode continuous-wave laser light in the third-order HBT interferometer is simple.
+(2019a).
+The third-order interference of single-mode continuous-wave laser light in a third-order HBT interferometer is simple.
 All the six different paths shown in Fig.
@@
 The superposition of these six probability amplitudes will not give three-photon interference.
-The reason is the same as the one for there is no two-photon interference for single-mode continuous-wave laser light in a HBT interferometer.
-No three-photon bunching can be observed when single-mode continuous-wave laser light in the third-order interferometer.
+The reason is the same as the one for there is no two-photon bunching for single-mode continuous-wave laser light in a HBT interferometer.
+No three-photon bunching for single-mode continuous-wave laser light exists in the third-order HBT interferometer.
 V.2 Third-order interference of two independent light beams
@@
 30.
-The first way is one photon is emitted by S1 and two photons are emitted by S2.
-The probability is $3I_{1}I_{2}^{2}/(3I_{1}I_{2}^{2}+I_{2}^{3})$, where $I_{1}$ and $I_{2}$ are the intensities emitted by S1 and S2, respectively.
+The first way is that one photon is emitted by S1 and two photons are emitted by S2.
+The probability is $3I_{1}I_{2}^{2}/(3I_{1}I_{2}^{2}+I_{2}^{3})$, where $I_{1}$ and $I_{2}$ are the intensities of light beams emitted by S1 and S2, respectively.
 There are three different paths and the corresponding three-photon probability amplitudes are $A_{11,22,23}$, $A_{12,21,23}$, and $A_{13,21,22}$, respectively.
@@
 The probability is $I_{2}^{3}/(I_{1}I_{2}^{2}+I_{2}^{3})$.
-There is one path and the corresponding three-photon probability is $A_{21,22,23}$.
+There is one path and the corresponding three-photon probability amplitude is $A_{21,22,23}$.
 Figure 30: Scheme for the third-order interference of two independent light beams.
+S1 and S2 are two light sources.
+D1, D2 and D3 are three single-photon detectors.
+The output signals of D1, D2 and D3 are sent into a three-photon coincidence count detection system, which is not shown.
 If these four different paths are indistinguishable, the three-photon probability function is
@@
 $+\sqrt{\frac{I_{2}^{3}}{3I_{1}I_{2}^{2}+I_{2}^{3}}}A_{21,22,23}|^{2}\rangle.$
-The initial phases of photons emitted by S1 are random and emitted by S2 are identical within one coherence volume, Eq. (96) can be simplified as
+The initial phases of photons emitted by the single-photon source, S1, are random.
+The initial phases of photons emitted by the single-mode continuous-wave laser, S2, are identical within one coherence volume.
+Equation (96) can be simplified as
 $P^{(3)}(\vec{r}_{1},t_{1};\vec{r}_{2},t_{2};\vec{r}_{3},t_{3})$ (97)
@@
 There is third-order interference pattern.
-The visibility of the third-order interference pattern is dependent on the ratio between the intensities of light emitted by single-photon source and single-mode continuous-wave laser.
+The visibility of the third-order interference pattern is dependent on the ratio between the intensities of light emitted by the single-photon source and single-mode continuous-wave laser.
 Similar method can be employed to calculate the third-order interference of two other types of light beams.
@@
 Assume S1, S2, and S3 are three identical and independent single-photon sources.
-The intensities, polarizations, and the frequencies are the same for the photons emitted by these three sources.
+The intensities, polarizations, and frequency spectra of the photons emitted by these three sources are the same.
 There are six different paths to trigger a three-photon coincidence count.
-The corresponding three probability amplitudes are $A_{11,22,33}$, $A_{11,23,31}$, $A_{12,21,33}$, $A_{12,23,31}$, $A_{13,22,31}$, and $A_{13,21,32}$, where $A_{11,22,33}$ is the probability amplitude for photon emitted by S1 is detected by D1, photon emitted by S2 is detected by D2, and photon emitted by S3 is detected by D3.
+The corresponding three-photon probability amplitudes are $A_{11,22,33}$, $A_{11,23,31}$, $A_{12,21,33}$, $A_{12,23,31}$, $A_{13,22,31}$, and $A_{13,21,32}$, where $A_{11,22,33}$ is the probability amplitude for that the photon emitted by S1 is detected by D1, the photon emitted by S2 is detected by D2, and the photon emitted by S3 is detected by D3.
 The meanings of other symbols are defined similarly.
-Figure 31: Scheme for the third-order interference of three independent light beams
-If six different paths are indistinguishable, three-photon probability function is
+Figure 31: Scheme for the third-order interference of three independent light beams. The meanings of the symbols are similar as the ones in Fig. 30.
+If all the six different paths are indistinguishable, three-photon probability function is
 $P^{(3)}(\vec{r}_{1},t_{1};\vec{r}_{2},t_{2};\vec{r}_{3},t_{3})$ (99)
-$=$ $\langle|A_{11,22,33}+A_{11,23,31}+A_{12,21,33}$
-$+A_{12,23,31}+A_{13,22,31}+A_{13,21,32}|^{2}\rangle$
-$=$ $\langle|e^{i(\varphi_{1}+\varphi_{2}+\varphi_{3})}(K_{11}K_{22}K_{33}+K_{11}K_{23}K_{31}+K_{12}K_{21}K_{33}$
-$+K_{12}K_{23}K_{31}+K_{13}K_{22}K_{31}+K_{13}K_{21}K_{32})|^{2}\rangle,$
+$=$ $\langle|\frac{1}{\sqrt{6}}(A_{11,22,33}+A_{11,23,31}+A_{12,21,33}$
+$+A_{12,23,31}+A_{13,22,31}+A_{13,21,32})|^{2}\rangle$
+$\propto$ $\langle|e^{i(\varphi_{1}+\varphi_{2}+\varphi_{3})}(K_{11}K_{22}K_{33}+K_{11}K_{23}K_{31}$
+$+K_{12}K_{21}K_{33}+K_{12}K_{23}K_{31}$
+$+K_{13}K_{22}K_{31}+K_{13}K_{21}K_{32})|^{2}\rangle,$
 where $\varphi_{1}$, $\varphi_{2}$, and $\varphi_{3}$ are the initial phases of photons emitted by S1, S2, and S3, respectively.
-With the same method and approximations above, one-dimension spatial three-photon probability function is
+With the same method and approximations above, one-dimension spatial three-photon probability function can be simplified as
 $P^{(3)}(x_{1};x_{2};x_{3})$
@@
 $P^{(3)}(x_{1};x_{2};x_{3}=x_{2})$
-$=$ $1+\cos\frac{2\pi d}{L\lambda}(x_{1}-x_{2})$
+$\propto$ $1+\cos\frac{2\pi d}{L\lambda}(x_{1}-x_{2})$
 $+\frac{1}{3}\cos\frac{4\pi d}{L\lambda}x_{1}\cos\frac{2\pi d}{L\lambda}(x_{2}-\frac{d}{2})\cos\frac{2\pi d}{L\lambda}(x_{2}+\frac{d}{2})$
@@
 $+\frac{1}{3}\cos\frac{4\pi d}{L\lambda}x_{2}\cos\frac{2\pi d}{L\lambda}(x_{2}-\frac{d}{2})\cos\frac{2\pi d}{L\lambda}(x_{1}+\frac{d}{2}).$
-Figure 32 shows the third-order interference pattern given by Eq. (V.3).
-Figure 32: Third-order interference pattern of three independent single-photon states.
-Similar method can be employed to calculate the third-order interference other three types of light beams and higher-order interference of light.
+Figure 32 shows the third-order interference pattern given by Eq. (V.3). The periodic structures can be observed in $x_{1}$, $x_{2}$, and $x_{1}-x_{2}$ directions.
+Figure 32: Third-order interference pattern of three independent single-photon states. $x_{1}$ and $x_{2}$ are the transverse positions of D1 and D2, respectively.
+Similar method can be employed to calculate the third-order interference other types of light beams and higher-order interference of light.
+In fact, all optical coherence problems can be calculated in the same method introduced above.
 VI Quantum atomic coherence theory based on Feynman’s path integral
+It is straightforward to generalize quantum optical coherence theory based on Feynman’s path integral to discuss the coherence of massive particles.
 The history of atomic coherence traces back to de Broglie in 1924 when he proposed matter wave assumption De Broglie (1924) by analogy of Planck’s Planck (1901) and Einstein’s Einstein (1905) quanta assumption of light.
@@
 with diffraction pattern of electrons Davisson and Germer (1927); Thomson and Reid (1927).
-Due to the very short de Broglie wavelength of matter wave, it is difficult to observe the first-order interference pattern like the one of light.
+Due to the very short de Broglie wavelength of matter wave, it was difficult to observe the first-order interference pattern like the one of light then.
 In 1953, Marton et al.
@@
 (1999) were also reported.
-The first-order interference pattern of two independent Bose Einstein Condensations (BEC) was also reported Andrews et al.
+The first-order interference pattern of two independent Bose Einstein Condensations (BEC) was observed Andrews et al.
 (1997).
-The first second-order interference experiments with matter wave was done by Yasuda and Shimizu, in which they observed two-atom bunching with cold atoms Yasuda and Shimizu (1996).
+The first experiment of the second-order interference with matter wave was done by Yasuda and Shimizu, in which they observed two-atom bunching with cold atoms Yasuda and Shimizu (1996).
 Burt et al.
-observed the degree of third-order coherence of cold atoms and BEC, finding that the ratio between these two equals $7.4\pm 2.6$ Burt et al.
+found that the ratio between the degrees of third-order coherence of cold atoms and BEC equals $7.4\pm 2.6$ Burt et al.
 (1997).
-The second-order interference of fermions was also reported by several groups Henny et al.
+The second-order interference of fermions was reported by several groups Henny et al.
 (1999); Oliver et al.
@@
 (2010, 2011).
-Storey and Cohen-Tannoudji provided a good introduction about how to employ Feynman’s path integral to calculate the first-order interference of atoms Storey and Cohen-Tannoudji (1994).
+Storey and Cohen-Tannoudji provided an introduction of how to employ Feynman’s path integral to calculate the first-order interference of atoms Storey and Cohen-Tannoudji (1994).
 Employing Feynman’s path integral to calculate the first-order interference of matter wave can be found in Refs.
 Deng (2006, 2008); Gondran and Gondran (2005), too.
-This section will summarize how to employ Feynman’s path integral to calculate the first-, second, and higher-order interference matter waves, with special attentions on the analogy between cold atoms (BEC) and thermal light (laser light).
+This section will focus on how to employ Feynman’s path integral to calculate the first-, second, and higher-order interference of matter waves, with special attentions on the analogy between cold atoms (BEC) and thermal light (laser light).
 VI.1 First-order interference of matter wave
@@
 O is the observation plane.
-For simplicity, free atoms are assumed.
-Figure 33: Scheme for the first-order interference of one matter wave beam.
+For simplicity, free atoms are assumed here and in the following discussions.
+Figure 33: Scheme for the first-order interference of one matter wave beam. S is a particle source. P1 and P2 are two pinholes. O is the observation plane.
 There are two different paths to trigger a particle detection event in the observation plane.
-The first one is the particle going through P1.
+The first path is the particle going through P1.
 The probability is 1/2 and the probability amplitude is $A_{S1x}$.
@@
 The probability is 1/2 and the probability amplitude is $A_{S2x}$.
-If these two different paths are indistinguishable, the probability function in the observation plane is
-$P^{(1)}(x,t)$ $=$ $\langle|A_{S1x}+A_{S2x}|^{2}\rangle$ (103)
-$=$ $\langle|e^{i\varphi_{j}}K_{S1x}+e^{i\varphi_{j}}K_{S2x}|^{2}\rangle$
+If these two different paths are indistinguishable, the probability function of particles in the observation plane is
+$P^{(1)}(x,t)$ $=$ $\langle|\frac{1}{\sqrt{2}}(A_{S1x}+A_{S2x})|^{2}\rangle$ (103)
+$\propto$ $\langle|e^{i\varphi_{j}}K_{S1x}+e^{i\varphi_{j}}K_{S2x}|^{2}\rangle$
 $=$ $\langle|K_{S1}K_{1x}+K_{S2}K_{2x}|^{2}\rangle,$
-where $\varphi_{j}$ is the initial phase of the $j$th detected particle, $K_{S1x}$ is the particle’s Feynman propagator from S to $x$ in the observation plane via P1, the meanings of other symbols are defined similarly.
-Particle’s Feynman propagator can be calculated in the same method as the one for photon and has simpler form for free particles Feynman et al. (2010)
+where $\varphi_{j}$ is the initial phase of the $j$th detected particle, $K_{S1x}$ is the particle’s Feynman propagator from S to $x$ via P1, the meanings of other symbols are defined similarly.
+Massive particle’s Feynman propagator can be calculated in the same method as the one for photon and has a simple form for free particles Feynman et al. (2010)
 $K(x_{a},t_{a};x_{b},t_{b})=\sqrt{\frac{m}{2\pi i\hbar(t_{b}-t_{a})}}\exp\frac{im(x_{b}-x_{a})^{2}}{2\hbar(t_{b}-t_{a})},$ (104)
-where $m$ is the mass of particle, $\hbar=h/(2\pi)$, $(x_{a},t_{a})$ and $(x_{b},t_{b})$ are the starting and ending space-time coordinates, respectively. For simplicity, one-dimension is considered.
-Assuming S is in the symmetrical position of P1 and P2, there is no path difference from S to two pinholes.
-The path difference is introduced by the pinholes to the observation plane.
+where $m$ is the mass of particle, $\hbar=h/(2\pi)$, $(x_{a},t_{a})$ and $(x_{b},t_{b})$ are the starting and ending space-time coordinates of the particle, respectively.
+For simplicity, one-dimension case is assumed.
+Assuming S is at the symmetrical position of P1 and P2, there is no path difference from S to two pinholes.
+The path difference is introduced by the distances from the pinholes to the observation plane.
 Substituting Eq.
@@
 $P^{(1)}(x)\propto 1+\cos\frac{2\pi d}{L\lambda_{D}}x,$ (105)
-in which paraxial approximation has been employed, $\lambda_{D}=h/(mv)$ is the de Broglie wavelength of particle, $v$ is the speed of atom and the speed of all atoms are assumed to be the same, $d$ is the distance between two pinholes, and $L$ is the distance between the pinhole and detection planes.
-There is first-order interference pattern when these two different paths are indistinguishable for particles, no matter these particles are fermions or bosons, in cold atom state or BEC state.
-Just like there is no difference for the first-order interference pattern of photons in a Young’s double-slit interferometer for thermal and laser light.
-The initial phases of particles does not contribute to the first-order interference pattern of one beam.
+in which paraxial approximation has been employed, $\lambda_{D}=h/(mv)$ is the de Broglie wavelength of the particle, $v$ is the speed of the particle and the speed of all the particles are assumed to be the same, $d$ is the distance between these two pinholes, and $L$ is the distance between the pinhole and observation planes.
+There is first-order interference pattern when these two different paths are indistinguishable.
+The patterns are the same for fermions and bosons.
+Just like there is no difference for the first-order interference patterns of thermal and laser light in a Young’s double-slit interferometer, the first-order interference patterns of bosons in thermal state or BEC state in a Young’s double-slit interferometer are the same.
+The initial phases of the particles do not influence the first-order interference pattern of one beam.
 VI.1.2 First-order interference of two independent matter wave beams
 Figure 34 shows the scheme for the first-order interference of two independent matter wave beams. S1 and S2 are two independent particle sources. O is the observation plane.
-Figure 34: Scheme for the first-order interference of two independent matter wave beams.
+Figure 34: Scheme for the first-order interference of two independent matter wave beams. S1 and S2 are two particle sources.
 There are two different paths to trigger a particle detection event in the observation plane.
-The first one is the detected particle is emitted S1.
+The first one is that the detected particle is emitted S1.
 The probability is 1/2 and the probability amplitude is $A_{1x}$.
-The second one is the detected particle is emitted S2.
+The second one is that the detected particle is emitted S2.
 The probability is 1/2 and the probability amplitude is $A_{2x}$.
-If these two different paths are indistinguishable, the probability function in the observation plane is
-$P^{(1)}(x,t)$ $=$ $\langle|A_{1x}+A_{2x}|^{2}\rangle$ (106)
-$=$ $\langle|e^{i\varphi_{1j}}K_{1x}+e^{i\varphi_{2j}}K_{2x}|^{2}\rangle,$
+If these two different paths are indistinguishable, the probability function of detecting a particle in the observation plane is
+$P^{(1)}(x,t)$ $=$ $\langle|\frac{1}{\sqrt{2}}(A_{1x}+A_{2x})|^{2}\rangle$ (106)
+$\propto$ $\langle|e^{i\varphi_{1j}}K_{1x}+e^{i\varphi_{2j}}K_{2x}|^{2}\rangle,$
 where $\varphi_{1j}$ and $\varphi_{2j}$ are the initial phases of particles emitted by S1 and S2, respectively.
-With the same approximations above, spatial probability function can be obtained via Eq.
+With the same approximations above, the spatial probability function can be obtained via Eq.
 (106),
 $P^{(1)}(x)$ $=$ $\langle 1+\cos[\frac{2\pi d}{L\lambda_{D}}x+(\varphi_{1j}-\varphi_{2j})]\rangle.$ (107)
-By analogy of the experimental results of the degree of second- and third-order coherence of cold atoms and BEC to thermal and laser light Yasuda and Shimizu (1996); Burt et al.
+The interference pattern is depend on the initial phases of particles.
+Comparing the experimental results of the degrees of second- and third-order coherence of cold atoms and BEC to the ones of thermal and laser light Yasuda and Shimizu (1996); Burt et al.
 (1997); Lopes et al.
-(2015), it is reasonable to assume the coherence properties of cold atoms are similar as the ones of thermal light and the coherence properties of BEC are similar as the one of laser light.
+(2015), it is reasonable to assume that the coherence properties of cold atoms are similar as the ones of thermal light and the coherence properties of BEC are similar as the ones of laser light.
 The initial phases of particles in cold atoms are random.
@@
 where $N$ is the number of detected particles and $\varphi$ is a random phase due to the sum of $\varphi_{1j}-\varphi_{2j}$ for $N$ detected particles.
-The first-order interference pattern can not observed.
+The first-order interference pattern of two independent cold atom sources can not be observed.
 For the superposition of of two independent BECs, the initial phases of particles in one BEC are identical. Equation (107) is simplified as
-$P^{(1)}(x)$ $=$ $1+\cos[\frac{2\pi d}{L\lambda_{D}}x+(\varphi_{1j}-\varphi_{2j})].$ (109)
+$P^{(1)}(x)$ $\propto$ $1+\cos[\frac{2\pi d}{L\lambda_{D}}x+(\varphi_{1j}-\varphi_{2j})].$ (109)
 $\varphi_{1j}-\varphi_{2j}$ is a random but fixed number for two independent BECs.
 The first-order interference pattern can be observed Andrews et al.
-(1997), just like the transient first-order interference pattern by superposing two independent laser light beams Magyar and Mandel (1963).
+(1997), just like there is transient first-order interference pattern by superposing two independent laser light beams Magyar and Mandel (1963).
 Similar method can be employed to calculate the first-order interference of multiple independent matter wave beams.
@@
 The scheme for the second-order interference of one matter wave beam is similar as the one shown in Fig.
-11 except the source S is a particle source, BS is a beam splitter for particles and these two detectors are particle detectors.
-First discuss the second-order interference of one boson particle beam.
+11 except that the source, S, is a particle source, BS is a beam splitter for particles and these two detectors are particle detectors.
+Let us first discuss the second-order interference of one boson beam.
 Assuming S is a cold atom source, there are two different paths to trigger a two-particle coincidence count.
-The first one is particle a is detected by D1 and particle b is detected by D2.
+The first path is that particle a is detected by D1 and particle b is detected by D2.
 The probability is 1/2 and the corresponding two-particle amplitude is $A_{a1,b2}$.
-The second path is particle a is detected by D2 and particle b is detected by D1.
+The second path is that particle a is detected by D2 and particle b is detected by D1.
 The probability is 1/2 and the corresponding two-particle amplitude is $A_{a2,b1}$.
@@
 $P^{(2)}(\vec{r}_{1},t_{1};\vec{r}_{2},t_{2})$ (110)
-$=$ $\langle|A_{a1,b2}+A_{a2,b1}|^{2}\rangle$
-$=$ $\langle|e^{i\varphi_{a}}K_{a1}e^{i\varphi_{b}}K_{b2}+e^{i\varphi_{a}}K_{a2}e^{i\varphi_{b}}K_{b1}|^{2}\rangle$
+$=$ $\langle|\frac{1}{\sqrt{2}}(A_{a1,b2}+A_{a2,b1})|^{2}\rangle$
+$\propto$ $\langle|e^{i\varphi_{a}}K_{a1}e^{i\varphi_{b}}K_{b2}+e^{i\varphi_{a}}K_{a2}e^{i\varphi_{b}}K_{b1}|^{2}\rangle$
 $=$ $\langle|K_{a1}K_{b2}+K_{a2}K_{b1}|^{2}\rangle.$
@@
 $P^{(2)}(x_{1}-x_{2})\propto 1+\text{sinc}^{2}\frac{\pi D}{L\lambda_{D}}(x_{1}-x_{2}),$ (111)
-where $D$ is the length of S and single speed particles are assumed. The temporal two-particle probability function is
+where $D$ is the length of S and single-speed particles are assumed.
+The temporal two-particle probability function is
 $P^{(2)}(t_{1}-t_{2})\propto 1+\text{sinc}^{2}\frac{\Delta\omega}{2}(t_{1}-t_{2}),$ (112)
 where $\Delta\omega=2mv_{0}\Delta v/\hbar$ is the angular frequency bandwidth, $v_{0}$ is the central speed, $\Delta v$ is the speed bandwidth. Point source is assumed to simplify the calculation.
-When these two detectors are at the same space-time coordinates, the probability to detect two particles is two times of the one when these two detectors are in different coherence volumes.
+When these two detectors are at the same space-time coordinate, the probability to detect two particles is two times of the one when these two detectors are in different coherence volumes.
 There is two-particle bunching for cold boson particles in a HBT interferometer Yasuda and Shimizu (1996), just like there is two-photon bunching for thermal light in a HBT interferometer.
-Assume S is a BEC in Fig.
+Assuming S is a BEC in Fig.
 11, there is only one path to trigger a two-particle coincidence count.
-Since the initial phases of particles in a BEC are identical, just like the one for single-mode continuous-wave laser light.
+The initial phases of the particles in a BEC are identical, just like the one for single-mode continuous-wave laser light.
 No two-particle bunching can be observed, which is consistent with Burt et al.’s experiments Burt et al.
 (1997).
-Then discuss the second-order interference of one fermion particle beam.
-There are also two different ways to trigger a two-particle coincidence count and the corresponding two-particle probability amplitude are $A_{a1,b2}$ and $A_{a2,b1}$.
+Then discuss the second-order interference of one fermion beam.
+There are also two different paths to trigger a two-particle coincidence count and the corresponding two-particle probability amplitudes are $A_{a1,b2}$ and $A_{a2,b1}$.
 If these two different ways are indistinguishable, the two-particle probability function is
 $P^{(2)}(\vec{r}_{1},t_{1};\vec{r}_{2},t_{2})$ (113)
-$=$ $\langle|A_{a1,b2}-A_{a2,b1}|^{2}\rangle$
-$=$ $\langle|e^{i\varphi_{a}}K_{a1}e^{i\varphi_{b}}K_{b2}-e^{i\varphi_{a}}K_{a2}e^{i\varphi_{b}}K_{b1}|^{2}\rangle$
+$=$ $\langle|\frac{1}{\sqrt{2}}(A_{a1,b2}-A_{a2,b1})|^{2}\rangle$
+$\propto$ $\langle|e^{i\varphi_{a}}K_{a1}e^{i\varphi_{b}}K_{b2}-e^{i\varphi_{a}}K_{a2}e^{i\varphi_{b}}K_{b1}|^{2}\rangle$
 $=$ $\langle|K_{a1}K_{b2}-K_{a2}K_{b1}|^{2}\rangle,$
-where the minus sign is due to fermions having exchange antisymmetry Griffiths and Schroeter (2018).
-By analogy of cold boson particles in a HBT interferometer, one-dimension spatial two-particle probability function is
+where the minus sign is due to the exchange antisymmetry of fermions Griffiths and Schroeter (2018).
+By analogy of cold boson particles in a HBT interferometer, one-dimension spatial two-particle probability function for fermions in a HBT interferometer is
 $P^{(2)}(x_{1}-x_{2})\propto 1-\text{sinc}^{2}\frac{\pi D}{L\lambda_{D}}(x_{1}-x_{2}),$ (114)
-where $D$ is the length of S and single speed is assumed. The temporal two-particle probability function of fermions in a HBT interferometer is
+where $D$ is the length of the source and single-speed particles are assumed.
+The temporal two-particle probability function of fermions in a HBT interferometer is
 $P^{(2)}(t_{1}-t_{2})\propto 1-\text{sinc}^{2}\frac{\Delta\omega}{2}(t_{1}-t_{2}),$ (115)
 where the meanings of symbols are similar as the ones in Eq. (112).
-When two detectors are at the same space-time coordinates, the probability to detect two fermions equals 0.
-As the distance between these two detectors increases, the probability to detect two fermions increases to 1.
-There is two-particle antibunching effect for fermion particles in a HBT interferometer Henny et al.
+When these two detectors are at the same space-time coordinates, the probability of detecting two fermions equals 0.
+As the distance between these two detectors increases, the probability of detecting two fermions increases to 1.
+There is two-particle antibunching effect for fermions in a HBT interferometer Henny et al.
 (1999); Jeltes et al.
-(2007), in contrast to the two-particle bunching effect for bosons in a HBT interferometer Jeltes et al.
+(2007), which is in contrast to the two-particle bunching effect for bosons in a HBT interferometer Jeltes et al.
 (2007); Brown and Twiss (1956b, a).
@@
 S1 and S2 are two independent particle sources.
-BS is a particle beam splitter.
+BS is a 1:1 particle beam splitter.
 D1 and D2 are two single-particle detectors.
 CC is two-particle coincidence count detection system.
-First discuss the second-order interference of two single-particle sources.
-Assuming S1 and S2 are two identical single-particle sources, there are two different paths to trigger a two-particle coincidence count.
-The first one is the particle emitted by S1 is detected by D1 and the particle emitted by S2 is detected by D2.
+First discuss the second-order interference of two single-boson sources.
+Assuming S1 and S2 are two identical and independent single-particle sources, there are two different paths to trigger a two-particle coincidence count.
+The first one is that the particle emitted by S1 is detected by D1 and the particle emitted by S2 is detected by D2.
 The probability is 1/2 and the two-particle probability amplitude is $A_{11,22}$.
-The second path is the particle emitted by S1 is detected by D2 and the particle emitted by S2 is detected by D1.
+The second path is that the particle emitted by S1 is detected by D2 and the particle emitted by S2 is detected by D1.
 The probability is 1/2 and the two-particle probability amplitude is $A_{12,21}$.
-If these different paths are indistinguishable, the two-particle probability function for bosons is
+If these two different paths are indistinguishable, the two-particle probability function for bosons is
 $P^{(2)}(\vec{r}_{1},t_{1};\vec{r}_{2},t_{2})$ (116)
-$=$ $\langle|A_{11,22}+A_{12,21}|^{2}\rangle$
-$=$ $\langle|e^{i\varphi_{1}}K_{11}e^{i\varphi_{2}}K_{22}+e^{i(\varphi_{1}+\frac{\pi}{2})}K_{12}e^{i(\varphi_{2}+\frac{\pi}{2})}K_{21}|^{2}\rangle$
+$=$ $\langle\frac{1}{\sqrt{2}}(|A_{11,22}+A_{12,21})|^{2}\rangle$
+$\propto$ $\langle|e^{i\varphi_{1}}K_{11}e^{i\varphi_{2}}K_{22}+e^{i(\varphi_{1}+\frac{\pi}{2})}K_{12}e^{i(\varphi_{2}+\frac{\pi}{2})}K_{21}|^{2}\rangle$
 $=$ $\langle|K_{11}K_{22}-K_{12}K_{21}|^{2}\rangle,$
-Substituting Eq. (104) into Eq. (116), one-dimension temporal two-particle probability function is
+where the extra $\pi/2$ phase is due to the reflection of BS comparing to the transmission like the one for photons.
+Substituting Eq.
+(104) into Eq.
+(116), temporal two-particle probability function is
 $P^{(2)}(t_{1}-t_{2})\propto 1-\text{sinc}^{2}\frac{\Delta\omega}{2}(t_{1}-t_{2}),$ (117)
-where point particle source is assumed.
-It is impossible to detect two boson particles when these two detectors are at symmetrical positions.
-HOM dip with visibility higher than 50% can be observed Lopes et al.
+where point particle sources are assumed.
+It is impossible to detect two boson particles when these two detectors are at symmetrical positions of BS.
+Particle HOM dip with visibility higher than 50% was observed Lopes et al.
 (2015).
-With the same method, one-dimension temporal two-particle probability function for fermions is
+With the same method, temporal two-particle probability function in Fig. 18 for fermions is
 $P^{(2)}(t_{1}-t_{2})\propto 1+\text{sinc}^{2}\frac{\Delta\omega}{2}(t_{1}-t_{2}),$ (118)
@@
 There are three different ways to trigger a two-particle coincidence count.
-The first one is these two particles are both emitted by S1.
+The first way is that these two particles are both emitted by S1.
 The probability is 1/4 and there is one path, $A_{11,12}$.
-The second one is these two particles are both emitted by S2.
+The second one is that these two particles are both emitted by S2.
 The probability is 1/4 and there is one path, $A_{21,22}$.
-The third way is these two particles are emitted by S1 and S2 one each.
+The third way is that these two particles are emitted by S1 and S2 one each.
 The probability is 1/2 and there are two different paths, $A_{11,22}$ and $A_{12,21}$.
@@
 $\propto$ $\langle|K_{11}K_{12}|^{2}\rangle+\langle|K_{21}K_{22}|^{2}\rangle+\langle|(K_{11}K_{22}-K_{12}K_{21})|^{2}\rangle.$
-Substituting Eq. (104) into Eq. (VI.2.2), it is straightforward to have temporal two-particle probability function
+Substituting Eq. (104) into Eq. (VI.2.2), it is straightforward to have temporal two-particle probability function,
 $P^{(2)}(t_{1}-t_{2})\propto 1-\frac{1}{2}\text{sinc}^{2}\frac{\Delta\omega}{2}(t_{1}-t_{2}).$ (121)
-Compared to the second-order interference of two single boson sources, Eq.
-(117), the visibility of the second-order interference of two independent BECs can not exceed 50%, which is similar as the conclusion for the visibility of the second-order interference of two independent classical and nonclassical light beams Mandel (1983).
+Comparing to the second-order interference of two single boson sources, the visibility of the second-order interference of two independent BECs can not exceed 50%, which is similar as the conclusion for the visibilities of the second-order interference of two independent classical and non-classical light beams Mandel (1983).
 VI.3 Third-order interference of matter wave
-Based on the results above, it is easy to predict that there is three-particle bunching when a cold boson particle beams in a third-order HBT interferometer and no three-particle bunching for a BEC Burt et al.
+Based on the results above, it is easy to predict that there is three-particle bunching for a cold boson beam and no three-particle bunching for a BEC in a third-order HBT interferometer Burt et al.
 (1997).
-This section will concentrate on the third-order interference of fermions.
+This section will focus on the third-order interference of fermions.
 The scheme for the third-order interference of one fermion beam is the same as the one shown in Fig.
@@
 $P^{(3)}(\vec{r}_{1},t_{1};\vec{r}_{2},t_{2};\vec{r}_{3},t_{3})$ (122)
-$=$ $\langle|A_{a1,b2,c3}-A_{a1,b3,c2}-A_{a2,b1,c3}$
-$+A_{a2,b3,c1}+A_{a3,b1,c2}-A_{a3,b2,c1}|^{2}\rangle,$
+$=$ $\langle|\frac{1}{\sqrt{6}}(A_{a1,b2,c3}-A_{a1,b3,c2}-A_{a2,b1,c3}$
+$+A_{a2,b3,c1}+A_{a3,b1,c2}-A_{a3,b2,c1})|^{2}\rangle,$
 where the minus signs are due to exchange antisymmetry of fermions. With similar method above, temporal three-particle probability function is
@@
 Figure 35: Third-order temporal coherence function of fermions in a third-order HBT interferometer.
+$g^{(3)}(t_{1},t_{2},t_{3})$ is the normalized third-order temporal coherence function.
+$t_{1}-t_{2}$ is the time difference between photon detection event at D1 and D2.
+Similar definition holds for $t_{2}-t_{3}$.
 Figure 35 shows the normalized third-order temporal coherence function of fermions in a third-order HBT interferometer, which is proportional to Eq.
 (VI.3).
-The probability to detect three particles equals 0 when these three detectors are at the symmetrical positions, which is consistent with the results obtained by Liu et al.
+The probability of detecting three particles when these three detectors are at the symmetrical positions equals 0, which is consistent with the results obtained by Liu et al.
 Liu (2016).
 Similar method can be employed to calculate the third- and higher-order interference of multiple matter wave beams.
-VII Summary
-In summary, we have presented a new quantum optical coherence theory based on Feynman’s path integral, which provides a useful tool for studying the physics of coherence besides classical optical coherence theory based on Maxwell’s electromagnetic theory and Glauber’s quantum optical coherence theory based on wave mechanics formalism of quantum mechanics.
+VII Conclusions
+It is interesting to notice that there is some difference between the points of view about the interference of light in classical optical coherence theory and quantum optical coherence theory based on Feynman’s path integral.
+In classical optical theory, the first-order interference of light is the foundation of the second- and higher-order interference of light.
+All the second- and higher-order interference of light is interpreted as the results of first-order interference of light.
+For instance, two-photon bunching of thermal light is interpreted as the intensity fluctuations correlation in classical optical coherence theory.
+In quantum optical coherence theory based on Feynman’s path integral, the first-, second-, and higher-order interference of light are interpreted by the same superposition principle.
+The first-order interference of light is not the foundation of second- and higher-order interference of light.
+There is no right and wrong about these two different points of view.
+It is just the interpretations about optical coherence are different in these two different theories.
+In summary, we have presented a novel quantum optical coherence theory based on Feynman’s path integral, which provides a useful tool for studying the physics of optical coherence besides classical optical coherence theory based on Maxwell’s electromagnetic theory and Glauber’s quantum optical coherence theory based on matrix mechanics formalism of quantum mechanics.
 Some interesting conclusions are drawn.
 For instance, unlike the well-accepted conclusion that the electric field models are the same for thermal and laser light within the coherence time, it is predicted by quantum optical coherence theory based on Feynman’s path integral that there may be some difference between these two.
-The physics behind two-photon bunching of thermal light in a HBT interferometer and HOM dip of entangled photon pairs in a HOM interferometer is the same, which can be interpreted by two-photon interference.
-In order to show how quantum optical coherence theory based on Feynman’s path integral can help us to understand the physics of optical coherence better, subwavelength interference is taken as an example to be illustrated and it is shown that there is no intrinsic difference between the observed second-order subwavelength interference for two detectors scanning in the same direction and opposite directions.
-Finally, quantum particle coherence theory based on Feynman’s path integral is also given by analogy of the one in optics.
-We believe quantum optical coherence theory based on Feynman’s path integral will be helpful to understand the physics of optical and atomic coherence in other fields, and eventually lead us to the answer to the question: what is photon?
+The physics behind two-photon bunching of thermal light in a HBT interferometer and HOM dip of entangled photon pairs in a HOM interferometer is the same, both of which can be interpreted by two-photon interference.
+In order to show how quantum optical coherence theory based on Feynman’s path integral can be helpful to understand the physics of optical coherence, subwavelength interference is taken as an example and it is shown that there is no intrinsic difference between the observed two-photon subwavelength interference for two detectors scanning in the same direction and opposite directions.
+Finally, quantum particle coherence theory based on Feynman’s path integral is also given by analogy of the developed quantum optical coherence theory based on Feynman’s path integral.
+We believe that quantum optical coherence theory based on Feynman’s path integral is helpful to understand the physics of optical and atomic coherence, and may eventually lead us to the answer to the question puzzling us for a very long time: what is light?
 Acknowledgments